### v1

# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

Abstract. Saturated sets and its reduced case the set of generic points are two types of important fractal-like sets in Multifractal analysis of dynamical systems. In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set generic points in both topological and measure-theoretic situations. For systems with specification property, we establish the certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets have full metric mean dimension. All results are useful for understanding the topological structures of dynamical systems with infinite topological entropy.
As applications, we further exhibit some qualitative aspects of metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
Keywords and phrases: Metric mean dimension; Saturated sets; Generic point; Variational principle; Multifractal analysis
1 Introduction
Let $(X,d,f)$ be a topological dynamical system (t.d.s. for short), where $(X,d)$ is a compact metric space with metric $d$, and $f:X\to X$ is a homeomorphism. Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(x)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively. Measure-theoretic entropy and topological entropy are two vital invariants for us to understand the topological complexity of dynamical systems, which are related by the classical variational principle:
$h_{top}(f)=\sup_{\mu\in M_{f}(X)}{h_{\mu}(f)},$
where $h_{top}(f)$ is the topological entropy of $(X,f)$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$. In 1973, analogous to the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space. The pioneering work directly contributes to the development of dimension theory of dynamical systems, especially as a powerful tool to study the “size” of fractal-like sets in multifractal analysis.
The multifractal analysis, which can date back to Besicovitch, considers the decomposition of the whole phase space into smaller subset consisting of the points with the similar dynamical behaviors, or known as multifractal spectrum, and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies. Such a procedure allows us to get more information about the dynamics and even recover the partial dynamical behaviors of the system, which we call such phenomenon multifractal rigidity. We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
Let us state some backgrounds about the dimensional entropies of saturated sets. Given $x\in X$, by $\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X)$, where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of the Birkhoff average w.r.t. $x$. Then the limit-point set $V_{f}(x)$ of $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ in the sense of weak-topology (cf. [DGS76, Proposition 3.8]). Specially, when $V_{f}(x)=\{\mu\}$ for some measure $\mu$, one says that $x$ is a generic point of $\mu$. The set of all generic points of $\mu$ is denoted by $G_{\mu}$. Then the Birkhoff ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic. Bowen [Bow73] showed for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f).$
For non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$ and even easily constructs the example showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$ such that Bowen’s result is not valid. It is natural to expect whether there are some proper conditions such that the analogous equality still holds for non-ergodic measures. A subset $D$ of $X$ is called a saturated set if $x\in D$ and $V_{f}(x)=V_{f}(y)$, then $y\in D$. More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, this defines a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$. For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems [LST13, TV17]. As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and uniform separation property (e.g. expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any compact connected non-empty set $K\subset M_{f}(X)$,
$h_{top}^{B}(T,G_{K})=\inf\{h_{\mu}(f):\mu\in K\}.$
Specially, only the assumption of $g$-almost product property they [PS07] showed
$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$
holds for any $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to any non-ergodic measures. Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional entropies of saturated sets. Packing entropy of non-compact subset was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension. Under a stronger condition of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] further established the following variational principle for any compact connected non-empty set $K\subset M_{f}(X)$,
$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\},$
where $h_{top}^{P}(f,K)$ is the packing entropy of $K$. Later, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
$h_{top}^{UC}(f,G_{K}^{C})=h_{top}(f),$
where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$. Besides, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle still hold for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points. Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results about the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
According to the possible values of topological entropy, dynamical systems can be divided into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems. Infinite entropy system always means the system may possess quiet complicated topological complexity, so the notion of topological entropy no longer provides more information about the dynamics of the system. In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00. It turns out that any finite entropy system has zero metric mean dimension. See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b]. The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, dues to Gutman and Śpiewak[GS21]. More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
$\underline{\rm MRID}_{\mu}(f,X,d):$ $=\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
$\overline{\rm MRID}_{\mu}(f,X,d):$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
which can be explained how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}\xi<\epsilon}\limits h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$. Some variational principles of metric mean dimension of $\mathbb{Z}$-actions can be found in [LW00, GS21, S22, YCZ23a]. Then the aforementioned work suggests the following questions for infinite entropy systems:
Questions:
  • (1)
    For ergodic measures, do the metric mean dimensions of the set of generic points equal to the mean Rényi information dimensions?
  • (2)
    For non-ergodic measures, under the assumption of specification-like property, are there some certain variational principles for metric mean dimensions of saturated sets? if not, can metric mean dimensions of saturated sets have the full metric mean dimension?
For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the setting of infinite entropy systems encounters some nontrivial challenges. Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ containing in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
However, for systems with specification property we can give a full answer to question (2). Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the following statements of main results involve more general (mixed) saturated sets. For simplicity, we write
$G_{K}^{\omega,U}:=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},$
where $U\subset X$ is a nonempty open subset, $K\subset M_{f}(X)$ is a nonempty closed connected (or convex) subset. If $G_{K}^{\omega,U}\not=\emptyset$, then there exists a transitive point in the open set $U$ such that $V_{f}(x)=K$.
Different types of metric mean dimensions of saturated sets exhibit different dynamical behaviors. There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated sets has full metric mean dimension (Theorem 1.4).
Theorem 1.1.
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$, and the infimum is taken over all finite Borel measurable partitions of $X$.
Theorem 1.2.
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
where $\overline{\rm mdim}_{M}^{B}(f,Z,d)$ is the Bowen upper metric mean dimension of $Z$.
Maybe, more satisfactory results are exchanging the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$), but until now the authors do not know whether there are some counter-examples of infinite entropy systems with specification property such that the following equalities are strict:
$\sup_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d),$
and
$\inf_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d).$
Inspired the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures. For any closed set $K$ of $M_{f}(X)$ and $\mu\in K$, by $M_{K}(\mu)$ we denote the space of sequences of Borel probability measures in $K$ which converge to $\mu$ in the weak-topology. We respectively define the lower and upper mean Rényi information dimensions of $\mu\in K$ as
$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi),$
$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
Then we have the following desired variational principles in terms of the modified lower and upper mean Rényi information dimensions.
Theorem 1.3.
Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\},$
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
We remark that Theorem 1.3 does not need the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
Theorem 1.4.
Let $(X,d,f)$ be a dynamical system with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set, then
$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}(f,X,d),$
where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$ defined by separated sets.
As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being just a single point. The Corollary 1.5 presents the metric mean dimensions of generic points of any invariant measures for systems with specification property, which extends the result given in [YCZ24, Theorem 1.1] to all non-ergodic measures.
Corollary 1.5.
Let $(X,d,f)$ be a dynamical system with specification property. Then for any $\mu\in M_{f}(X)$,
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),$
$\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$, then we have
$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
The “Consequently” part can realize for some dynamical systems. For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$. The concept of maximal metric mean dimension measure was proposed in [YCZ23a], yet for what kind of infinite entropy systems can have maximal metric mean dimension measure is still a difficult problem. See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms. For some certain systems, by Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
Corollary 1.6.
Let $(X,d,f)$ be a dynamical system with specification property. An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$ if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
Some comments are in order.
  • (1)
    There are plenty of infinite entropy systems with specification property, for instance, the usual the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
  • (2)
    In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension. The other aforementioned theorems are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
  • (3)
    We believe it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs. The novelty of our work not only presents a proper and full extension of the previous work in [PS07, ZCC12, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for the multifractal analysis of other types of fractal-like sets in infinite entropy systems. It remains unclear whether there are some proper substitutions of the uniform separation property so that all results are true for $g$-almost product property, or beyond $\mathbb{Z}$-actions.
The paper is organized as follows. In Section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property. In section 3, we prove the main results. In Section 4, we give two applications of main results. In section 5, we show the metric mean dimensions of generic points of full shift over compact metric spaces.
2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and the precise definition of specification property in dynamical systems.
2.1 Three types of metric mean dimensions
In this subsection, we recall the definitions of upper capacity metric mean dimension on subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension on subsets [YCZ22].
Let $(X,d,f)$ be a dynamical system. For $n\in N$, we define the Bowen metric $d_{n}$ on $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\}$
for any $x,y\in X$. Then the Bowen open ball and closed ball of $x$ in the metric $d_{n}$ with radius $\epsilon$ are respectively given by
$B_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)<\epsilon\},$
$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}.$
Let $Z$ be a nonempty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if $x,y\in E$ with $x\neq y$ implies $d_{n}(x,y)>\epsilon$. Let $s_{n}(Z,\epsilon)$ denote the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
$h_{top}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
Analogous to the definition of box dimension in fractal geometry, we can formulate the definition of upper capacity metric mean dimension of subsets using separated sets.
Definition 2.1.
We define the upper capacity metric mean dimension of $Z$ as
$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$ (2.1)
When $Z=X$, we sometimes write $h_{top}(f,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, $\overline{\rm mdim}_{M}^{UC}(f,X,d)=\overline{\rm mdim}_{M}(f,X,d)$ respectively, then in this case it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00]. Such a formulation allows us to extend the metric mean dimensions to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Recall that the topological entropy of $f$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
$h_{top}(f)=\lim_{\epsilon\to 0}h_{top}(f,\epsilon)=\sup_{\epsilon>0}h_{top}(f,\epsilon).$
For any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity. Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite entropy of the system as $\epsilon$ tends to 00. More precisely, for sufficiently $\epsilon>0$, we may think of
$h_{top}^{UC}(f,Z,\epsilon)\approx\overline{\rm mdim}_{M}^{UC}(f,Z,d)\cdot|\log\epsilon|.$
It is easy to see that any finite entropy system has zero metric mean dimension. This further verifies that metric mean dimension is a good candidate to capture the dynamics of infinite systems.
Similar to the definitions of Hausdorff dimension and packing dimension in fractal geometry, as an analogue of the classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which is defined by Carathéodory-Pesin structures, the corresponding dimensional characterizations of metric mean dimensions via Carathéodory-Pesin structures are Bowen and packing metric mean dimensions, which has been introduced in [LV21, YCZ22, W21] and studied by several authors in [BR23, CLS21, LL24, FKO24, YCZ24].
Definition 2.2.
For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i}\exp(-n_{i}s)\right\},$
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}$ such that $Z\subset\cup_{i}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
The quantity $M(Z,s,N,\epsilon)$ does not decrease as $N$ increase. So we set
$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon).$
There is a critical value of parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
$h_{top}^{B}(f,Z,\epsilon)$ $=\inf\{s>0:M(Z,s,\epsilon)=0\}$
$=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
We define the Bowen upper metric mean dimension of $Z$ as
$\overline{\rm mdim}^{B}_{M}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{B}(f,Z,\epsilon)}{|\log\epsilon|}.$
Definition 2.3.
For $Z\subseteq X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i}\exp(-sn_{i})\right\},$
where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i$.
Since the quantity $P(Z,s,N,\epsilon)$ does not increase as $N$ increases, the following limit exists:
$P(Z,s,\epsilon)=\lim_{N\to\infty}P(Z,s,N,\epsilon).$
Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So some modifications are needed to obtain the property:
$\mathcal{P}(Z,s,\epsilon):=\inf\left\{\sum_{i=1}^{\infty}P(Z_{i},s,\epsilon):\bigcup_{i=1}^{\infty}Z_{i}\supseteq Z\right\}.$
There is a critical value of parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
$h_{top}^{P}(f,Z,\epsilon)$ $=\inf\{s>0:\mathcal{P}(Z,s,\epsilon)=0\}$
$=\sup\{s>0:\mathcal{P}(Z,s,\epsilon)=\infty\}.$
We define the packing upper metric mean dimension of $Z$ as
$\overline{\rm mdim}_{M}^{P}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,Z,\epsilon)}{|\log\epsilon|}.$
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to defined the corresponding three types of lower metric mean dimensions on subsets. We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6] as follows:
Proposition 2.1.
$(1)$ If $Z_{1}\subset Z_{2}$, then
$\overline{\rm mdim}_{M}^{UC}(f,Z_{1},d)$ $\leq\overline{\rm mdim}_{M}^{UC}(f,Z_{2},d)$
$\overline{\rm mdim}^{B}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{B}_{M}(f,Z_{2},d)$
$\overline{\rm mdim}^{P}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{P}_{M}(f,Z_{2},d).$
$(2)$ If $Z=\bigcup_{i=1}^{\infty}Z_{i}$, then for every $\epsilon>0$,
$h_{top}^{B}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{B}(f,Z_{i},\epsilon)$
$h_{top}^{P}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{P}(f,Z_{i},\epsilon),$
and $h_{top}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then
$\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{UC}(f,Z_{i},d)$
$\overline{\rm mdim}^{B}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{B}_{M}(f,Z_{i},d)$
$\overline{\rm mdim}^{P}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{P}_{M}(f,Z_{i},d).$
$(3)$ If $Z$ is an any non-empty subset of $X$, then for every $10<\epsilon<$ 1, we have
$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}(f,Z,\epsilon).$
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$; if $Z$ is $f$-invariant and compact, then the equalities hold.
Observe that the metric mean dimension of a given subset is in general hard to calculate its precise value. However, the Entropy Distribution Principle allows us to get the lower bounds of the $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems. Consequently, the powerful tool enables us to get the lower bounds of entropy and metric mean dimension of the fractal-like sets.
Lemma 2.1.
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exists a Borel probability measure $\mu$ and a constant $C>0$ satisfying:
$(1)$ for every $n$, $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for each Bowen ball $B_{n}(x,\epsilon)$ such that $B_{n}(x,\epsilon)\cap Z\neq\emptyset$;
$(2)$ $\mu(Z)>0$,
then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
$\epsilon$-packing topological entropy also admits an analogous Entropy Distribution Principle as follows:
Lemma 2.2.
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exists a Borel probability measure $\mu$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
$(1)$ $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$;
$(2)$ $\mu(Z)>0$,
then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
Proof.
For sufficiently small $\delta>0$, by definition it is enough to prove that
$P\left(E,s-\delta,\frac{\epsilon}{6}\right)=\infty$
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$. For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$. Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint sub-collection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that
$E\subset\bigcup_{j}B_{n_{i}}(x_{j},\epsilon).$
Hence by condition $(1)$, one has
$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$. ∎
2.2 Measure-theoretic $\epsilon$-entropy
In this subsection, we recall the definitions of Kolmogorov-Sinai $\epsilon$-entropy and Katok $\epsilon$-entropy of invariant measures.
Let $C(X)$ denote the space of continuous functions on $X$ equipped with the supermum norm. The Riesz representation theorem on compact metric space gives us an approach to endow with the weak-topology on $M(X)$, which is metrizable by the following metric
$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$
where $\{\phi_{k}\}$ is a countable dense subset of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in sense of metric $D$ is always less than $1$, and for fixed $\mu$ and $\nu$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural projection from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in sense of weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, we say a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$. The partition entropy of $\xi$ w.r.t. $\mu$ is defined by
$H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$
Then we define measure-theoretic entropy of $f$ w.r.t. $\xi$ as
$h_{\mu}(f,\xi)=\lim_{n\to+\infty}\frac{1}{n}H_{\mu}(\xi^{n}),$
where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $f$ w.r.t. $\mu$ is defined as
$h_{\mu}(f)=\sup_{\xi}h_{\mu}(f,\xi),$
where the supremum is taken over all finite Borel measurable partitions $\xi$ of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions on subsets is the measure-theoretic metric mean dimensions of invariant measures. Unlike topological entropy, measure-theoretic entropy of ergodic measures admit several equivalent definitions by considering the limits of the measure-theoretic $\epsilon$-entropies: for instance, Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ24], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf. [YCZ24, Theorem 1.1]). Here, we only recall two types of them.
  • (a)
    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu$ is given by
    $\inf_{{\rm diam}\xi<\epsilon}h_{\mu}(f,\xi).$
  • (b)
    For $\mu\in M_{f}(X)$, $\epsilon>0$, $\delta\in(0,1)$ and $n\in\mathbb{N}$, we put
    $N_{\mu}(n,\epsilon,\delta):=\min\{\#E:E\subset X~{}\emph{and}~{}\mu(\cup_{x\in E}B_{n}(x,\epsilon))>1-\delta\}.$
    The (lower) Katok $\epsilon$-entropy of $\mu$ is given by
    $\underline{h}_{\mu}^{K}(f,\epsilon,\delta)=\liminf_{n\to\infty}\frac{1}{n}\log N_{\mu}(n,\epsilon,\delta).$
The relation between Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy is related by the following inequalities:
Lemma 2.3.
[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})<4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$. Then for any $\delta\in(0,1)$, one has
$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
Consequently, $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
The next variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
Theorem 2.4.
(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with metric $d$. Then for any $\delta\in(0,1)$,
$\overline{\rm mdim}_{M}(f,X,d)$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(X)}F(\mu,\epsilon),$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}^{e}(X)}F(\mu,\epsilon),$
where $F(\mu,\epsilon)\in\{\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),\underline{h}^{K}_{\mu}(f,\epsilon,\delta)\}$.
A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is given in [YCZ24, Theorems 1.2 and 1.3].
2.3 Specification property
Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments using specification-like property.
Definition 2.4 (Specification property).
One says that $f$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
$d(f^{p+a_{j}}x,f^{p}x_{j})<\epsilon$
for any $p=0,\cdots,b_{j}-a_{j}$, $j=1,\cdots,k$.
In other words, that point $x$ $\epsilon$-shadows the $k$ pieces of orbits
$\{x_{j},f(x_{j}),\ldots,f^{b_{j}-a_{j}}(x_{j})\},\ j=1,\ldots,k.$
For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $[0,1]^{\mathbb{Z}}$ have specification property. More examples of infinite entropy systems with specification property can be found in [BR23, Section IV].
A point $x\in X$ is almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, it holds that $f^{k}(x)\in U$ for some $n\leq k\leq n+N$. Let $AP(X)$ denote the set of all almost periodic points of $X$.
The following result, which was proved for systems with $g$-almost product property [HTZ23, Proposition 2.11], gives a characterization of the structure of $AP(X)$.
Proposition 2.2.
If the dynamical system $(X,f)$ has specification property, then the almost periodic set AP(X) is dense in $X$.
3 Proofs of main results
This section aims to prove the main results of this paper. Essentially, the difficult part is the estimation of lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$. We shall construct a Moran fractal set containing in $G_{K}^{\omega,U}$, which is a limit set with geometric construction. Then such set is suitable for the purpose of the computation of metric mean dimensions of saturated sets.
3.1 Packing metric mean dimension of saturated sets
We prove Theorem 1.1. To arrive this goal, we divide the proof ot Theorem 1.1 into two parts.
3.1.1 Upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$
We adapt a general method presented in [PS07, Theorem 4.1] [ZCC12, Theorem 1.1 $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
Lemma 3.1.
[Y23, Lemma 3.1] Let $(X,f)$ be a dynamical system and $\epsilon>0$. Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$ and $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality. We define
$v_{n}:=\frac{1}{\#E_{n}}\sum_{x\in E_{n}}\mathcal{E}_{n}(x).$
Then for any limit point $\mu$ of the sequence $\{\nu_{n}\}$, one has
$\limsup_{n\to\infty}\frac{1}{n}\log\#E_{n}\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
This essentially follows the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 9.10]). ∎
For $K\subset M_{f}(X)$, we set $G^{K}:=\{x\in X:V_{f}(x)\subset K\}$. Obviously, we have the fact $G_{K}\subset G^{K}$.
Lemma 3.2.
Let $(X,f)$ be a t.d.s..
$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$, we have
$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$
$(2)$ If $\mu\in M_{f}(X)$, then for any $\epsilon>0$,
$h_{top}^{P}(f,G_{\mu},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Proof.
It suffices to show $(1)$, since one can take $K=\{\mu\}$ as in $(1)$ and use the fact $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
For $n\in\mathbb{N}$, $\delta>0$, we put
$\mathcal{R}(K,\delta,n):=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{R}(K,\delta,n)$. Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}(K,\delta,n)$ holds for every $\delta$, it follows from Proposition 2.1 that
$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}(K,\delta,n),\epsilon)\leq\sup_{k}h_{top}(f,\mathcal{R}(K,\delta,n),\epsilon).$
Next we shall prove for any $\gamma>0$, there is $\mu\in K$ such that
$\Theta(K,\epsilon)-\gamma\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Take a strictly decreasing positive sequence $\{\delta_{k}\}$ with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
$S(K,\delta_{k},n_{k},\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}(K,\delta_{k},n_{k})$, and set
$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
Choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed convex set. By Lemma 3.1, we have
$\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\geq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\geq\Theta(K,\epsilon)-\gamma.$
This completes the proof. ∎
Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of saturated set:
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,G_{K},\epsilon)}{|\log\epsilon|}\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.1)
3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
For any small enough $\epsilon^{\ast}>0$, we shall prove that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$ (3.2)
One may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$. Then take $\mu^{\ast}\in K$ such that
$\inf_{\mathrm{diam}(\xi)<20\epsilon^{\ast}}h_{\mu^{\ast}}(f,\xi)>H-\zeta.$ (3.3)
Since $K$ is connected and compact, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
[…]
$D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})$ $\leq\frac{1}{k},\ \mu_{(k,Q_{k})}=\mu^{\ast}.$
Now set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order. Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$. For $(k,j)\in\mathcal{A}$, we shall use $(k,j)^{\ast}$ and $(k,j)_{\ast}$ to represent the one before and the one after $(k,j)$, respectively. Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$. We shall write $\delta_{k}:=\delta_{(k,Q_{k})}$. For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is a positive integer given in the definition of specification property.
For any non-empty open set $U$, there exist $\tilde{\epsilon}>0$, a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$. One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$. Set
$\mathcal{C}_{(0,0)}=\{a_{0}\},\ n_{(0,0)}=1,\ N_{(0,0)}=1.$
For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$ which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2. Then according to specification property, there exists $a_{k}\in X$ such that
$d\left(f^{(i-1)m_{(k,0)}+i-1}(a_{k}),a_{i}^{k}\right)<\epsilon_{(k,0)},\ i=1,2,\ldots,q_{k}.$
Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$, $N_{(k,0)}=1$.
Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. The following measure approximation derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
Lemma 3.3.
For each $k\in\mathbb{N}$, there is a measure $\nu_{k}\in M_{f}(X)$ satisfying
$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\{\lambda_{i}^{k}:1\leq i\leq j_{k}\}$ are rational positive numbers, $\forall 1\leq i\leq j_{k}$ with $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$ and $\nu_{i}^{k}\in M_{f}^{e}(X)$;
$(2)$ one has
$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta.$ (3.4)
$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$.
By the ergodicity of $\nu_{i}^{k}$, there exists $N\in\mathbb{N}$ such that the set
$Y_{i}(N)=\left\{x\in X:D(\mathcal{E}_{n}(x),\nu_{i}^{k})<\delta_{k},\ \forall n>N\right\}$
has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3 one has
$\underline{h}_{\nu_{i}^{k}}^{K}(f,20\epsilon^{\ast},\zeta)\leq\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)\leq\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)<\infty.$
Now take an integer $n_{k}^{\prime}\in\mathbb{N}$ such that
$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer holds for all $1\leq i\leq j_{k}$, and
$\frac{n_{k}^{\prime}}{n_{k}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\},$ (3.5)
where $n_{k}=n_{(k,Q_{k})}=n_{k}^{\prime}+(j_{k}-1)m_{(k,Q_{k})}$.
$(2)$ there exists a subset $E_{k,i}$ which is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
$(3)$ for every $x\in E_{k,i}$,
$D(\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x),\nu_{i}^{k})<\delta_{k}.$
By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ so that
[…]
where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$. One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniform continuous.
For any distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$. Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$,
$5\epsilon^{\ast}<d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},x_{i}^{\prime}\right)\leq d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},f^{T_{i}}y\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y^{\prime},x_{i}^{\prime}\right)$
$<\frac{\epsilon^{\ast}}{2^{k+1}}+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right),$
which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$. Let
$\mathcal{C}_{(k,Q_{k})}=\left\{y(x_{1},x_{2},\ldots,x_{j_{k}}):\ (x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}\right\}.$
Then $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set and
[…]
The last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and formulas (3.3), (3.5).
For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$,
$\left|\int\varphi\mathop{}\!\mathrm{d}\mathcal{E}_{n_{k}}(y)-\int\varphi\mathop{}\!\mathrm{d}\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y))\right)\right|$
$\leq\left|\frac{1}{n_{k}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))-\frac{1}{n_{k}^{\prime}}\sum_{i=1}^{j_{k}}S_{\lambda_{i}^{k}n_{k}^{\prime}}\varphi(f^{T_{i}}(y))\right|+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\|\varphi\|$
$\leq\left[\left|\frac{1}{n_{k}^{\prime}}-\frac{1}{n_{k}}\right|\cdot n_{k}^{\prime}+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\right]\cdot\|\varphi\|<\frac{2}{k}\|\varphi\|,$
which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Recall we have
$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k}.$
By the construction, one has
$D\left(\mathcal{E}_{n_{k}}(y),\mu^{\ast}\right)\leq\frac{2}{k}+D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i})\right)+2\delta_{k}<\frac{3}{k}+2\delta_{k}.$ (3.7)
Until now, using the specification property we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the remaining cases for $1\leq j<Q_{k}$.
Since the system has specification property, $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$. Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist a point $x_{k,j}$ and $n_{(k,j)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,j)}}(x_{k,j}),\mu_{(k,j)})<\delta_{(k,j)},$
and
$\frac{n_{(k,j)}}{n_{(k,j)}+m_{(k,j)}}>1-\frac{1}{k}.$
Now set $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
Step $1$. Construction of intermediate sets $\mathcal{S}_{(k,j)}$.
Recall that we have defined $N_{(k,0)}=1$ for any $k\geq 0$. Now choose a strictly increasing sequence of integers $\{N_{(k,j)}\}_{(k,j)\in\mathcal{A},~{}j\neq 0}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k},\\ n_{k}N_{(k,Q_{k})}&\geq(1-\zeta)\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}].\end{split}$ (3.8)
For $(k,j)\in\mathcal{A}$ with $j=0$, set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
Case $1$. $j\neq Q_{k}$.
Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$. By specification property, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with gap $m_{(k,j)}$. Now set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
Case $2$. $j=Q_{k}$.
Enumerate the points in $\mathcal{C}_{(k,Q_{k})}$ as
$\mathcal{C}_{(k,Q_{k})}=\{y_{i}^{k}:i=1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}.$
For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$. Then define
$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}.$
Write $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$, then it follows from the facts that $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set, and $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set. In particular, by (3.6) and (3.8), one has
#$\#\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.9)
Similar to the argument in (3.7) one has
$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ y_{k,j}\in\mathcal{S}_{(k,j)}\ \text{for }k\geq 1\ \text{and }1\leq j<Q_{k},$
and
$D(\mathcal{E}_{c_{k,Q_{k}}}(y),\mu^{\ast})<\frac{5}{k}+2\delta_{k},\ \forall y\in\mathcal{S}_{(k,Q_{k})}.$
Hence, in an unified way we always have
$D(\mathcal{E}_{c_{k,j}}(y),\mu_{(k,j)})<\frac{5}{k}+2\delta_{(k,j)},\ \forall y\in\mathcal{S}_{(k,j)}\ \text{with }k\geq 1,\ 1\leq j\leq Q_{k}.$ (3.10)
Step $2$. Constructions of $\mathbb{F}_{(k,j)}$ and Moran set $\mathbb{F}$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how construct $\mathcal{I}_{(k,j)}$. Let
$M_{(k,j)}$ $=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j};$
[…]
Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in formula (3.9) has the estimation:
#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$
For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by specification property let $z=z(x,y)\in X$ be some point such that
$d_{M_{(k,j)^{\ast}}}(x,z)<\epsilon_{(k,j)},\ d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}}+m_{(k,j)}}(z))<\epsilon_{(k,j)}.$
We set $\mathcal{I}_{(k,j)}=\{z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$. For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
$d_{M_{k,Q_{k}}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
which leads to
#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$
Moreover, according to (3.10), for $z\in\mathcal{I}_{(k,j)}$ with $j\neq 0$, we have
$D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(z)),\mu_{(k,j)})<\frac{6}{k}+2\delta_{(k,j)}.$ (3.11)
Now set
$\mathbb{F}_{(k,j)}=\bigcup_{x\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)}).$
It is easy to see that for each $(k,j)\in\mathcal{A}$ the following is satisfied:
$(1)$ $\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)})\cap\overline{B}_{M_{(k,j)}}(x^{\prime},2\epsilon_{(k,j)})=\emptyset$ holds for any distinct $x,x^{\prime}\in\mathcal{I}_{(k,j)}$;
$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has $\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}})$. That is $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$.
Let
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}=\bigcap_{k\in\mathbb{N}}\mathbb{F}_{(k,Q_{k})}.$
Lemma 3.4.
One has the fractal set $\mathbb{F}$ is contained in $\mathbb{F}\subset G_{K}^{\omega,U}$, that is,
$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:X\subset\omega_{f}(x)\}.$
Proof.
Firstly, we know $\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U$. For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there are some $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:X\subset\omega_{f}(x)\}$.
For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, in order to estimate $D(\mathcal{E}_{n}(p),K)$, we split into four cases to discuss.
Case $(1)$. If there exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,Q_{k})}$ such that $p\in\overline{B}_{M_{(k,Q_{k})}}(z,\frac{\epsilon_{(k,Q_{k})}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{c_{k,Q_{k}}}{n}D(\mathcal{E}_{c_{k,Q_{k}}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast})$
$+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast}).$
Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.11), we have $D(\mathcal{E}_{c_{k,Q_{k}}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$. Thus,
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,Q_{k})}.$
Case $(2)$. If there exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
Then there is some $z\in\mathcal{I}_{(k,j)}$ such that $p\in\overline{B}_{M_{(k,j)}}(z,\frac{\epsilon_{(k,j)}}{2})$, and
$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{c_{k,j}}{n}D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
$+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)}).$
The similar argument as in Case $(1)$ yields that
$D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,j)}.$
Case $(3)$. If there exists $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
One has
$D($ $\mathcal{E}_{n}(p),\mu_{(k,1)})\leq\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}D(\mathcal{E}_{M_{(k-1,Q_{k-1})^{\ast}}}(p),\mu_{(k,1)})$
$+\frac{c_{k-1,Q_{k-1}}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})+\frac{c_{k,0}}{n}D(\mathcal{E}_{c_{k,0}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
$+\frac{n_{(k,1)}+m_{(k,1)}}{n}\sum_{l=0}^{i-1}D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})$
$+\frac{n-M_{(k,1)}^{i}}{n}D(\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),\mu_{(k,1)}).$
Note that
$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
$D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})\leq\frac{5}{k}+2\delta_{(k,1)},$
and
$D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ $\leq D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})$
$\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
This gives us
$D(\mathcal{E}_{n}(p),K)\leq\frac{16}{k-1}+4\delta_{(k-1,0)}.$
Case $(4)$. If there exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
In this case, Using the expression
$\mathcal{E}_{n}(z)=$ $\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}}{n}\mathcal{E}_{c_{k,j-1}}(f^{M_{(k,j-1)^{\ast}}}(p))$
$+$ $\frac{n_{(k,j)}+m_{(k,j)}}{n}\sum_{l=0}^{i-1}\mathcal{E}_{n_{(k,j)}+m_{(k,j)}}(f^{M_{(k,j)}^{l}}(p))+\frac{n-M_{(k,1)}^{i}}{n}\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),$
and repeating the similar argument as above, one can obtain
$D(\mathcal{E}_{n}(p),K)\leq\frac{15}{k}+4\delta_{(k,j)}.$
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on above arguments, one can also show sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
Step $3$. Construction of measure.
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying that the $\rho$-measure of some appropriate dynamical balls decay exponentially fast along sequence $\{M_{k}\}$, where $M_{k}=M_{k,Q_{k}}$. Then we can use the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
Define
$\rho_{k}=\frac{1}{\#\mathcal{I}_{(k,Q_{k})}}\sum_{x\in\mathcal{I}_{(k,Q_{k})}}\delta_{x}.$
For simplicity, we write $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ instead of $\mathcal{I}_{k}$, $\mathbb{F}_{k}$, respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$, which implies that $\rho(\mathbb{F})=1$. By (3.9), we have
#$\#\mathcal{I}_{k}\geq\exp\left[M_{k}\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
Lemma 3.5.
Let $k_{0}\geq 1$, and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball which intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
$\rho_{k}\left(B_{M_{k_{0}}}\left(z,\frac{\epsilon^{\ast}}{2}\right)\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}.$
Proof.
If $k=k_{0}+1$, it is easy to check that
#$\#(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
and hence
$\rho_{k_{0}+1}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})}}{\#\mathcal{I}_{k_{0}+1}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
In general, for any $k>k_{0}$,
$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
For large enough $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
$\rho\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)$ $\leq\liminf_{k\to\infty}\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}$
$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
Using Lemma 2.2, one has
$(1-\zeta)^{2}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
The arbitrariness of $\zeta$ implies that
$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
Now dividing $|\log\epsilon^{\ast}|$ in both sides and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d).$ (3.12)
Finally, by (3.1) and (3.12) we get
$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
3.2 Bowen metric mean dimension of saturated sets
We prove Theorems 1.2 and 1.3 in this subsection.
3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already given in [Y23, Lemma 3.3,(3)]
Lemma 3.6.
Let $(X,f)$ be a t.d.s.. Let $K\subset M_{f}(X)$ be a nonempty closed connected subset. Then for any $\epsilon>0$, we have
$h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
Thus, by the Lemma 3.6 we obtain an upper bound of the Bowen metric mean dimension of saturated set:
$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.13)
3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof of lower bound is similar to that of in subsection 3.1.2. So we only give a sketch for the estimation of lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
We shall show that for any small enough $\epsilon^{\ast}>0$, it holds that
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$ (3.14)
Fix $\zeta>0$ and pick a sequence $\{\mu_{(k,i)}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that
$K\subset\cup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})<\frac{1}{k},\ D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})<\frac{1}{k}.$
Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow $\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$ with lexicographic order again. Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
Assume that $H:=\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Then one has
$\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu_{(k,i)}}(f,\xi)>H-\zeta,\ \forall(k,i)\in\mathcal{A}^{\prime}.$
Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination of ergodic $f$-invariant probability measures with rational coefficients such that
$D(\mu_{(k,i)},\nu_{(k,i)})<\delta_{(k,i)},\ \ \inf_{\xi\succ\mathcal{U}}h_{\mu_{(k,i)}}(f,\xi)\leq\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)+\zeta.$
Applying Katok’s entropy formula and repeating the similar argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that
$(1)$ for any $1\leq l\leq p_{k,i}$, $\lambda_{k,i,l}n_{(k,i)}^{\prime}$ is an integer and
$\frac{n_{(k,i)}^{\prime}}{n_{(k,i)}+m_{(k,i)}}>\max\{1-\zeta,1-\frac{1}{k}\},$
where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$.
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$ which are $(n_{(k,i)}^{\prime},\epsilon^{\ast})$-separated and
#$\#E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
$(3)$ for any $x\in E_{k,i,l}$,
$D(\mathcal{E}_{\lambda_{k,i,l}n_{(k,i)}^{\prime}}(x),\nu_{l}^{k,i})<\delta_{(k,i)}.$
Let
$\mathcal{C}_{(k,i)}=\left\{y=y(x_{1},\ldots,x_{p_{k,i}}):(x_{1},\ldots,x_{p_{k,i}})\in E_{k,i,1}\times\cdots\times E_{k,i,p_{k,i}}\right\},$
where $y(x_{1},\ldots,x_{p_{k,i}})$ $\epsilon_{(k,i)}$-shadows the pieces of orbits
[…]
Then $\mathcal{C}_{(k,i)}$ is $(n_{(k,i)},4\epsilon^{\ast})$-separated with
#$\#\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
For $k\geq 0$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $N_{(k,0)}$ be defined as in subsection 3.1.2. Choose a strictly increasing sequence $\{N_{(k,i)}\}_{(k,i)\in\mathcal{A}^{\prime}}$ such that $N_{(1,1)}=1$ and
$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists $z=z(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})$ $\epsilon_{(k,i)}$-shadows the pieces of orbits
$\{y_{k,i,l},f(y_{k,i,l}),\ldots,f^{n_{(k,i)}-1}(y_{k,i,l})\},\ l=1,\cdots,N_{(k,i)},$
with gap $m_{(k,i)}$. Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points, which is $(c_{k,i},3\epsilon^{\ast})$-separated, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$. Suppose that we have already constructed set $\mathcal{I}_{(k,j)^{\ast}}$. Let $M_{(k,j)}=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j}$. For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, let $z=z(x,y)$ be the shadowing point such that $d_{M_{(k,j)^{\ast}}}(z,x)<\epsilon_{(k,j)}$ and $d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}+m_{(k,j)}}}(z))<\epsilon_{(k,j)}$. Then set $\mathcal{I}_{(k,j)}=\{z=z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$, and $\mathbb{F}=\cap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}$, where $\mathbb{F}_{(k,j)}=\cup_{z\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)})$. Then one has $\mathbb{F}\subset G_{K}^{\omega,U}$. Define a sequence of measures $\{\rho_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ by
$\rho_{(k,j)}=\frac{1}{\#\mathcal{I}_{(k,j)}}\sum_{x\in\mathcal{I}_{(k,j)}}\delta_{x},$
and assuming $\rho_{(k,j)}\to\rho$. One can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with $\mathbb{F}$, it holds
$\rho(B_{n}(z,\frac{\epsilon^{\ast}}{2}))\leq\exp\left[-n(1-\zeta)^{2}(H-2\zeta)\right].$
Then according to Lemma 2.1, we have
$(1-\zeta)^{2}(H-2\zeta)\leq h_{top}^{B}(f,\mathbb{F},\frac{\epsilon^{\ast}}{2})\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$
Letting $\zeta\to 0$ gives us
$\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2})$
Therefore, by (3.13) and (3.14) we get
$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$ (3.15)
3.2.3 Proof of Theorem 1.3
Firstly, we show
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}.$ (3.16)
Fix $\mu\in K$ and $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$. Then, by Lemma 3.6, for every $\epsilon>0$,
$h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
This shows
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)$
and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\underline{\rm mrid}_{\mu}(f,K,d)$ for every $\mu\in K$.
On the other hand, we may assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove. Let $\gamma>0$. Then according to Theorem 1.2 we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for any $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Assume that $\mu_{k}\rightarrow\mu$. Then $\mu\in M_{f}(X)$ since $K$ is closed. This shows
$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
Letting $\gamma\to 0$, we have
$\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
This completes the proof of (3.16).
Using formula (3.2), the inequality
$\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d).$
is clear. On the other hand, let $\gamma>0$ and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}$. Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for sufficiently large $k$,
$M_{\gamma}<\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi).$
Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows
$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
Letting $\gamma\to 0$, we have
$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
3.3 Upper capacity metric mean dimension of saturated sets
We prove the Theorem 1.4 in this subsection.
According to Theorem 2.4, the proof of Theorem 1.4 will be done, once we prove that for $\epsilon^{\ast}>0$, $\zeta>0$ small enough and $\delta\in(0,1)$, it holds
$h_{top}(f,G_{K}^{\omega,U},\epsilon^{\ast})>\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
That is, we must pick out enough separated points. Take $\mu^{\ast}\in M_{f}^{e}(X)$ with
$\underline{h}_{\mu^{\ast}}^{K}(f,3\epsilon^{\ast},\delta)>H-\zeta,$
where $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)>0$. Then for large enough $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
#$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
Now fix such $\mathcal{N}$, and write $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
Since $K$ is a connected closed subset, a standard argument shows that there exists sequence $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
$K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{k,i},\frac{1}{k}),\ D(\mu_{k,i},\mu_{k,i+1})<\frac{1}{k},\ \text{and }D(\mu_{k,Q_{k}},\mu_{k+1,1})<\frac{1}{k}.$
Consider the set $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq k\leq Q_{k}\}$, $\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$ with lexicographic order. The sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed similar as in subsection 3.1.2.
For nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$. Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$. For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be same as in subsection 3.1.2. For $(k,i)\in\mathcal{A}^{\prime}$, since $G_{\mu_{k,i}}\neq\emptyset$, one can pick $x_{k,i}\in X$ and $n_{(k,i)}\in\mathbb{N}$ such that
$D(\mathcal{E}_{n_{(k,i)}}(x),\mu_{k,i})<\delta_{(k,i)},\ \frac{n_{(k,i)}}{n_{(k,i)}+m_{(k,i)}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\}.$
Set $\mathcal{C}_{(k,i)}=\{x_{k,i}\}$.
Now choose a strictly increasing sequence of integers $\{N_{(i,j)}\}_{(k,j)\in\mathcal{A}\ \text{with }j\neq 0}$ with $N_{(1,1)}=1$ and
$\begin{split}n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right];\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
Based on $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar with previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ and the fractal set
$\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}\subset G_{K}^{\omega,U}.$
It is worth mentioning that for any $(k,j)>(0,0)$, #$\#\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
$d_{M_{(0,0)}}(x,x^{\prime})>2\epsilon^{\ast}.$
Recall that $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$. Hence for $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$, and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$ one has
$d_{M_{(0,0)}}(y,y^{\prime})>d_{M_{(0,0)}}(x,x^{\prime})-d_{M_{(0,0)}}(x,y)-d_{M_{(0,0)}}(x^{\prime},y^{\prime})>\epsilon^{\ast}.$
Hence for any large enough $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus
$h_{top}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
4 Applications
In this section, we apply the main results to multifractal analysis and mean Li-Yorke chaos. The main results are Theorems 4.1 and 4.3.
4.1 Variational principle of level sets
In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ in the sense of an abstract framework.
Let $(X,f)$ be a dynamical system, and let $\varphi:M_{f}(X)\to\mathbb{R}$ be a continuous function. Let
$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right],\$
For $\alpha\in L$, let $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$, and let
$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\}$
denote the level set of $\alpha$ w.r.t. $\varphi$ consisting of the set of points of $X$ such that $\varphi$ is taking constant value $\alpha$ on $V_{f}(x)$. Then continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
Condition A: For each given $\varphi$ and $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
Some examples of $\varphi$ satisfying Condition A are given as follows:
  • (a)
    If $\varphi=\alpha\in\mathbb{R}$ is a constant, then $L_{\varphi}=\{\alpha\}$, $K_{\alpha}=X$, $M_{f}(\varphi,\alpha)=M_{f}(X)$;
  • (b)
    Let $\psi$, $\phi\in C(X)$ such that $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ holds for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$). We define
    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu}.$
    Then for any $\alpha\in L_{\varphi}$,
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\},$
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\};$
  • (c)
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i,e. for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
    $\varphi:M_{f}(X)\to\mathbb{R},~{}\mu\mapsto\lim_{n\to\infty}\int\frac{\phi_{n}(x)}{n}\mathop{}\!\mathrm{d}\mu(x).$
    Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for any $\Phi$ and $\alpha\in L_{\Phi}$.
    Specially, when sequence $\log\phi_{n}(x)=S_{n}\psi(x)$ is exactly an additive function, where $\psi\in C(X)$ and $S_{n}\psi(x)=\sum_{i=0}^{n-1}\psi(f^{i}(x))$, in this case we have
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\varphi\mathop{}\!\mathrm{d}\mu=\alpha\},$
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\varphi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Given the nonempty open subset $U\subset X$, we put
$K_{\alpha}^{\omega,U}=K_{\alpha}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}.$
Theorem 4.1.
Let $(X,d,f)$ be a dynamical system with specification property and $U\subset X$ be a nonempty open subset. Suppose that let $\varphi:M_{f}(X)\to\mathbb{R}$ be a continuous function satisfying Condition A. For any $\alpha\in L$ and $S\in\{B,P\}$, we have
$\overline{\rm mdim}_{M}^{S}(f,K_{\alpha}^{\omega,U},d)=$ $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha},d)$
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
[…]
Proof.
Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$ which is not dependent on $\mu$, such that
$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$ (4.1)
holds for all $\mu\in M_{f}(X)$ and any small enough $\epsilon>0$. Together with the fact $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha}^{\omega,U},\epsilon)$
for all sufficiently small $\epsilon>0$, which implies that
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha}^{\omega,U},d).$ (4.2)
Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$. Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, then $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$. By Lemma 3.2 $(1)$, we get
$\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (4.3)
By (4.2), (4.3), and the fact stated in Proposition 2.1 $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)$ for any $Z\subset X$, we complete the proof. ∎
A direct consequence of Theorem 4.1 is the variational principles for Bowen metric mean dimension of level sets obtained by Backes and Rodrigues [BR23, Theorem A], but we can strength the result which states that the variational principle is still valid for packing metric mean dimension of level sets.
Corollary 4.2.
Let $(X,d,f)$ be a dynamical system with specification property and $\psi$ be a continuous function on $X$. Let
$E_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}$
denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then we have
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}\xi<\epsilon}h_{\mu}(f,\xi),$
[…]
Proof.
[…]
$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}$
satisfies Condition A. Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We get the desired result by Theorem 4.1. ∎
Readers can turn to [O03, Tho09, ZC13] and references therein for the multifractal analysis of level sets using topological pressure. Then by Theorem 4.1 one can get analogous results for the metric mean dimensions of level sets of some special functions determined by the dynamical behaviors of the system.
4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In this subsection, we link the concepts of metric mean dimension and mean Li-Yorke chaos in the sense of infinite entropy. We give an quantitative characterization for the chaotic phenomenon of infinite entropy systems. Namely, we show the packing metric mean dimension of the set of mean Li-Yorke pairs is bounded from blow by the metric mean dimension of the phase space (Theorem 4.3).
Let $(X,d,f)$ be a dynamical system. This naturally gives a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is compact endowed with the product metric $d\times d$ defined by $d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\}$, and the map
$f\times f:X\times X\rightarrow X\times X$
given by $(f\times f)(x,y):=(fx,fy)$ is continuous.
It is well-known that if two systems admit the specification property, then the product system also has the specification property(cf. [DGS76, Proposition 21.4]). A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
The set of mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$. A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair. We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set.
Theorem 4.3.
Let $(X,d,f)$ be a dynamical system with specification property with positive topological entropy. Then we have
$\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
Proof.
Claim: there exists $\mu_{0}\in M_{f\times f}(X\times X)$ with $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
Proof of the claim.
By the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$. We put $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then we can choose an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfying the desired claim. ∎
Now take $\nu\in M_{f}(X)$ and set $\mu_{1}\in M_{f\times f}(X\times X)$ defined by $\mu_{1}=\phi_{\ast}\nu$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. Then one has $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. Assume that $\overline{\rm mdim}_{M}(f,X,d)$ is finite. For any $\gamma>0$, by Theorem 2.4 we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $\rho_{k}\in M_{f}(X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then it is easy to check that for every $k$,
$\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)=\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta).$ (4.4)
Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ satisfying ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
$H_{\rho_{k}^{\ast}}(\eta^{n})$ $=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=-\sum_{C\in\xi^{n}}\rho_{k}^{\ast}(C\times C)\log\rho_{k}^{\ast}(C\times C)$
$=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=H_{\rho_{k}}(\xi^{n}),$
which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\leq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. Similarly, if $\eta$ is a finite Borel partition of $X\times X$, then $\xi=\{\phi^{-1}(D\cap\Delta):D\in\eta\}$ is a finite partition of $\Delta=\{(x,x):x\in X\}$, such that ${\rm diam}(\xi)\leq{\rm diam}(\eta)$. Moreover, one has
$H_{\rho_{k}}(\xi^{n})=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=H_{\rho_{k}^{\ast}}(\eta^{n}),$
which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\geq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. This shows (4.4).
Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e., the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. For any $(x_{0},y_{0})\in G_{K}$, there are sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$. Hence
$\lim_{n\to\infty}\frac{1}{t_{n}^{0}}\sum_{i=0}^{t_{n}^{0}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}>0,$
and
$\lim_{n\to\infty}\frac{1}{t_{n}^{1}}\sum_{i=0}^{t_{n}^{1}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0.$
This shows $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
$\overline{\rm mdim}_{M}^{P}$ $(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}^{P}(f\times f,G_{K},d\times d)$
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}\geq\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
Since $\gamma$ is arbitrary, one has
$\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
The case of $\overline{\rm mdim}_{M}(f,d)=\infty$ is still valid by considering a slight modification. ∎
5 Examples
Applying the main results, we give the specific formula of the metric mean dimensions of generic points of full shift over compact metric space.
Example 5.1.
Endow the product space $X=[0,1]^{\mathbb{Z}}$ with product metric
$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
and let $\sigma:X\to X$ be the left shift map defined by $(x_{n})_{n\in\mathbb{Z}}\mapsto(x_{n+1})_{n\in\mathbb{Z}}$. It is shown that the system $(X,d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm\overline{mdim}}_{M}(\sigma,X,d)=1$ [LT18]. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is the product measure, which is $\sigma$-ergodic. Then the ergodicity enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in X.$
Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ24, Theorem 1.1], we have
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,X,d).$
Then by Corollary 1.5, we have
${\rm\overline{mdim}}_{M}^{B}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{P}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{UC}(\sigma,G_{\mu},d)=1.$
We extend the Example 5.1 to more general full shift.
Example 5.2.
Let $K$ be a compact metric space with metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
$d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
For $n\geq 0$, let $\pi_{n}$ be the projection from $K^{\mathbb{Z}}$ to its first $n$ coordinates.
It was proved in [S22, Theorem 15]
$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K),$
where $\overline{\rm dim}_{B}(F)$ is the upper box dimension of $F$ under the metric $d^{q}$.
Theorem 5.3.
Let $K$ be a compact metric space with metric $d$. There exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(F)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(\sigma,\xi).$
Proof.
It is known that the system $(K^{\mathbb{Z}},\sigma)$ has specification property. Let $S_{\epsilon}$ be a maximal $\epsilon$-separated subset of $K$ with respect to metric $d$. For $n\geq 1$, we put
$\hat{\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
and consider the set $\mathcal{S}_{n}\subset X$ consisting of one preimage of each element of $\hat{\mathcal{S}_{n}}$ under $\pi_{n}$. Then $\hat{\mathcal{S}}_{n}$ is an $(n,\epsilon)$-separated set of $K^{\mathbb{Z}}$ and #$\#{\hat{\mathcal{S}}_{n}}=(\#S_{\epsilon})^{n}$. Define
$\mu_{n}:=\frac{1}{\#\hat{\mathcal{S}}_{n}}\sum_{x\in\hat{\mathcal{S}}_{n}}\mathcal{E}_{n}(x).$
We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(X)$. Then by Lemma 3.1, for each $\epsilon>0$ we have
$\log\#S_{\epsilon}=\limsup_{n\to\infty}\frac{1}{n}\log\#\hat{\mathcal{S}}_{n}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(\sigma,\xi).$
Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$. Combining with Theorem 1.3, one has
$\overline{\rm dim}_{B}(K)=\limsup_{\epsilon\to 0}\frac{\log\#S_{\epsilon}}{|\log\epsilon|}\leq\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm dim}_{B}(K).$
Finally, by Corollary 1.5 we complete the proof.
Availability of data and materials: Not applicable.
Competing interests: The authors declare that they have no competing interests.
Acknowledgments: The first and second authors were supported by NNSF of China (12101340). The third author was supported by the China Postdoctoral Science Foundation (Grant Number: 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University.

### v2

@@
 Abstract.
-Saturated sets and its reduced case the set of generic points are two types of important fractal-like sets in Multifractal analysis of dynamical systems.
-In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set generic points in both topological and measure-theoretic situations.
+Saturated set and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems.
+In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set of generic points in both topological and measure-theoretic perspectives.
 For systems with specification property, we establish the certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets have full metric mean dimension.
@@
 As applications, we further exhibit some qualitative aspects of metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
-Keywords and phrases: Metric mean dimension; Saturated sets; Generic point; Variational principle; Multifractal analysis
+Keywords and phrases: Metric mean dimension; Saturated set; Generic point; Variational principle; Multifractal analysis
 1 Introduction
 Let $(X,d,f)$ be a topological dynamical system (t.d.s.
-for short), where $(X,d)$ is a compact metric space with metric $d$, and $f:X\to X$ is a homeomorphism.
+for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism.
 Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(x)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively.
@@
 The pioneering work directly contributes to the development of dimension theory of dynamical systems, especially as a powerful tool to study the “size” of fractal-like sets in multifractal analysis.
-The multifractal analysis, which can date back to Besicovitch, considers the decomposition of the whole phase space into smaller subset consisting of the points with the similar dynamical behaviors, or known as multifractal spectrum, and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies.
-Such a procedure allows us to get more information about the dynamics and even recover the partial dynamical behaviors of the system, which we call such phenomenon multifractal rigidity.
+The multifractal analysis, which can date back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of the points with the similar dynamical behaviors(or known as multifractal spectrum), and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies.
+Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, which is called phenomenon multifractal rigidity.
 We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
-Let us state some backgrounds about the dimensional entropies of saturated sets.
+Let us mention some relevant results about the dimensional entropies of saturated sets.
 Given $x\in X$, by $\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X)$, where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of the Birkhoff average w.r.t.
@@
 The set of all generic points of $\mu$ is denoted by $G_{\mu}$.
-Then the Birkhoff ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
-Bowen [Bow73] showed for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
+Then the Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
+Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
 $h_{top}^{B}(f,G_{\mu})=h_{\mu}(f).$
 For non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$ and even easily constructs the example showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$ such that Bowen’s result is not valid.
-It is natural to expect whether there are some proper conditions such that the analogous equality still holds for non-ergodic measures.
+It is natural to expect whether there are some proper conditions such that the equality still holds for non-ergodic measures.
 A subset $D$ of $X$ is called a saturated set if $x\in D$ and $V_{f}(x)=V_{f}(y)$, then $y\in D$.
-More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, this defines a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$.
-For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems [LST13, TV17].
-As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and uniform separation property (e.g.
-expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any compact connected non-empty set $K\subset M_{f}(X)$,
+More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we can obtain a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$.
+For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17].
+As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and with uniform separation property (e.g.
+expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
 $h_{top}^{B}(T,G_{K})=\inf\{h_{\mu}(f):\mu\in K\}.$
@@
 $h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$
-holds for any $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to any non-ergodic measures.
+holds for all $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to non-ergodic measures.
 Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional entropies of saturated sets.
-Packing entropy of non-compact subset was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension.
-Under a stronger condition of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] further established the following variational principle for any compact connected non-empty set $K\subset M_{f}(X)$,
+Packing entropy of non-compact subsets was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension.
+Additionally, under the (stronger) conditions of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] showed the following variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
 $h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\},$
 where $h_{top}^{P}(f,K)$ is the packing entropy of $K$.
-Later, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
+Afterwards, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
 $h_{top}^{UC}(f,G_{K}^{C})=h_{top}(f),$
 where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$.
-Besides, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle still hold for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points.
-Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results about the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
+Furthermore, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle is valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points.
+Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in the setting of nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results concerning the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
 According to the possible values of topological entropy, dynamical systems can be divided into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems.
@@
 See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b].
-The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, dues to Gutman and Śpiewak[GS21].
+The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21].
 More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
@@
 Some variational principles of metric mean dimension of $\mathbb{Z}$-actions can be found in [LW00, GS21, S22, YCZ23a].
-Then the aforementioned work suggests the following questions for infinite entropy systems:
+Then the aforementioned work suggests the following questions for saturated sets in infinite entropy systems:
 Questions:
@@
 If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
-For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the setting of infinite entropy systems encounters some nontrivial challenges.
-Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ containing in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
+For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges.
+Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ contained in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
 $\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-However, for systems with specification property we can give a full answer to question (2).
+Prior to the comparison of different types of specification-like property, we only focus on the systems with specification property and give a full answer to question (2).
 Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the following statements of main results involve more general (mixed) saturated sets.
@@
 If $G_{K}^{\omega,U}\not=\emptyset$, then there exists a transitive point in the open set $U$ such that $V_{f}(x)=K$.
-Different types of metric mean dimensions of saturated sets exhibit different dynamical behaviors.
-There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated sets has full metric mean dimension (Theorem 1.4).
+Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions.
+There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated set has full metric mean dimension (Theorem 1.4).
 Theorem 1.1.
@@
 $\inf_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d).$
-Inspired the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures.
+Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures.
 For any closed set $K$ of $M_{f}(X)$ and $\mu\in K$, by $M_{K}(\mu)$ we denote the space of sequences of Borel probability measures in $K$ which converge to $\mu$ in the weak-topology.
-We respectively define the lower and upper mean Rényi information dimensions of $\mu\in K$ as
+We respectively define the (modified) lower and upper mean Rényi information dimensions of $\mu\in K$ as
 $\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi),$
@@
 $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
-We remark that Theorem 1.3 does not need the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
+We remark that Theorem 1.3 does not require the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
 Theorem 1.4.
@@
 where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$ defined by separated sets.
-As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being just a single point.
+As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being exactly a single point.
 The Corollary 1.5 presents the metric mean dimensions of generic points of any invariant measures for systems with specification property, which extends the result given in [YCZ24, Theorem 1.1] to all non-ergodic measures.
@@
 $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
-The “Consequently” part can realize for some dynamical systems.
+The “Consequently” part can be realized for some dynamical systems.
 For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$.
@@
 See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms.
-For some certain systems, by Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
+By Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
 Corollary 1.6.
@@
   • (1)
-    There are plenty of infinite entropy systems with specification property, for instance, the usual the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
+    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
   • (2)
@@
   • (3)
-    We believe it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs.
+    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs.
 The novelty of our work not only presents a proper and full extension of the previous work in [PS07, ZCC12, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for the multifractal analysis of other types of fractal-like sets in infinite entropy systems.
@@
 In section 3, we prove the main results.
-In Section 4, we give two applications of main results.
+In section 4, we give two applications of the main results.
 In section 5, we show the metric mean dimensions of generic points of full shift over compact metric spaces.
 2 Preliminaries
-In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and the precise definition of specification property in dynamical systems.
+In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of specification property in dynamical systems.
 2.1 Three types of metric mean dimensions
 In this subsection, we recall the definitions of upper capacity metric mean dimension on subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension on subsets [YCZ22].
-Let $(X,d,f)$ be a dynamical system. For $n\in N$, we define the Bowen metric $d_{n}$ on $X$ as
+Let $(X,d,f)$ be a dynamical system. For $n\in\mathbb{N}$, we define the Bowen metric $d_{n}$ on $X$ as
 $d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\}$
@@
 The $\epsilon$-upper capacity topological entropy of $Z$ is given by
-$h_{top}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
-Analogous to the definition of box dimension in fractal geometry, we can formulate the definition of upper capacity metric mean dimension of subsets using separated sets.
+$h_{top}^{UC}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
+Analogous to the definition of box dimension in fractal geometry, we can formulate the definition for upper capacity metric mean dimension of subsets using separated sets.
 Definition 2.1.
@@
 $\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$ (2.1)
-When $Z=X$, we sometimes write $h_{top}(f,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, $\overline{\rm mdim}_{M}^{UC}(f,X,d)=\overline{\rm mdim}_{M}(f,X,d)$ respectively, then in this case it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00].
-Such a formulation allows us to extend the metric mean dimensions to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
-Recall that the topological entropy of $f$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
-$h_{top}(f)=\lim_{\epsilon\to 0}h_{top}(f,\epsilon)=\sup_{\epsilon>0}h_{top}(f,\epsilon).$
-For any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity.
-Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite entropy of the system as $\epsilon$ tends to 00.
+When $Z=X$, we write $h_{top}(f,X,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)=\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively.
+In this case, it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00].
+Such a formulation allows us to extend the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
+Recall that the upper capacity topological entropy of $Z$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
+$h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
+Especially, for any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity.
+Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite (upper capacity) topological entropy as $\epsilon$ tends to 00.
 More precisely, for sufficiently $\epsilon>0$, we may think of
 $h_{top}^{UC}(f,Z,\epsilon)\approx\overline{\rm mdim}_{M}^{UC}(f,Z,d)\cdot|\log\epsilon|.$
-It is easy to see that any finite entropy system has zero metric mean dimension. This further verifies that metric mean dimension is a good candidate to capture the dynamics of infinite systems.
+It is easy to see that any finite entropy system has zero metric mean dimension, and hence verifies that metric mean dimension is a good candidate to capture the dynamics of infinite entropy systems.
 Similar to the definitions of Hausdorff dimension and packing dimension in fractal geometry, as an analogue of the classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which is defined by Carathéodory-Pesin structures, the corresponding dimensional characterizations of metric mean dimensions via Carathéodory-Pesin structures are Bowen and packing metric mean dimensions, which has been introduced in [LV21, YCZ22, W21] and studied by several authors in [BR23, CLS21, LL24, FKO24, YCZ24].
@@
 where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}$ such that $Z\subset\cup_{i}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
-The quantity $M(Z,s,N,\epsilon)$ does not decrease as $N$ increase. So we set
+The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So we set
 $M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon).$
@@
 where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i$.
-Since the quantity $P(Z,s,N,\epsilon)$ does not increase as $N$ increases, the following limit exists:
+Since the quantity $P(Z,s,N,\epsilon)$ is non-increasing as $N$ increases, and thus the following limit exists:
 $P(Z,s,\epsilon)=\lim_{N\to\infty}P(Z,s,N,\epsilon).$
@@
 $\overline{\rm mdim}_{M}^{P}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,Z,\epsilon)}{|\log\epsilon|}.$
-One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to defined the corresponding three types of lower metric mean dimensions on subsets.
-We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6] as follows:
+One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to obtain the corresponding three types of lower metric mean dimensions on subsets, which are denoted by $\underline{\rm mdim}_{M}^{S}(f,Z,d)$ with $S\in\{UC,B,P\}$.
+If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$.
+We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6].
 Proposition 2.1.
@@
 $h_{top}^{P}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{P}(f,Z_{i},\epsilon),$
-and $h_{top}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
+and $h_{top}^{UC}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}^{UC}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
 $(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then
@@
 $\overline{\rm mdim}^{P}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{P}_{M}(f,Z_{i},d).$
-$(3)$ If $Z$ is an any non-empty subset of $X$, then for every $10<\epsilon<$ 1, we have
-$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}(f,Z,\epsilon).$
+$(3)$ If $Z$ is a nonempty subset of $X$, then for every $10<\epsilon<$ 1, we have
+$h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}^{UC}(f,Z,\epsilon).$
 Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$; if $Z$ is $f$-invariant and compact, then the equalities hold.
-Observe that the metric mean dimension of a given subset is in general hard to calculate its precise value.
-However, the Entropy Distribution Principle allows us to get the lower bounds of the $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems.
-Consequently, the powerful tool enables us to get the lower bounds of entropy and metric mean dimension of the fractal-like sets.
+The metric mean dimension of a given subset is in general not easy to calculate its precise value.
+However, the Entropy Distribution Principle allows us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems.
+Consequently, the powerful tool gives the lower bounds of entropy and metric mean dimension of the fractal-like sets.
 Lemma 2.1.
-[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exists a Borel probability measure $\mu$ and a constant $C>0$ satisfying:
+[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$ and a constant $C>0$ satisfying:
 $(1)$ for every $n$, $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for each Bowen ball $B_{n}(x,\epsilon)$ such that $B_{n}(x,\epsilon)\cap Z\neq\emptyset$;
@@
 Let $Z$ be a Borel set of $X$ and $s\geq 0$.
-If there exists a Borel probability measure $\mu$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
+If there exist a Borel probability measure $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
 $(1)$ $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$;
@@
 For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$.
-Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint sub-collection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that
+Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that
 $E\subset\bigcup_{j}B_{n_{i}}(x_{j},\epsilon).$
@@
 $D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$
-where $\{\phi_{k}\}$ is a countable dense subset of $C(X)$ taking values in $[0,1]$.
+where $\{\phi_{k}\}$ is a family of countable dense subset of $C(X)$ taking values in $[0,1]$.
 Thus, the diameter of $M(X)$ in sense of metric $D$ is always less than $1$, and for fixed $\mu$ and $\nu$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$.
@@
 Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in sense of weak-topology.
-Given an invariant measure $\mu\in M_{f}(X)$, we say a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$.
+Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$.
 The partition entropy of $\xi$ w.r.t.
@@
     The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu$ is given by
-    $\inf_{{\rm diam}\xi<\epsilon}h_{\mu}(f,\xi).$
+    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
   • (b)
@@
 Theorem 2.4.
-(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with metric $d$. Then for any $\delta\in(0,1)$,
+(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with a metric $d$. Then for any $\delta\in(0,1)$,
 $\overline{\rm mdim}_{M}(f,X,d)$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(X)}F(\mu,\epsilon),$
@@
 Definition 2.4 (Specification property).
-One says that $f$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
+One says that a t.d.s.
+$(X,f)$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
 $d(f^{p+a_{j}}x,f^{p}x_{j})<\epsilon$
@@
 We shall construct a Moran fractal set containing in $G_{K}^{\omega,U}$, which is a limit set with geometric construction.
-Then such set is suitable for the purpose of the computation of metric mean dimensions of saturated sets.
+Then such set is suitable for the purpose of the computation of the lower bound of metric mean dimensions of saturated sets.
 3.1 Packing metric mean dimension of saturated sets
-We prove Theorem 1.1. To arrive this goal, we divide the proof ot Theorem 1.1 into two parts.
-3.1.1 Upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$
-We adapt a general method presented in [PS07, Theorem 4.1] [ZCC12, Theorem 1.1 $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
+We first prove Theorem 1.1. To arrive this goal, we divide the proof of the Theorem 1.1 into two parts.
+3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
+We adapt a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
 Lemma 3.1.
 [Y23, Lemma 3.1] Let $(X,f)$ be a dynamical system and $\epsilon>0$.
-Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$ and $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality.
+Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality.
 We define
@@
 Let $(X,f)$ be a t.d.s..
-$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$, we have
-$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$
+$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$,
+$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 $(2)$ If $\mu\in M_{f}(X)$, then for any $\epsilon>0$,
@@
 For $n\in\mathbb{N}$, $\delta>0$, we put
-$\mathcal{R}(K,\delta,n):=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
-and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{R}(K,\delta,n)$.
-Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}(K,\delta,n)$ holds for every $\delta$, it follows from Proposition 2.1 that
-$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}(K,\delta,n),\epsilon)\leq\sup_{k}h_{top}(f,\mathcal{R}(K,\delta,n),\epsilon).$
+$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)~{}\text{for infinitely many }n>k\right\},$
+$\mathcal{Q}_{K}(\delta,n):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
+and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$.
+Fix $x\in G^{K}$.
+Then, we can choose a subsequence $\{n_{i}\}$ such that $\mathcal{E}_{n_{i}}(x)\to\mu$ for some $\mu\in K$.
+This implies for every $\delta>0$, there exists a positive integer $k$ such that $\mathcal{E}_{n_{i}}(x)\in B(K,\delta)$ for sufficiently large $n_{i}$ greater than $k$.
+Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ holds for every $\delta$, it follows from Proposition 2.1 that
+$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon),$
+which implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon).$
 Next we shall prove for any $\gamma>0$, there is $\mu\in K$ such that
@@
 $S(K,\delta_{k},n_{k},\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
-Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}(K,\delta_{k},n_{k})$, and set
-$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
+Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$, and set
+$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x)\in B(K,\delta_{k}).$
 Choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed convex set. By Lemma 3.1, we have
-$\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\geq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\geq\Theta(K,\epsilon)-\gamma.$
+$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 This completes the proof. ∎
-Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of saturated set:
+Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of $G_{K}$:
 $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,G_{K},\epsilon)}{|\log\epsilon|}\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.1)
@@
 $D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})$ $\leq\frac{1}{k},\ \mu_{(k,Q_{k})}=\mu^{\ast}.$
-Now set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order.
+Now the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order.
 Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$.
@@
 For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is a positive integer given in the definition of specification property.
-For any non-empty open set $U$, there exist $\tilde{\epsilon}>0$, a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$.
+For any nonempty open set $U$, there exist $\tilde{\epsilon}>0$ and a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$.
 One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$.
@@
 Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$.
-The following measure approximation derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
+The following measure approximation of $\mu^{*}$ derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
 Lemma 3.3.
@@
 $(2)$ there exists a subset $E_{k,i}$ which is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
-#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
+#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
 $(3)$ for every $x\in E_{k,i}$,
@@
 The last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and formulas (3.3), (3.5).
-For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$,
+For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$ taking values in $[0,1]$,
 $\left|\int\varphi\mathop{}\!\mathrm{d}\mathcal{E}_{n_{k}}(y)-\int\varphi\mathop{}\!\mathrm{d}\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y))\right)\right|$
@@
 Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$.
-Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how construct $\mathcal{I}_{(k,j)}$.
+Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$.
 Let
@@
 One has the fractal set $\mathbb{F}$ is contained in $\mathbb{F}\subset G_{K}^{\omega,U}$, that is,
-$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:X\subset\omega_{f}(x)\}.$
+$\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\}.$
 Proof.
@@
 For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$.
-Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there are some $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$.
-By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:X\subset\omega_{f}(x)\}$.
+Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there is a $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$.
+By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}$.
 For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, in order to estimate $D(\mathcal{E}_{n}(p),K)$, we split into four cases to discuss.
-Case $(1)$. If there exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
+Case $(1)$. There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
 Then there is some $z\in\mathcal{I}_{(k,Q_{k})}$ such that $p\in\overline{B}_{M_{(k,Q_{k})}}(z,\frac{\epsilon_{(k,Q_{k})}}{2})$, and
@@
 $D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,Q_{k})}.$
-Case $(2)$. If there exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
+Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
 Then there is some $z\in\mathcal{I}_{(k,j)}$ such that $p\in\overline{B}_{M_{(k,j)}}(z,\frac{\epsilon_{(k,j)}}{2})$, and
@@
 $D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,j)}.$
-Case $(3)$. If there exists $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
+Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
 One has
@@
 $D(\mathcal{E}_{n}(p),K)\leq\frac{16}{k-1}+4\delta_{(k-1,0)}.$
-Case $(4)$. If there exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
+Case $(4)$. There exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
 In this case, using the expression
@@
 $\rho_{k}=\frac{1}{\#\mathcal{I}_{(k,Q_{k})}}\sum_{x\in\mathcal{I}_{(k,Q_{k})}}\delta_{x}.$
-For simplicity, we write $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ instead of $\mathcal{I}_{k}$, $\mathbb{F}_{k}$, respectively.
+For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$, respectively.
 Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$.
@@
 We prove Theorems 1.2 and 1.3 in this subsection.
-3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
+3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
 The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already given in [Y23, Lemma 3.3,(3)]
@@
 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
-The proof of lower bound is similar to that of in subsection 3.1.2. So we only give a sketch for the estimation of lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
+The proof of the lower bound is similar to that of in subsection 3.1.2. So we only give a sketch for the estimation of lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
 We shall show that for any small enough $\epsilon^{\ast}>0$, it holds that
@@
 $K\subset\cup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})<\frac{1}{k},\ D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})<\frac{1}{k}.$
-Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow $\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$ with lexicographic order again.
+Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow
+$\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$
+with the lexicographic order again.
 Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
@@
 $\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu_{(k,i)}}(f,\xi)>H-\zeta,\ \forall(k,i)\in\mathcal{A}^{\prime}.$
-Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$.
+Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$.
 For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination of ergodic $f$-invariant probability measures with rational coefficients such that
 $D(\mu_{(k,i)},\nu_{(k,i)})<\delta_{(k,i)},\ \ \inf_{\xi\succ\mathcal{U}}h_{\mu_{(k,i)}}(f,\xi)\leq\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)+\zeta.$
-Applying Katok’s entropy formula and repeating the similar argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that
+Applying Katok’s $\epsilon$-entropy formula and repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
 $(1)$ for any $1\leq l\leq p_{k,i}$, $\lambda_{k,i,l}n_{(k,i)}^{\prime}$ is an integer and
@@
 where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$.
-$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$ which are $(n_{(k,i)}^{\prime},\epsilon^{\ast})$-separated and
+$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$ which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
 #$\#E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
@@
 $\mathcal{C}_{(k,i)}=\left\{y=y(x_{1},\ldots,x_{p_{k,i}}):(x_{1},\ldots,x_{p_{k,i}})\in E_{k,i,1}\times\cdots\times E_{k,i,p_{k,i}}\right\},$
-where $y(x_{1},\ldots,x_{p_{k,i}})$ $\epsilon_{(k,i)}$-shadows the pieces of orbits
+where $y(x_{1},\ldots,x_{p_{k,i}})$ is the point that $\epsilon_{(k,i)}$-shadows the pieces of orbits
 […]
-Then $\mathcal{C}_{(k,i)}$ is $(n_{(k,i)},4\epsilon^{\ast})$-separated with
+Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
 #$\#\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
@@
 Choose a strictly increasing sequence $\{N_{(k,i)}\}_{(k,i)\in\mathcal{A}^{\prime}}$ such that $N_{(1,1)}=1$ and
-$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
+$\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
 Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$.
-For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists $z=z(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})$ $\epsilon_{(k,i)}$-shadows the pieces of orbits
+For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
 $\{y_{k,i,l},f(y_{k,i,l}),\ldots,f^{n_{(k,i)}-1}(y_{k,i,l})\},\ l=1,\cdots,N_{(k,i)},$
-with gap $m_{(k,i)}$. Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points, which is $(c_{k,i},3\epsilon^{\ast})$-separated, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
+with the same gap $m_{(k,i)}$.
+Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points, which is $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
 Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$.
-Suppose that we have already constructed set $\mathcal{I}_{(k,j)^{\ast}}$.
+Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$.
 Let $M_{(k,j)}=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j}$.
-For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, let $z=z(x,y)$ be the shadowing point such that $d_{M_{(k,j)^{\ast}}}(z,x)<\epsilon_{(k,j)}$ and $d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}+m_{(k,j)}}}(z))<\epsilon_{(k,j)}$.
-Then set $\mathcal{I}_{(k,j)}=\{z=z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$, and $\mathbb{F}=\cap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}$, where $\mathbb{F}_{(k,j)}=\cup_{z\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)})$.
-Then one has $\mathbb{F}\subset G_{K}^{\omega,U}$.
-Define a sequence of measures $\{\rho_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ by
-$\rho_{(k,j)}=\frac{1}{\#\mathcal{I}_{(k,j)}}\sum_{x\in\mathcal{I}_{(k,j)}}\delta_{x},$
-and assuming $\rho_{(k,j)}\to\rho$. One can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with $\mathbb{F}$, it holds
+For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, let $z=z(x,y)$ be the shadowing point such that $d_{M_{(k,j)^{\ast}}}(z,x)<\epsilon_{(k,j)}$ and
+$d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}+m_{(k,j)}}}(z))<\epsilon_{(k,j)}.$
+We set $\mathcal{I}_{(k,j)}=\{z=z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$, and then define
+$\mathbb{F}=\cap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)},$
+where $\mathbb{F}_{(k,j)}=\cup_{z\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})$. One can verify that $\mathbb{F}\subset G_{K}^{\omega,U}$.
+We proceed to define a sequence of measures $\{\rho_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ by
+$\rho_{(k,j)}:=\frac{1}{\#\mathcal{I}_{(k,j)}}\sum_{x\in\mathcal{I}_{(k,j)}}\delta_{x}.$
+Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$.
+Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with $\mathbb{F}$, it holds that
 $\rho(B_{n}(z,\frac{\epsilon^{\ast}}{2}))\leq\exp\left[-n(1-\zeta)^{2}(H-2\zeta)\right].$
-Then according to Lemma 2.1, we have
+Then by the Lemma 2.1, we have
 $(1-\zeta)^{2}(H-2\zeta)\leq h_{top}^{B}(f,\mathbb{F},\frac{\epsilon^{\ast}}{2})\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$
@@
 is clear.
-On the other hand, let $\gamma>0$ and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}$.
-Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for sufficiently large $k$,
+On the other hand, let $\gamma>0$ be sufficiently small and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}$.Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for sufficiently large $k$,
 $M_{\gamma}<\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi).$
 Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows
-$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
+$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
 Letting $\gamma\to 0$, we have
@@
 We prove the Theorem 1.4 in this subsection.
-According to Theorem 2.4, the proof of Theorem 1.4 will be done, once we prove that for $\epsilon^{\ast}>0$, $\zeta>0$ small enough and $\delta\in(0,1)$, it holds
-$h_{top}(f,G_{K}^{\omega,U},\epsilon^{\ast})>\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
-That is, we must pick out enough separated points. Take $\mu^{\ast}\in M_{f}^{e}(X)$ with
-$\underline{h}_{\mu^{\ast}}^{K}(f,3\epsilon^{\ast},\delta)>H-\zeta,$
-where $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)>0$.
+The inequality
+$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)\leq\overline{\rm mdim}_{M}(f,X,d)$
+is clear. Thus, it suffices to show
+$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right).$
+Once we prove that for $\epsilon^{\ast}>0$, $\zeta>0$ small enough and $\delta\in(0,1)$, it holds that
+$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}),$
+then by Theorem 2.4, the proof of Theorem 1.4 will be done.
+Since the proof is similar to the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ in subsection 3.1.2, we give a sketch for it.
+We write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
+$\underline{h}_{\mu^{\ast}}^{K}(f,3\epsilon^{\ast},\delta)>H-\zeta.$
 Then for large enough $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
@@
 Now fix such $\mathcal{N}$, and write $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
-Since $K$ is a connected closed subset, a standard argument shows that there exists sequence $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
+Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
 $K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{k,i},\frac{1}{k}),\ D(\mu_{k,i},\mu_{k,i+1})<\frac{1}{k},\ \text{and }D(\mu_{k,Q_{k}},\mu_{k+1,1})<\frac{1}{k}.$
-Consider the set $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq k\leq Q_{k}\}$, $\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$ with lexicographic order.
-The sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed similar as in subsection 3.1.2.
-For nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$.
+Consider the sets $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, and
+$\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$
+with the lexicographic order.
+The sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed similarly as in subsection 3.1.2.
+For the nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$.
 Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$.
-For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be same as in subsection 3.1.2.
+For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be the same as in subsection 3.1.2.
 For $(k,i)\in\mathcal{A}^{\prime}$, since $G_{\mu_{k,i}}\neq\emptyset$, one can pick $x_{k,i}\in X$ and $n_{(k,i)}\in\mathbb{N}$ such that
 $D(\mathcal{E}_{n_{(k,i)}}(x),\mu_{k,i})<\delta_{(k,i)},\ \frac{n_{(k,i)}}{n_{(k,i)}+m_{(k,i)}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\}.$
-Set $\mathcal{C}_{(k,i)}=\{x_{k,i}\}$.
-Now choose a strictly increasing sequence of integers $\{N_{(i,j)}\}_{(k,j)\in\mathcal{A}\ \text{with }j\neq 0}$ with $N_{(1,1)}=1$ and
-$\begin{split}n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right];\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
-Based on $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar with previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ and the fractal set
+Set $\mathcal{C}_{(k,i)}=\{x_{k,i}\}$. Now choose a strictly increasing sequence of integers $\{N_{(i,j)}\}_{(k,j)\in\mathcal{A}\ \text{with }j\neq 0}$ with $N_{(1,1)}=1$ and
+$\begin{split}&n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right],\\ &\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
+Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain the fractal set
 $\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}\subset G_{K}^{\omega,U}.$
@@
 Thus
-$h_{top}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
+$h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
 4 Applications
@@
 $\varphi$ consisting of the set of points of $X$ such that $\varphi$ is taking constant value $\alpha$ on $V_{f}(x)$.
-Then continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
-Condition A: For each given $\varphi$ and $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
+Then the continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
+Condition A: Given a continuous function $\varphi$ on $M_{f}(X)$, for every $\alpha\in L_{\varphi}$ we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
 Some examples of $\varphi$ satisfying Condition A are given as follows:
@@
   • (c)
-    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i,e.
+    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i.e.
 for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$.
@@
     Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for any $\Phi$ and $\alpha\in L_{\Phi}$.
-    Specially, when sequence $\log\phi_{n}(x)=S_{n}\psi(x)$ is exactly an additive function, where $\psi\in C(X)$ and $S_{n}\psi(x)=\sum_{i=0}^{n-1}\psi(f^{i}(x))$, in this case we have
-    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\varphi\mathop{}\!\mathrm{d}\mu=\alpha\},$
-    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\varphi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
-Given the nonempty open subset $U\subset X$, we put
+    Specially, when the sequence $\{\phi_{n}\}$ is an additive function $\{S_{n}\psi\}$ with $\psi\in C(X)$, in this case we have
+    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\psi\mathop{}\!\mathrm{d}\mu=\alpha\},$
+    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
+Given a nonempty open subset $U\subset X$, we put
 $K_{\alpha}^{\omega,U}=K_{\alpha}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}.$
@@
 Let $(X,d,f)$ be a dynamical system with specification property and $U\subset X$ be a nonempty open subset.
-Suppose that let $\varphi:M_{f}(X)\to\mathbb{R}$ be a continuous function satisfying Condition A.
-For any $\alpha\in L$ and $S\in\{B,P\}$, we have
+Suppose that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A.
+Then for any $\alpha\in L$ and $S\in\{B,P\}$, we have
 $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha}^{\omega,U},d)=$ $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha},d)$
-$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 […]
 Proof.
-Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$ which is not dependent on $\mu$, such that
+Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$, which is not dependent on $\mu$, such that
 $\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$ (4.1)
@@
 $E_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}$
-denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then we have
-$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}\xi<\epsilon}h_{\mu}(f,\xi),$
+denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then
+$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
 […]
@@
 […]
-$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}$
-satisfies Condition A.
+$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
 Fix $x\in X$.
@@
 ∎
-Readers can turn to [O03, Tho09, ZC13] and references therein for the multifractal analysis of level sets using topological pressure.
-Then by Theorem 4.1 one can get analogous results for the metric mean dimensions of level sets of some special functions determined by the dynamical behaviors of the system.
+Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure.
+Then, by Theorem 4.1, one can get analogous variational principles for the metric mean dimensions of level sets of some special functions, which are determined by the dynamical behaviors of the system.
 4.2 Linking metric mean dimension and mean Li-Yorke chaotic
 In this subsection, we link the concepts of metric mean dimension and mean Li-Yorke chaos in the sense of infinite entropy.
-We give an quantitative characterization for the chaotic phenomenon of infinite entropy systems.
+We give a quantitative characterization for the chaotic phenomenon of infinite entropy systems.
 Namely, we show the packing metric mean dimension of the set of mean Li-Yorke pairs is bounded from blow by the metric mean dimension of the phase space (Theorem 4.3).
@@
 Theorem 4.3.
-Let $(X,d,f)$ be a dynamical system with specification property with positive topological entropy. Then we have
-$\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
+Let $(X,d,f)$ be a dynamical system with specification property. Then
+$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d),$
+$\overline{\rm mdim}_{M}^{UC}(f\times f,{\rm MLY}(f),d\times d)=\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d).$
 Proof.
@@
 Proof of the claim.
-By the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$.
+Specification property implies that the system has positive entropy.
+Then by the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$.
 We put $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$.
-Then we can choose an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfying the desired claim.
+Then any an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
 ∎
-Now take $\nu\in M_{f}(X)$ and set $\mu_{1}\in M_{f\times f}(X\times X)$ defined by $\mu_{1}=\phi_{\ast}\nu$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$.
+Now take $\nu\in M_{f}(X)$ and put $\mu_{1}=\phi_{\ast}\nu\in M_{f\times f}(X\times X)$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$.
 Then one has $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$.
@@
 For any $\gamma>0$, by Theorem 2.4 we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $\rho_{k}\in M_{f}(X)$ such that for each $k$,
-$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
-Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then it is easy to check that for every $k$,
+$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$ (4.4)
+Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then for every $k$,
 $\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)=\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta).$ (4.5)
-Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ satisfying ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
+Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ so that ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
 $H_{\rho_{k}^{\ast}}(\eta^{n})$ $=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=-\sum_{C\in\xi^{n}}\rho_{k}^{\ast}(C\times C)\log\rho_{k}^{\ast}(C\times C)$
@@
 which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\geq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. This shows (4.5).
-Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e., the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$.
+Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e.
+the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$.
 Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$.
-For any $(x_{0},y_{0})\in G_{K}$, there are sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$.
+For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$.
 Hence
@@
 $\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq$ $\overline{\rm mdim}_{M}^{P}(f\times f,G_{K},d\times d)$
-$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}\geq\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
-$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
-Since $\gamma$ is arbitrary, one has
-$\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
-The case of $\overline{\rm mdim}_{M}(f,d)=\infty$ is still valid by considering a slight modification. ∎
+$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
+$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
+$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}\geq\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
+Since $\gamma$ is arbitrary, by Proposition 2.1 one has
+$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
+The case of $\overline{\rm mdim}_{M}(f,X,d)=\infty$ is still valid by considering a slight modification.
+For sufficiently small $\gamma>0$, let $c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}$.
+By Theorem 2.4, we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
+$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f,\xi)>c_{\gamma}.$ (4.6)
+Put $\hat{K}=\overline{{\rm Conv}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}}$. Then the similar argument gives us
+$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d),$
+which implies the desired result by Proposition 2.1.
+Finally, we give an example to verify the Corollary 4.2 and Theorem 4.3.
+Example 4.4.
+Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric
+$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
+and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})\mapsto(x_{n+1})_{n\in\mathbb{Z}}$.
+It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf.
+[DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
+  • (1)
+    Thus, by the product formula of metric mean dimension one has ${\rm mdim}_{M}(\sigma\times\sigma,[0,1]^{\mathbb{Z}}\times[0,1]^{\mathbb{Z}},d\times d)=2$. Consequently, by Theorem 4.3, we have
+    $\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
+  • (2)
+    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, we have
+    $\overline{\rm mdim}_{M}^{B}(\sigma,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
+[…]
 5 Examples
-Applying the main results, we give the specific formula of the metric mean dimensions of generic points of full shift over compact metric space.
+Applying the main results, we give the specific formulae of the metric mean dimensions of generic points of full shift over compact metric space.
 Example 5.1.
-Endow the product space $X=[0,1]^{\mathbb{Z}}$ with product metric
-$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
-and let $\sigma:X\to X$ be the left shift map defined by $(x_{n})_{n\in\mathbb{Z}}\mapsto(x_{n+1})_{n\in\mathbb{Z}}$.
-It is shown that the system $(X,d,\sigma)$ has the specification property (cf.
-[DGS76, Proposition 21.2]), and ${\rm\overline{mdim}}_{M}(\sigma,X,d)=1$ [LT18].
-Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is the product measure, which is $\sigma$-ergodic.
-Then the ergodicity enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
-$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in X.$
+Consider the Example 4.4 again.
+Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ be the product measure, which is a $\sigma$-ergodic measure on $[0,1]^{\mathbb{Z}}$.
+Then the ergodicity of $\mu$ enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
+$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in[0,1]^{\mathbb{Z}}.$
 Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ24, Theorem 1.1], we have
-$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,X,d).$
-Then by Corollary 1.5, we have
+$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d).$
+Then, by Corollary 1.5, we have
 ${\rm\overline{mdim}}_{M}^{B}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{P}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{UC}(\sigma,G_{\mu},d)=1.$
@@
 Example 5.2.
-Let $K$ be a compact metric space with metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
+Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
 $d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
@@
 $\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K),$
-where $\overline{\rm dim}_{B}(F)$ is the upper box dimension of $F$ under the metric $d^{q}$.
+where $\overline{\rm dim}_{B}(K)$ is the upper box dimension of $K$ under the metric $d$.
 Theorem 5.3.
-Let $K$ be a compact metric space with metric $d$. There exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
-$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(F)$
-$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(\sigma,\xi).$
+Let $K$ be a compact metric space with a metric $d$. Then there exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
+$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K)$
+$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
 Proof.
 It is known that the system $(K^{\mathbb{Z}},\sigma)$ has specification property. Let $S_{\epsilon}$ be a maximal $\epsilon$-separated subset of $K$ with respect to metric $d$. For $n\geq 1$, we put
-$\hat{\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
-and consider the set $\mathcal{S}_{n}\subset X$ consisting of one preimage of each element of $\hat{\mathcal{S}_{n}}$ under $\pi_{n}$.
+${\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
+and consider the set $\hat{\mathcal{S}}_{n}\subset K^{\mathbb{Z}}$ consisting of one preimage of each element in ${\mathcal{S}_{n}}$ under $\pi_{n}$.
 Then $\hat{\mathcal{S}}_{n}$ is an $(n,\epsilon)$-separated set of $K^{\mathbb{Z}}$ and #$\#{\hat{\mathcal{S}}_{n}}=(\#S_{\epsilon})^{n}$.
@@
 $\mu_{n}:=\frac{1}{\#\hat{\mathcal{S}}_{n}}\sum_{x\in\hat{\mathcal{S}}_{n}}\mathcal{E}_{n}(x).$
-We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(X)$. Then by Lemma 3.1, for each $\epsilon>0$ we have
+We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(K^{\mathbb{Z}})$. Then, by Lemma 3.1, for each $\epsilon>0$ we have
 $\log\#S_{\epsilon}=\limsup_{n\to\infty}\frac{1}{n}\log\#\hat{\mathcal{S}}_{n}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(\sigma,\xi).$
-Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$. Combining with Theorem 1.3, one has
+Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$ as $\epsilon\to 0$. Combining with the Theorem 1.3, one has
 $\overline{\rm dim}_{B}(K)=\limsup_{\epsilon\to 0}\frac{\log\#S_{\epsilon}}{|\log\epsilon|}\leq\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm dim}_{B}(K).$
-Finally, by Corollary 1.5 we complete the proof.
+Finally, by the Corollary 1.5 we complete the proof.
 Availability of data and materials: Not applicable.
@@
 The third author was supported by the China Postdoctoral Science Foundation (Grant Number: 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University.
+The authors also thank Prof.
+Ercai Chen for many useful comments.

### v3

@@
-Yong Ji1, Junye Li1 and Rui Yang∗2,3,4
+Yong Ji1, Junye Li1 and Rui Yang∗2,3
 Abstract.
-Saturated set and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems.
-In the context of infinite entropy systems, this paper aims to give some qualitative aspects of saturated sets and the set of generic points in both topological and measure-theoretic perspectives.
-For systems with specification property, we establish the certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets have full metric mean dimension.
-All results are useful for understanding the topological structures of dynamical systems with infinite topological entropy.
-As applications, we further exhibit some qualitative aspects of metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
+Saturated sets and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems.
+In the context of infinite entropy systems, we aim to provide some qualitative aspects of saturated sets and the set of generic points from both topological and measure-theoretic perspectives.
+For systems with the specification property, we establish certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets has full metric mean dimension.
+As applications, we further present some qualitative aspects of the metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
 Keywords and phrases: Metric mean dimension; Saturated set; Generic point; Variational principle; Multifractal analysis
@@
 Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively.
-Measure-theoretic entropy and topological entropy are two vital invariants for us to understand the topological complexity of dynamical systems, which are related by the classical variational principle:
+Quantitative analysis using some entropy-like quantities is of great significance in understanding the dynamics of abstract dynamical systems.
+Measure-theoretic entropy and topological entropy are two crucial invariants that helps us understand the topological complexity of dynamical systems, which are related by the classical variational principle:
 $h_{top}(f)=\sup_{\mu\in M_{f}(X)}{h_{\mu}(f)},$
-where $h_{top}(f)$ is the topological entropy of $(X,f)$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$.
-In 1973, analogous to the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space.
-The pioneering work directly contributes to the development of dimension theory of dynamical systems, especially as a powerful tool to study the “size” of fractal-like sets in multifractal analysis.
-The multifractal analysis, which can date back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of the points with the similar dynamical behaviors(or known as multifractal spectrum), and then from the geometrical or topological viewpoint describes the “size” of those subsets by using fractal dimensions and dimensional entropies.
-Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, which is called phenomenon multifractal rigidity.
+where $h_{top}(f)$ is the topological entropy of $X$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$.
+In 1973, resembling the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space $X$.
+The pioneering work has directly contributed to the development of dimension theory in dynamical systems, especially serving as a powerful tool for studying the “size” of fractal-like sets in multifractal analysis.
+The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors(or known as multifractal spectrum), and then, from a geometrical or topological perspective, describes the “size” of those subsets by using fractal-like dimensions and dimensional-like entropies.
+Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, a phenomenon known as multifractal rigidity.
 We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
-Let us mention some relevant results about the dimensional entropies of saturated sets.
-Given $x\in X$, by $\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X)$, where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of the Birkhoff average w.r.t.
-$x$.
-Then the limit-point set $V_{f}(x)$ of $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ in the sense of weak-topology (cf.
+Let us mention some relevant results about the dimensional entropies of saturated sets to clarify the motivation.
+Given $x\in X$, by
+$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X),$
+where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of $x$.
+Then, in the sense of weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ (cf.
 [DGS76, Proposition 3.8]).
-Specially, when $V_{f}(x)=\{\mu\}$ for some measure $\mu$, one says that $x$ is a generic point of $\mu$.
-The set of all generic points of $\mu$ is denoted by $G_{\mu}$.
-Then the Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
-Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ is equal to $h_{\mu}(f)$, that is,
+Specially, when $V_{f}(x)=\{\mu\}$ for some invariant measure $\mu$, then $x$ is called a generic point of $\mu$.
+Denote by $G_{\mu}$ the set of all generic points of $\mu$.
+The Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
+In 1973, Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
 $h_{top}^{B}(f,G_{\mu})=h_{\mu}(f).$
-For non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$ and even easily constructs the example showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$ such that Bowen’s result is not valid.
-It is natural to expect whether there are some proper conditions such that the equality still holds for non-ergodic measures.
-A subset $D$ of $X$ is called a saturated set if $x\in D$ and $V_{f}(x)=V_{f}(y)$, then $y\in D$.
-More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we can obtain a saturated set $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$.
-For systems with specification property, including hyperbolic systems, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, which was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17].
-As for the estimation of Bowen topological entropy of saturated sets, for systems with $g$-almost product property that is weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments, and with uniform separation property (e.g.
-expansive and more generally asymptotically $h$-expansive systems), Pfister and Sullivan [PS07] established the certain variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
+For the case of non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there even exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$.
+Therefore, Bowen’s result fails for non-ergodic measures.
+It is natural to ask whether there are some proper conditions such that the Bowen’s equality still holds for non-ergodic measures.
+A positive answer to the above question is due to Pfister and Sullivan [PS07], who provided it within a more general framework.
+A subset $K$ of $X$ is called a saturated set if $x\in K$ and $V_{f}(x)=V_{f}(y)$, then $y\in K$.
+More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we have a saturated set by defining $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$.
+For any topological dynamical system (t.d.s) possessing the specification property, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, and then was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17].
+In 2007, Pfister and Sullivan [PS07] introduced the $g$-almost product property that is strictly weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments and the uniform separation property (e.g.
+expansive and more generally asymptotically $h$-expansive systems), and established the variational principle of Bowen topological entropy of nonempty compact connected set $K\subset M_{f}(X)$:
 $h_{top}^{B}(T,G_{K})=\inf\{h_{\mu}(f):\mu\in K\}.$
-Specially, only the assumption of $g$-almost product property they [PS07] showed
-$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$
-holds for all $\mu\in M_{f}(X)$, which extends Bowen’s remarkable result to non-ergodic measures.
-Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional entropies of saturated sets.
-Packing entropy of non-compact subsets was introduced by Feng and Huang [FH12] in a way resembling the definition of packing dimension.
-Additionally, under the (stronger) conditions of specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] showed the following variational principle for any nonempty compact connected set $K\subset M_{f}(X)$,
-$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\},$
-where $h_{top}^{P}(f,K)$ is the packing entropy of $K$.
-Afterwards, weakening the specification property, for systems with $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only proved that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed the upper capacity topological entropy of the saturated sets have full topological entropy, i.e.
+Furthermore, under only the assumption of $g$-almost product property, they [PS07] showed that for all $\mu\in M_{f}(X)$,
+$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f),$
+which extends Bowen’s remarkable result to non-ergodic measures.
+Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional-like entropies of saturated sets.
+Under the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained the variational principle of packing topological entropy of nonempty compact connected set $K\subset M_{f}(X)$,
+$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\}.$
+By weakening the specification property and considering systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only demonstrated that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed that the upper capacity topological entropy of the saturated sets has full topological entropy, i.e.,
 $h_{top}^{UC}(f,G_{K}^{C})=h_{top}(f),$
 where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$.
-Furthermore, for systems satisfying $g$-almost product property and uniform separation property, Huang, Tian and Wang [HTW19] showed Pfister and Sullivan’s variational principle is valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ denotes the set of transitive points.
-Readers can turn to [JCZ22, LLS17] for more results involving the dimensional entropies of saturated sets in the setting of nonuniformly hyperbolic systems, and [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for some results concerning the amenable Bowen and packing topological entropies of saturated sets in the context of the actions of amenable groups.
-According to the possible values of topological entropy, dynamical systems can be divided into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems.
-Infinite entropy system always means the system may possess quiet complicated topological complexity, so the notion of topological entropy no longer provides more information about the dynamics of the system.
+Additionally, Huang, Tian and Wang [HTW19] showed that Pfister and Sullivan’s variational principle is still valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ means the set of transitive points.
+The generalizations of the aforementioned results can be found in [JCZ22, LLS17] for the scenarios of nonuniformly hyperbolic systems, and in [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for the cases of the actions of amenable groups
+Based on the possible values of topological entropy, dynamical systems can be can be classified into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems.
+Infinite entropy systems typically imply that these systems may exhibit rather complicated topological complexity.
+As a result, the concept of topological entropy fails to offer more information about the dynamics of the system.
 In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00.
-It turns out that any finite entropy system has zero metric mean dimension.
+It has been shown that every system with finite entropy has a metric mean dimension of zero.
 See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b].
@@
 $\overline{\rm MRID}_{\mu}(f,X,d):$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
-which can be explained how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}\xi<\epsilon}\limits h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$.
-Some variational principles of metric mean dimension of $\mathbb{Z}$-actions can be found in [LW00, GS21, S22, YCZ23a].
-Then the aforementioned work suggests the following questions for saturated sets in infinite entropy systems:
+where the infimum ranges over the finite Borel measurable partition with diameter at most $\epsilon$.
+This measures how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$.
+Several variational principles regarding the metric mean dimension of $\mathbb{Z}$-actions can be found in the works [LW00, GS21, S22, YCZ23a, YCZ25].
+Then, in the context of the infinite entropy systems, the aforementioned work suggests the following questions for saturated sets:
 Questions:
   • (1)
-    For ergodic measures, do the metric mean dimensions of the set of generic points equal to the mean Rényi information dimensions?
+    For ergodic measures, does the metric mean dimension of the set of generic points coincide with the mean Rényi information dimension?
   • (2)
-    For non-ergodic measures, under the assumption of specification-like property, are there some certain variational principles for metric mean dimensions of saturated sets?
+    For non-ergodic measures, under the assumption of specification-like property, are there certain variational principles for metric mean dimensions of saturated sets?
 If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
-For (1), the authors in [YCZ24] showed the packing metric mean dimension of the set of generic points equals to the mean Rényi information dimension, however it holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula in infinite entropy systems (cf.[YCZ24, §5, Question,(1)]); for (2), noting the uniform separation property implies the system has finite topological entropy, so extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges.
-Only the assumption of $g$-almost product property, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$, and showed that for any $\{\mu_{1},\mu_{2},...,\mu_{k}\}\subset M_{f}(X)$ and any non-empty compact connected set $K$ contained in the convex closure ${\rm cov}\{\mu_{1},...,\mu_{k}\}$,
-$\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Prior to the comparison of different types of specification-like property, we only focus on the systems with specification property and give a full answer to question (2).
-Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the following statements of main results involve more general (mixed) saturated sets.
+For (1), the authors in [YCZ25] showed that the packing metric mean dimension of the set of generic points is equal to the mean Rényi information dimension.
+However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems (cf.[YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges.
+Under the assumption of $g$-almost product property, for any non-empty compact connected set $K$ contained in the convex closure of finite invariant measures, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$.
+Prior to the comparison of the different types of specification-like property, we concentrate solely on the systems possessing the specification property and provide a comprehensive answer to question (2).
+Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the subsequent statements of main results involve more general (mixed) saturated sets.
 For simplicity, we write
@@
 Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions.
-There exist certain variational principles for Bowen and packing metric metric mean dimensions of saturated sets (Theorems 1.1 and 1.2), while the upper capacity metric mean dimension of saturated set has full metric mean dimension (Theorem 1.4).
+There are indeed certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3.
+Although the $\mu$-measure of $G_{\mu}$ is zero for every $\mu\in M_{f}(X)$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
 Theorem 1.1.
-Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
-$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
-$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
-where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t.
-$\xi$, and the infimum is taken over all finite Borel measurable partitions of $X$.
+Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
+$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$.
 Theorem 1.2.
-Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
-$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}\left(f,G_{K}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\},d\right)$
-$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
+$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
 where $\overline{\rm mdim}_{M}^{B}(f,Z,d)$ is the Bowen upper metric mean dimension of $Z$.
-Maybe, more satisfactory results are exchanging the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$), but until now the authors do not know whether there are some counter-examples of infinite entropy systems with specification property such that the following equalities are strict:
+Perhaps, more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$).
+However, up to now, the authors remain unaware of whether there exist any counterexamples among infinite entropy systems with the specification property for which the following equalities hold strictly
 $\sup_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d),$
@@
 Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures.
-For any closed set $K$ of $M_{f}(X)$ and $\mu\in K$, by $M_{K}(\mu)$ we denote the space of sequences of Borel probability measures in $K$ which converge to $\mu$ in the weak-topology.
-We respectively define the (modified) lower and upper mean Rényi information dimensions of $\mu\in K$ as
-$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi),$
-$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
-Then we have the following desired variational principles in terms of the modified lower and upper mean Rényi information dimensions.
+Let $K$ be a closed set of $M_{f}(X)$.
+For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology.
+We define the (modified) lower and upper mean Rényi information dimensions of $\mu$ as
+$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
+$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}$
+respectively. Then this allows us to obtain the the variational principles in terms of the (modified) lower and upper mean Rényi information dimensions.
 Theorem 1.3.
-Let $(X,d,f)$ be a dynamical system with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
+Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, then
 $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\},$
@@
 Theorem 1.4.
-Let $(X,d,f)$ be a dynamical system with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set, then
-$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}(f,X,d),$
-where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$ defined by separated sets.
-As a direct consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case is $K$ being exactly a single point.
-The Corollary 1.5 presents the metric mean dimensions of generic points of any invariant measures for systems with specification property, which extends the result given in [YCZ24, Theorem 1.1] to all non-ergodic measures.
+Let $(X,d,f)$ be a t.d.s. with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set. Then
+$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}(f,X,d),$
+where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$.
+As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K$ consists of exactly a single point.
+The Corollary 1.5 extends the result given in [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with specification property.
 Corollary 1.5.
-Let $(X,d,f)$ be a dynamical system with specification property. Then for any $\mu\in M_{f}(X)$,
+Let $(X,d,f)$ be a dynamical system with specification property. Then for every $\mu\in M_{f}(X)$,
 $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),$
 $\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
-Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$, then we have
+Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure in the sense of $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$, then we have
 $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
@@
 For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$.
-The concept of maximal metric mean dimension measure was proposed in [YCZ23a], yet for what kind of infinite entropy systems can have maximal metric mean dimension measure is still a difficult problem.
+The concept of maximal metric mean dimension measure was proposed in [YCZ23a].
+However, determining what types of infinite entropy systems can possess a maximal metric mean dimension measure remains a challenging task.
 See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms.
-By Gutman- Śpiewak variational principle (see Theorem 2.4) and Corollary 1.5 we have following (abstract) criterion for maximal metric mean dimension measure for systems with specification property.
+By Gutman- Śpiewak variational principle (as presented in Theorem 2.4) and Corollary 1.5, we have an abstract criterion for maximal metric mean dimension measure for identifying maximal metric mean dimension measures in systems with the specification property.
 Corollary 1.6.
-Let $(X,d,f)$ be a dynamical system with specification property.
-An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure such that $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi)$ if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
+Let $(X,d,f)$ be a t.d.s.
+with specification property.
+An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
 Some comments are in order.
   • (1)
-    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV] such that our main results can be applied to such systems.
+    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV], and our main results can be applied to these systems.
   • (2)
     In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension.
-The other aforementioned theorems are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
+The other theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
   • (3)
-    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by considering some modifications of the present proofs.
-The novelty of our work not only presents a proper and full extension of the previous work in [PS07, ZCC12, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for the multifractal analysis of other types of fractal-like sets in infinite entropy systems.
-It remains unclear whether there are some proper substitutions of the uniform separation property so that all results are true for $g$-almost product property, or beyond $\mathbb{Z}$-actions.
+    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by making some modifications to the current proofs.
+The novelty of our work not only gives a proper and full extension of the previous work in [PS07, ZCC12, ZC13, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for studying the multifractal analysis of other types of fractal-like sets in infinite entropy systems.
+It remains unclear whether there are some possible substitutions of the uniform separation property so that all the above results are true for $g$-almost product property, or even for cases beyond $\mathbb{Z}$.
 The paper is organized as follows.
@@
 In section 3, we prove the main results.
-In section 4, we give two applications of the main results.
-In section 5, we show the metric mean dimensions of generic points of full shift over compact metric spaces.
+In section 4, we consider some applications of the main results, and apply it to investigate the level sets of Birkhoff average of a continuous potential, Li-Yorke chaos and the metric mean dimensions of generic points of the full shift over compact metric spaces.
 2 Preliminaries
@@
 2.1 Three types of metric mean dimensions
-In this subsection, we recall the definitions of upper capacity metric mean dimension on subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension on subsets [YCZ22].
-Let $(X,d,f)$ be a dynamical system. For $n\in\mathbb{N}$, we define the Bowen metric $d_{n}$ on $X$ as
-$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\}$
-for any $x,y\in X$. Then the Bowen open ball and closed ball of $x$ in the metric $d_{n}$ with radius $\epsilon$ are respectively given by
+In this subsection, we review the definitions of upper capacity metric mean dimension of subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension [YCZ22] of subsets.
+Let $(X,d,f)$ be a t.d.s.. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on the space $X$ as
+$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\},$
+where $x,y\in X$. Then the Bowen open ball and closed ball centered at $x$ with radius $\epsilon$ in the metric $d_{n}$ are given by
 $B_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)<\epsilon\},$
-$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}.$
+$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
+respectively.
 Let $Z$ be a nonempty subset of $X$.
 Fix $\epsilon>0$.
-A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if $x,y\in E$ with $x\neq y$ implies $d_{n}(x,y)>\epsilon$.
-Let $s_{n}(Z,\epsilon)$ denote the largest cardinality of $(n,\epsilon)$-separated sets of $Z$.
+A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if any distinct $x,y\in E$ implies $d_{n}(x,y)>\epsilon$.
+Denote by $s_{n}(Z,\epsilon)$ the largest cardinality of $(n,\epsilon)$-separated sets of $Z$.
 The $\epsilon$-upper capacity topological entropy of $Z$ is given by
 $h_{top}^{UC}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
-Analogous to the definition of box dimension in fractal geometry, we can formulate the definition for upper capacity metric mean dimension of subsets using separated sets.
+Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, i.e.,
+$h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
+Specifically, topological entropy fails to distinguish different levels of topological complexity within the infinite entropy systems
+Analogous to the definition of box dimension in fractal geometry, we formulate the refined entropy-like quantity for infinite entropy systems using $\epsilon$-upper capacity topological entropy.
 Definition 2.1.
-We define the upper capacity metric mean dimension of $Z$ as
-$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$ (2.1)
-When $Z=X$, we write $h_{top}(f,X,\epsilon)=h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)=\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively.
-In this case, it is exactly reduced to Lindenstrauss and Weiss’s classical notion introduced in [LW00].
-Such a formulation allows us to extend the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
-Recall that the upper capacity topological entropy of $Z$ is limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, that is,
-$h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
-Especially, for any system admitting the infinite topological entropy, topological entropy does not provide the additional information of the dynamics except we have known the system has pretty complicated topological complexity.
-Thus, (upper capacity) metric mean dimension can be viewed as a quantity to detect how fact the $\epsilon$-upper capacity topological entropy diverges to the infinite (upper capacity) topological entropy as $\epsilon$ tends to 00.
-More precisely, for sufficiently $\epsilon>0$, we may think of
+The upper capacity metric mean dimension of $Z$ is defined by
+$\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$
+In the case where $Z$ is equal to $X$, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$ respectively.
+Then the upper capacity metric mean dimension of $X$ is reduced to Lindenstrauss-Weiss’s definition introduced in [LW00].
+Such a formulation extends the concept of the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
+Notice that the (upper capacity) metric mean dimension is defined by detecting how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon$ tends to 00.
+That is, for sufficiently small $\epsilon>0$, we may think of
 $h_{top}^{UC}(f,Z,\epsilon)\approx\overline{\rm mdim}_{M}^{UC}(f,Z,d)\cdot|\log\epsilon|.$
-It is easy to see that any finite entropy system has zero metric mean dimension, and hence verifies that metric mean dimension is a good candidate to capture the dynamics of infinite entropy systems.
-Similar to the definitions of Hausdorff dimension and packing dimension in fractal geometry, as an analogue of the classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which is defined by Carathéodory-Pesin structures, the corresponding dimensional characterizations of metric mean dimensions via Carathéodory-Pesin structures are Bowen and packing metric mean dimensions, which has been introduced in [LV21, YCZ22, W21] and studied by several authors in [BR23, CLS21, LL24, FKO24, YCZ24].
+It is straightforward to observe that any system with finite topological entropy has zero metric mean dimension.
+This, in turn, shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
+The classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogue of the Hausdorff dimension and packing dimension in fractal geometry.
+The corresponding dimensional characterizations of metric mean dimensions, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the Bowen and packing metric mean dimensions.
+One may refer to [BR23, CLS21, LL24, FKO24, YCZ25] for the results involving Bowen and packing metric mean dimensions.
+For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
+$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i\in I}\exp(-n_{i}s)\right\},$
+where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
+The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases, this ensures that the limit
+$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon)$
+exists.
+There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00.
+That is, the critical value is given by
+$h_{top}^{B}(f,Z,\epsilon)=\inf\{s>0:M(Z,s,\epsilon)=0\}=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
 Definition 2.2.
-For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
-$M(Z,s,N,\epsilon)=\inf\left\{\sum_{i}\exp(-n_{i}s)\right\},$
-where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}$ such that $Z\subset\cup_{i}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
-The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So we set
-$M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon).$
-There is a critical value of parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
-$h_{top}^{B}(f,Z,\epsilon)$ $=\inf\{s>0:M(Z,s,\epsilon)=0\}$
-$=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
-We define the Bowen upper metric mean dimension of $Z$ as
+The Bowen upper metric mean dimension of $Z$ is defined by
 $\overline{\rm mdim}^{B}_{M}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{B}(f,Z,\epsilon)}{|\log\epsilon|}.$
-Definition 2.3.
 For $Z\subseteq X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
-$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i}\exp(-sn_{i})\right\},$
-where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i$.
+$P(Z,s,N,\epsilon)=\sup\left\{\sum_{i\in I}\exp(-sn_{i})\right\},$
+where the supremum is taken over all finite or countable pairwise disjoint families $\{\overline{B}_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ with $x_{i}\in Z$, $n_{i}\geq N$ for all $i\in I$.
 Since the quantity $P(Z,s,N,\epsilon)$ is non-increasing as $N$ increases, and thus the following limit exists:
@@
 $\mathcal{P}(Z,s,\epsilon):=\inf\left\{\sum_{i=1}^{\infty}P(Z_{i},s,\epsilon):\bigcup_{i=1}^{\infty}Z_{i}\supseteq Z\right\}.$
-There is a critical value of parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00, that is, the critical value is given by
+There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00.
+That is, the critical value is given by
 $h_{top}^{P}(f,Z,\epsilon)$ $=\inf\{s>0:\mathcal{P}(Z,s,\epsilon)=0\}$
 $=\sup\{s>0:\mathcal{P}(Z,s,\epsilon)=\infty\}.$
-We define the packing upper metric mean dimension of $Z$ as
+Definition 2.3.
+The packing upper metric mean dimension of $Z$ is defined by
 $\overline{\rm mdim}_{M}^{P}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,Z,\epsilon)}{|\log\epsilon|}.$
@@
 Proposition 2.1.
-$(1)$ If $Z_{1}\subset Z_{2}$, then
-$\overline{\rm mdim}_{M}^{UC}(f,Z_{1},d)$ $\leq\overline{\rm mdim}_{M}^{UC}(f,Z_{2},d)$
-$\overline{\rm mdim}^{B}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{B}_{M}(f,Z_{2},d)$
-$\overline{\rm mdim}^{P}_{M}(f,Z_{1},d)$ $\leq\overline{\rm mdim}^{P}_{M}(f,Z_{2},d).$
+$(1)$ If $Z_{1}\subset Z_{2}\subset X$, then for each $S\in\{UC,B,P\}$,
+$\overline{\rm mdim}_{M}^{S}(f,Z_{1},d)\leq\overline{\rm mdim}_{M}^{S}(f,Z_{2},d).$
 $(2)$ If $Z=\bigcup_{i=1}^{\infty}Z_{i}$, then for every $\epsilon>0$,
@@
 and $h_{top}^{UC}(f,Z,\epsilon)=\max_{1\leq i\leq N}h_{top}^{UC}(f,Z_{i},\epsilon)$ if $Z=\bigcup_{i=1}^{N}Z_{i}$.
-$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then
-$\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{UC}(f,Z_{i},d)$
-$\overline{\rm mdim}^{B}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{B}_{M}(f,Z_{i},d)$
-$\overline{\rm mdim}^{P}_{M}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}^{P}_{M}(f,Z_{i},d).$
+$(3)$ If $Z=\bigcup_{i=1}^{N}Z_{i}$, then for each $S\in\{UC,B,P\}$,
+$\overline{\rm mdim}_{M}^{S}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{S}(f,Z_{i},d).$
 $(3)$ If $Z$ is a nonempty subset of $X$, then for every $10<\epsilon<$ 1, we have
 $h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}^{UC}(f,Z,\epsilon).$
-Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$; if $Z$ is $f$-invariant and compact, then the equalities hold.
-The metric mean dimension of a given subset is in general not easy to calculate its precise value.
-However, the Entropy Distribution Principle allows us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of some fractal-like sets in multifractal analysis of dynamical systems.
-Consequently, the powerful tool gives the lower bounds of entropy and metric mean dimension of the fractal-like sets.
+Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$.
+Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions are equal.
+Given a subset of phase space, it is not easy to calculate its precise Bowen and packing metric mean dimensions.
+However, in the multifractal analysis of dynamical systems the Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets.
+Consequently, the powerful tool yields the lower bounds of metric mean dimension of the fractal-like sets.
+$\epsilon$-Bowen and packing topological entropies of Entropy Distribution Principles are stated as follows:
 Lemma 2.1.
-[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Let $Z$ be a Borel set of $X$ and $s\geq 0$. If there exist a Borel probability measure $\mu\in M(X)$ and a constant $C>0$ satisfying:
-$(1)$ for every $n$, $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for each Bowen ball $B_{n}(x,\epsilon)$ such that $B_{n}(x,\epsilon)\cap Z\neq\emptyset$;
-$(2)$ $\mu(Z)>0$,
-then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
-$\epsilon$-packing topological entropy also admits an analogous Entropy Distribution Principle as follows:
+[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s.
+and $\epsilon>0$.
+Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$.
+If there exist a $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and for every $n$ and for each Bowen ball $B_{n}(x,\epsilon)$ with $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, one has $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$, then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
 Lemma 2.2.
@@
 and $\epsilon>0$.
-Let $Z$ be a Borel set of $X$ and $s\geq 0$.
-If there exist a Borel probability measure $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ of positive integer that converges to $\infty$ satisfying:
-$(1)$ $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$;
-$(2)$ $\mu(Z)>0$,
-then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
+Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$.
+If there exist a $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ that converges to $\infty$ satisfying $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
 Proof.
-For sufficiently small $\delta>0$, by definition it is enough to prove that
+By definition, it is enough to prove that for sufficiently small $\delta>0$,
 $P\left(E,s-\delta,\frac{\epsilon}{6}\right)=\infty$
 holds for any Borel subset $E\subset Z$ with $\mu(E)>0$.
-For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$.
-Then by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint, such that
-$E\subset\bigcup_{j}B_{n_{i}}(x_{j},\epsilon).$
-Hence by condition $(1)$, one has
-$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
-Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$. ∎
-2.2 Measure-theoretic $\epsilon$-entropy
+For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of the set $E$.
+Then, by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ Hence, one has
+$P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j\in J}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
+Letting $i\to\infty$, we have $P(E,s-\delta,\frac{\epsilon}{6})=\infty$.
+This shows that $\mathcal{P}(Z,s-\delta,\frac{\epsilon}{6})=\infty$, and hence $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$ by letting $\delta\to 0$.
+∎
+2.2 Measure-theoretic $\epsilon$-entropy of invariant measures
 In this subsection, we recall the definitions of Kolmogorov-Sinai $\epsilon$-entropy and Katok $\epsilon$-entropy of invariant measures.
-Let $C(X)$ denote the space of continuous functions on $X$ equipped with the supermum norm.
-The Riesz representation theorem on compact metric space gives us an approach to endow with the weak-topology on $M(X)$, which is metrizable by the following metric
-$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$
-where $\{\phi_{k}\}$ is a family of countable dense subset of $C(X)$ taking values in $[0,1]$.
-Thus, the diameter of $M(X)$ in sense of metric $D$ is always less than $1$, and for fixed $\mu$ and $\nu$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$.
-It follows from the weak-topology on $M(X)$ that the natural projection from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous.
-Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in sense of weak-topology.
-Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\bigcap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\bigcup_{i=1}^{k}C_{i})=0$.
-The partition entropy of $\xi$ w.r.t.
+Let $C(X)$ denote the space consisting of real-valued continuous functions defined on $X$, equipped with the supremum norm.
+By virtue of the Riesz representation theorem for compact metric spaces, we are able to endow $M(X)$ with the weak-topology, which is metrizable by the following metric:
+$D(\mu,\nu):=\sum_{k=1}^{\infty}\frac{|\int\phi_{k}\mathop{}\!\mathrm{d}\mu-\int\phi_{k}\mathop{}\!\mathrm{d}\nu|}{2^{k}},$ (2.1)
+where $\{\phi_{k}\}$ is a countable family of dense subsets of $C(X)$ taking values in $[0,1]$.
+Thus, the diameter of $M(X)$ in the sense of the metric $D$ is always less than $1$, and for fixed $\mu,\nu\in M(X)$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$.
+It follows from the weak-topology on $M(X)$ that the natural embedding from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous.
+Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the sense of the weak-topology.
+Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\cap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\cup_{i=1}^{k}C_{i})=0$.
+Then the partition entropy of $\xi$ w.r.t.
 $\mu$ is defined by
 $H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$
-Then we define measure-theoretic entropy of $f$ w.r.t. $\xi$ as
+We define the measurable partition of $\mu\in M_{f}(X)$ w.r.t. $\xi$ as
 $h_{\mu}(f,\xi)=\lim_{n\to+\infty}\frac{1}{n}H_{\mu}(\xi^{n}),$
-where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $f$ w.r.t. $\mu$ is defined as
+where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
 $h_{\mu}(f)=\sup_{\xi}h_{\mu}(f,\xi),$
-where the supremum is taken over all finite Borel measurable partitions $\xi$ of $X$.
-For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions on subsets is the measure-theoretic metric mean dimensions of invariant measures.
-Unlike topological entropy, measure-theoretic entropy of ergodic measures admit several equivalent definitions by considering the limits of the measure-theoretic $\epsilon$-entropies: for instance, Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others.
-A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ24], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf.
-[YCZ24, Theorem 1.1]).
-Here, we only recall two types of them.
+where the supremum is taken over all finite Borel measurable partitions of $X$.
+For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions is the measure-theoretic metric mean dimensions of invariant measures by measuring the divergent rates of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others.
+A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf.
+[YCZ25, Theorem 1.1]).
+Here, we only recall two types of the measure-theoretic $\epsilon$-entropies.
   • (a)
-    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu$ is given by
+    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu\in M_{f}(X)$ is given by
     $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
     $N_{\mu}(n,\epsilon,\delta):=\min\{\#E:E\subset X~{}\emph{and}~{}\mu(\cup_{x\in E}B_{n}(x,\epsilon))>1-\delta\}.$
-    The (lower) Katok $\epsilon$-entropy of $\mu$ is given by
+    The lower Katok $\epsilon$-entropy of $\mu$ is given by
     $\underline{h}_{\mu}^{K}(f,\epsilon,\delta)=\liminf_{n\to\infty}\frac{1}{n}\log N_{\mu}(n,\epsilon,\delta).$
-The relation between Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy is related by the following inequalities:
+The Katok $\epsilon$-entropy and Kolmogorov-Sinai $\epsilon$-entropy are related by the following inequalities:
 Lemma 2.3.
-[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})<4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$.
+[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$.
 Then for any $\delta\in(0,1)$, one has
@@
 Consequently, $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
-The next variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
+The following variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
 Theorem 2.4.
@@
 where $F(\mu,\epsilon)\in\{\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),\underline{h}^{K}_{\mu}(f,\epsilon,\delta)\}$.
-A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is given in [YCZ24, Theorems 1.2 and 1.3].
+The results are also valid for lower metric mean dimension.
+A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is summarized in [YCZ25, Theorems 1.2 and 1.3].
 2.3 Specification property
-Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments using specification-like property.
-Definition 2.4 (Specification property).
-One says that a t.d.s.
-$(X,f)$ has specification property if for any $\epsilon>0$, there is an integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists $x\in X$ such that
-$d(f^{p+a_{j}}x,f^{p}x_{j})<\epsilon$
+Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as specification property.
+Definition 2.4.
+A t.d.s.
+$(X,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
+$d(f^{p}x_{j},f^{p+a_{j}}x)<\epsilon$
 for any $p=0,\cdots,b_{j}-a_{j}$, $j=1,\cdots,k$.
-In other words, that point $x$ $\epsilon$-shadows the $k$ pieces of orbits
+In other words, the point $x$ $\epsilon$-shadows the $k$ pieces of finite orbits of $x_{j}$:
 $\{x_{j},f(x_{j}),\ldots,f^{b_{j}-a_{j}}(x_{j})\},\ j=1,\ldots,k.$
@@
 More examples of infinite entropy systems with specification property can be found in [BR23, Section IV].
-A point $x\in X$ is almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, it holds that $f^{k}(x)\in U$ for some $n\leq k\leq n+N$.
+A point $x\in X$ is called almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, there exists an integer $k$ with $n\leq k\leq n+N$ for which $f^{k}(x)\in U$.
 Let $AP(X)$ denote the set of all almost periodic points of $X$.
-The following result, which was proved for systems with $g$-almost product property [HTZ23, Proposition 2.11], gives a characterization of the structure of $AP(X)$.
+The following statement, proved in [HTZ23, Proposition 2.11] for systems with $g$-almost product property, gives a characterization of the structure of $AP(X)$.
 Proposition 2.2.
@@
 3 Proofs of main results
-This section aims to prove the main results of this paper.
-Essentially, the difficult part is the estimation of lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$.
-We shall construct a Moran fractal set containing in $G_{K}^{\omega,U}$, which is a limit set with geometric construction.
-Then such set is suitable for the purpose of the computation of the lower bound of metric mean dimensions of saturated sets.
+This section proves the main results of this paper.
+The challenging aspect lies in the estimation of the lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$.
+We will construct a Moran fractal set, which is geometrically constructed and is contained in $G_{K}^{\omega,U}$.
+Consequently, this set is well-suited for computing the lower bound of the metric mean dimensions of saturated sets.
 3.1 Packing metric mean dimension of saturated sets
-We first prove Theorem 1.1. To arrive this goal, we divide the proof of the Theorem 1.1 into two parts.
+We first prove Theorem 1.1. To arrive at this goal, we divide its proof into two parts.
 3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
-We adapt a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to obtain the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K})$.
+We employ a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to derive the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$.
 Lemma 3.1.
-[Y23, Lemma 3.1] Let $(X,f)$ be a dynamical system and $\epsilon>0$.
+[Y23, Lemma 3.1] Let $(X,f)$ be a t.d.s.
+and $\epsilon>0$.
 Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality.
@@
 Proof.
-This essentially follows the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 9.10]). ∎
-For $K\subset M_{f}(X)$, we set $G^{K}:=\{x\in X:V_{f}(x)\subset K\}$. Obviously, we have the fact $G_{K}\subset G^{K}$.
+This essentially adheres to the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
+Recall $G_{K}=\{x\in X:V_{f}(x)=K\}$. For $K\subset M_{f}(X)$, we set
+$G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
+Obviously, we have $G_{K}\subset G^{K}$.
 Lemma 3.2.
@@
 Proof.
-It suffices to show $(1)$, since one can take $K=\{\mu\}$ as in $(1)$ and use the fact $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
-For $n\in\mathbb{N}$, $\delta>0$, we put
-$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)~{}\text{for infinitely many }n>k\right\},$
+It suffices to show $(1)$ since one can take $K=\{\mu\}$ as in $(1)$ and use $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
+For $n\in\mathbb{N}$ and $\delta>0$, we put
+$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~{}\forall n\geq k\right\},$
 $\mathcal{Q}_{K}(\delta,n):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
-and $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(K,\delta,n,\epsilon),$ where $S(K,\delta,n,\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$.
+and
+$\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon),$
+where $s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$.
 Fix $x\in G^{K}$.
-Then, we can choose a subsequence $\{n_{i}\}$ such that $\mathcal{E}_{n_{i}}(x)\to\mu$ for some $\mu\in K$.
-This implies for every $\delta>0$, there exists a positive integer $k$ such that $\mathcal{E}_{n_{i}}(x)\in B(K,\delta)$ for sufficiently large $n_{i}$ greater than $k$.
-Then by the relation $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ holds for every $\delta$, it follows from Proposition 2.1 that
-$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon),$
-which implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon).$
+Since we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$, then the index $n$ with $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite.
+This shows that $x\in\mathcal{R}_{K}(\delta,k)$ for some $k$.
+Then, by $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
+$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon).$ (3.1)
+For any $\tau>0$, there exists $\delta_{0}>0$ such that
+$\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
+For any $k\geq 1$, by definitions, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
+$h_{top}^{UC}(f,\mathcal{R}_{K}(\delta_{0},k),\epsilon)\leq\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
+Combining (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
 Next we shall prove for any $\gamma>0$, there is $\mu\in K$ such that
 $\Theta(K,\epsilon)-\gamma\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Take a strictly decreasing positive sequence $\{\delta_{k}\}$ with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
-$S(K,\delta_{k},n_{k},\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
-Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$, and set
-$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x)\in B(K,\delta_{k}).$
-Choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$. Then $\mu\in K$ since $K$ is a closed convex set. By Lemma 3.1, we have
-$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log S(K,\delta_{k_{j}},n_{k_{j}},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-This completes the proof. ∎
-Thus, by the Lemma 3.2 we obtain an upper bound of the packing metric mean dimension of $G_{K}$:
-$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{P}(f,G_{K},\epsilon)}{|\log\epsilon|}\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.1)
+Take a strictly decreasing sequence $\{\delta_{k}\}$ of positive real numbers with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
+$s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
+Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$. We set
+$\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
+Then $\mu_{k}\in B(K,\delta_{k})$ since $K$ is convex.
+In the weak-topology, choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$.
+Then $\mu\in K$ since $K$ is a closed.
+By Lemma 3.1, we have
+$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+This implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
+Thus, by the Lemma 3.2 the upper bound of the packing metric mean dimension of $G_{K}$ is given by
+$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.2)
 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
-For any small enough $\epsilon^{\ast}>0$, we shall prove that
+For any small enough $\epsilon^{\ast}>0$, we prove that
 $\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$ (3.3)
-One may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$. Then take $\mu^{\ast}\in K$ such that
+Then the above inequality yields the desired lower bound.
+Without loss of generality, one may assume $H:=\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)>0$. Let $\zeta>0$ and then take $\mu^{\ast}\in K$ such that
 $\inf_{\mathrm{diam}(\xi)<20\epsilon^{\ast}}h_{\mu^{\ast}}(f,\xi)>H-\zeta.$ (3.4)
-Since $K$ is connected and compact, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
+Since $K$ is a connected and compact set, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
 […]
 $D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})$ $\leq\frac{1}{k},\ \mu_{(k,Q_{k})}=\mu^{\ast}.$
-Now the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with lexicographic order.
+Now endow the set $\mathcal{A}=\{(0,0)\}\cup\{(k,j):k\geq 1,0\leq j\leq Q_{k}\}$ with the lexicographic order.
 Denote by $\iota(k,j)$ the position of $(k,j)\in\mathcal{A}$.
@@
 Let $\{\delta_{(k,j)}\}$ be a strictly decreasing sequence indexed by $\mathcal{A}$ with $\delta_{(0,0)}<\zeta$, $\delta_{(k,j)}\to 0$ as $(k,j)\to\infty$.
-We shall write $\delta_{k}:=\delta_{(k,Q_{k})}$.
-For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is a positive integer given in the definition of specification property.
+We shall simply write $\delta_{k}=\delta_{(k,Q_{k})}$.
+For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is chosen as in the definition of specification property.
 For any nonempty open set $U$, there exist $\tilde{\epsilon}>0$ and a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$.
@@
 $\mathcal{C}_{(0,0)}=\{a_{0}\},\ n_{(0,0)}=1,\ N_{(0,0)}=1.$
-For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$ which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2.
-Then according to specification property, there exists $a_{k}\in X$ such that
+For any $k\in\mathbb{N}$, there exists $\Delta_{k}=\{a_{1}^{k},\ldots,a_{q_{k}}^{k}\}\subset AP(X)$, which is $\epsilon_{(k,0)}$-dense in $X$ by Proposition 2.2.
+By the specification property, there exists $a_{k}\in X$ such that
 $d\left(f^{(i-1)m_{(k,0)}+i-1}(a_{k}),a_{i}^{k}\right)<\epsilon_{(k,0)},\ i=1,2,\ldots,q_{k}.$
-Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$, $N_{(k,0)}=1$.
+Write $\mathcal{C}_{(k,0)}=\{a_{k}\}$, $n_{(k,0)}=q_{k}+(q_{k}-1)m_{(k,0)}$ and $N_{(k,0)}=1$.
 Let $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$ and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$.
-The following measure approximation of $\mu^{*}$ derived from a slight modification of [LL24, Lemma 2.1], so we omit the proof.
+The following measure approximation of $\mu^{*}$ is a slight modification of [LL24, Lemma 2.1], so we omit the proof.
 Lemma 3.3.
-For each $k\in\mathbb{N}$, there is a measure $\nu_{k}\in M_{f}(X)$ satisfying
+For each $k\in\mathbb{N}$, there exists $\nu_{k}\in M_{f}(X)$ satisfying
 $(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\{\lambda_{i}^{k}:1\leq i\leq j_{k}\}$ are rational positive numbers, $\forall 1\leq i\leq j_{k}$ with $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$ and $\nu_{i}^{k}\in M_{f}^{e}(X)$;
 $(2)$ one has
-$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta.$ (3.4)
-$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$.
-By the ergodicity of $\nu_{i}^{k}$, there exists $N\in\mathbb{N}$ such that the set
+$\inf_{\xi\succ\mathcal{U}}h_{\mu^{\ast}}(f,\xi)\leq\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)+\zeta;$ (3.5)
+$(3)$ $D(\mu^{\ast},\nu_{k})<\delta_{k}$, where $D$ is the compatible metric on $M(X)$ given in (2.1).
+Since $\nu_{i}^{k}$ is ergodic, by Birkhoff ergodic theorem there exists $N\in\mathbb{N}$ such that the set
 $Y_{i}(N)=\left\{x\in X:D(\mathcal{E}_{n}(x),\nu_{i}^{k})<\delta_{k},\ \forall n>N\right\}$
-has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3 one has
+has $\nu_{i}^{k}$-measure greater than $1-\zeta$ for all $1\leq i\leq j_{k}$. Applying Lemma 2.3, one has
 $\underline{h}_{\nu_{i}^{k}}^{K}(f,20\epsilon^{\ast},\zeta)\leq\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)\leq\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)<\infty.$
 Now take an integer $n_{k}^{\prime}\in\mathbb{N}$ such that
-$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer holds for all $1\leq i\leq j_{k}$, and
+$(1)$ $\lambda_{i}^{k}n_{k}^{\prime}$ is an integer for all $1\leq i\leq j_{k}$, and
 $\frac{n_{k}^{\prime}}{n_{k}}>\max\left\{1-\zeta,1-\frac{1}{k}\right\},$ (3.6)
 where $n_{k}=n_{(k,Q_{k})}=n_{k}^{\prime}+(j_{k}-1)m_{(k,Q_{k})}$.
-$(2)$ there exists a subset $E_{k,i}$ which is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
-#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right].$
+$(2)$ there exists a subset $E_{k,i}$ that is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
+#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
+where the second inequality is derived from Lemma 2.3.
 $(3)$ for every $x\in E_{k,i}$,
@@
 where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$.
-One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniform continuous.
-For any distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$.
-Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$,
+One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniformly continuous.
+For any two distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$.
+Indeed, if $x_{i}\neq x_{i}^{\prime}$ for some $1\leq i\leq j_{k}$, then
 $5\epsilon^{\ast}<d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},x_{i}^{\prime}\right)\leq d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(x_{i},f^{T_{i}}y\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right)+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y^{\prime},x_{i}^{\prime}\right)$
@@
 $\mathcal{C}_{(k,Q_{k})}=\left\{y(x_{1},x_{2},\ldots,x_{j_{k}}):\ (x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}\right\}.$
-Then $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set and
+Then $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set and
+#$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
 […]
-The last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and formulas (3.3), (3.5).
-For $n\in\mathbb{N}$, $\varphi\in C(X)$, write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$. Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X,[0,1])$ taking values in $[0,1]$,
+${\geq}\exp\left[n_{k}(1-\zeta)\left(H-3\zeta\right)\right],$ (3.7)
+where the last inequality holds since ${\rm diam}(\xi)<20\epsilon^{\ast}$ and (3.4), (3.6).
+For $n\in\mathbb{N}$ and $\varphi\in C(X)$, we write $S_{n}\varphi=\sum_{l=0}^{n-1}\varphi\circ f^{i}$ as the $n$-th Birkhoff sum of $\varphi$.
+Then for $y\in\mathcal{C}_{(k,Q_{k})}$ and $\varphi\in C(X)$ taking values in $[0,1]$,
 $\left|\int\varphi\mathop{}\!\mathrm{d}\mathcal{E}_{n_{k}}(y)-\int\varphi\mathop{}\!\mathrm{d}\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y))\right)\right|$
@@
 $\leq$ $\left[\left|\frac{1}{n_{k}^{\prime}}-\frac{1}{n_{k}}\right|\cdot n_{k}^{\prime}+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\right]\cdot\|\varphi\|<\frac{2}{k}\|\varphi\|,$
-which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Recall we have
-$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k}.$
-By the construction, one has
+which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Since
+$D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k},$
+one has that for any $y\in\mathcal{C}_{(k,Q_{k})}$,
 $D\left(\mathcal{E}_{n_{k}}(y),\mu^{\ast}\right)\leq\frac{2}{k}+D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i})\right)+2\delta_{k}<\frac{3}{k}+2\delta_{k}.$ (3.8)
-Until now, using the specification property we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the remaining cases for $1\leq j<Q_{k}$.
-Since the system has specification property, $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$.
-Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist a point $x_{k,j}$ and $n_{(k,j)}\in\mathbb{N}$ such that
+Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the cases for $1\leq j<Q_{k}$.
+Since the system has specification property, the generic point $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$ [PS07].
+Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
 $D(\mathcal{E}_{n_{(k,j)}}(x_{k,j}),\mu_{(k,j)})<\delta_{(k,j)},$
@@
 Now set $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
-Step $1$. Construction of intermediate sets $\mathcal{S}_{(k,j)}$.
+With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready for the construction:
+Step $1$. Construct the intermediate sets $\mathcal{S}_{(k,j)}$.
 Recall that we have defined $N_{(k,0)}=1$ for any $k\geq 0$. Now choose a strictly increasing sequence of integers $\{N_{(k,j)}\}_{(k,j)\in\mathcal{A},~{}j\neq 0}$ such that $N_{(1,1)}=1$ and
 $\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k},\\ n_{k}N_{(k,Q_{k})}&\geq(1-\zeta)\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}].\end{split}$ (3.9)
-For $(k,j)\in\mathcal{A}$ with $j=0$, set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$.
+For $(k,j)\in\mathcal{A}$ with $j=0$, we set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$.
 We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
 Case $1$. $j\neq Q_{k}$.
-Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
-By specification property, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with gap $m_{(k,j)}$.
-Now set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
+Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$ is defined above.
+By specification property again, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$.
+Now we set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
 Case $2$. $j=Q_{k}$.
@@
 $\mathcal{C}_{(k,Q_{k})}=\{y_{i}^{k}:i=1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}.$
-For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$.
-Then define
-$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}.$
-Write $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$, then it follows from the facts that $\mathcal{C}_{(k,Q_{k})}$ is a $(n_{k},4\epsilon^{\ast})$-separated set, and $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set.
-In particular, by (3.6) and (3.8), one has
+For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$.
+We write
+$\mathcal{S}_{(k,Q_{k})}=\left\{y=y(i_{1},\ldots,i_{N_{(k,Q_{k})}}):(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}\right\}$
+and $c_{k,Q_{k}}=n_{k}\cdot N_{(k,Q_{k})}+(N_{(k,Q_{k})}-1)m_{(k,Q_{k})}$.
+Since $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set, this implies that $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set.
+Then, by (3.1.2) and (3.9), one has the estimation for the cardinality of $\mathcal{S}_{(k,Q_{k})}$:
 #$\#\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.10)
-Similar to the argument in (3.7) one has
-$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ y_{k,j}\in\mathcal{S}_{(k,j)}\ \text{for }k\geq 1\ \text{and }1\leq j<Q_{k},$
+Similar to the arguments in (3.8) one has
+$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ \forall\ y_{k,j}\in\mathcal{S}_{(k,j)}~{}\text{with}~{}k\geq 1\ \text{and }1\leq j<Q_{k},$
 and
 $D(\mathcal{E}_{c_{k,Q_{k}}}(y),\mu^{\ast})<\frac{5}{k}+2\delta_{k},\ \forall\ y\in\mathcal{S}_{(k,Q_{k})}.$
-Hence, in an unified way we always have
+Hence, noting that $\mu_{(k,Q_{k})}=\mu^{*}$, in an unified manner we always have
 $D(\mathcal{E}_{c_{k,j}}(y),\mu_{(k,j)})<\frac{5}{k}+2\delta_{(k,j)},\ \forall\ y\in\mathcal{S}_{(k,j)}\ \text{with }k\geq 1,\ 1\leq j\leq Q_{k}.$ (3.11)
-Step $2$. Constructions of $\mathbb{F}_{(k,j)}$ and Moran set $\mathbb{F}$.
+Step $2$. Construct the $\mathbb{F}_{(k,j)}$ and the Moran set $\mathbb{F}$.
+One may regard this process as an analogue of the construction of the middle three Cantor set of the unit interval $[0,1]$.
 Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$.
-Now suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$.
+Suppose that we have already constructed the set $\mathcal{I}_{(k,j)^{\ast}}$, and then we will describe how to construct $\mathcal{I}_{(k,j)}$.
 Let
@@
 Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in formula (3.10) has the estimation:
-#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$
-For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by specification property let $z=z(x,y)\in X$ be some point such that
+#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)
+For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by the specification property let $z=z(x,y)\in X$ be some point such that
 $d_{M_{(k,j)^{\ast}}}(x,z)<\epsilon_{(k,j)},\ d_{c_{k,j}}(y,f^{M_{(k,j)^{\ast}}+m_{(k,j)}}(z))<\epsilon_{(k,j)}.$
 We set $\mathcal{I}_{(k,j)}=\{z(x,y):x\in\mathcal{I}_{(k,j)^{\ast}},y\in\mathcal{S}_{(k,j)}\}$.
-For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
-$d_{M_{k,Q_{k}}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
+For $x\in\mathcal{I}_{(k,Q_{k})^{\ast}}$ and two distinct $y,y^{\prime}\in\mathcal{S}_{(k,Q_{k})}$, since $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set, one has
+$d_{M_{(k,Q_{k})}}(z(x,y),z(x,y^{\prime}))>2\epsilon^{\ast},$
 which leads to
-#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$
+#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
 Moreover, according to (3.11), for $z\in\mathcal{I}_{(k,j)}$ with $j\neq 0$, we have
@@
 $\mathbb{F}_{(k,j)}=\bigcup_{x\in\mathcal{I}_{(k,j)}}\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)}).$
-It is easy to see that for each $(k,j)\in\mathcal{A}$ the following is satisfied:
+Then for each $(k,j)\in\mathcal{A}$, the following statements are satisfied:
 $(1)$ $\overline{B}_{M_{(k,j)}}(x,2\epsilon_{(k,j)})\cap\overline{B}_{M_{(k,j)}}(x^{\prime},2\epsilon_{(k,j)})=\emptyset$ holds for any distinct $x,x^{\prime}\in\mathcal{I}_{(k,j)}$;
-$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has $\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}})$.
-That is $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$.
-Let
+$(2)$ if $z\in\mathcal{I}_{(k,j)}$ descends from $x\in\mathcal{I}_{(k,j)^{\ast}}$, one has
+$\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}}).$
+This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$. Let
 $\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}=\bigcap_{k\in\mathbb{N}}\mathbb{F}_{(k,Q_{k})}.$
 Lemma 3.4.
-One has the fractal set $\mathbb{F}$ is contained in $\mathbb{F}\subset G_{K}^{\omega,U}$, that is,
+The fractal set $\mathbb{F}$ is contained in $G_{K}^{\omega,U}$, i.e.,
 $\mathbb{F}\subset G_{K}^{\omega,U}=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\}.$
-Proof.
-Firstly, we know $\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U$.
+Proof of Lemma 3.4.
+By the construction of $\mathbb{F}$, we know
+$\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset\overline{B(a_{0},\tilde{\epsilon})}\subset U.$
 For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$.
-Hence for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there is a $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$, which implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$.
-By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have $\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}$.
-For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, in order to estimate $D(\mathcal{E}_{n}(p),K)$, we split into four cases to discuss.
+Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$.
+This implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$.
+By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have
+$\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}.$
+For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
 Case $(1)$. There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
@@
 $D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,Q_{k})}.$
-Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$, such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
+Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
 Then there is some $z\in\mathcal{I}_{(k,j)}$ such that $p\in\overline{B}_{M_{(k,j)}}(z,\frac{\epsilon_{(k,j)}}{2})$, and
@@
 $D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,j)}.$
-Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$, such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
+Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
 One has
@@
 $D(\mathcal{E}_{n}(p),K)\leq\frac{16}{k-1}+4\delta_{(k-1,0)}.$
-Case $(4)$. There exists $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$, such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
+Case $(4)$. There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
 In this case, using the expression
@@
 To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$.
-Based on above arguments, one can also show sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points.
+Based on the above arguments, one can also show that the sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points.
 Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$.
@@
 ∎
-Step $3$. Construction of measure.
-Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying that the $\rho$-measure of some appropriate dynamical balls decay exponentially fast along sequence $\{M_{k}\}$, where $M_{k}=M_{k,Q_{k}}$.
-Then we can use the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
+Step $3$. Construct the measure $\rho$ supported on $\mathbb{F}$.
+Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying the $\rho$-measure of some appropriate dynamical balls decays exponentially fast along the sequence $\{M_{k}\}$, where $M_{k}:=M_{(k,Q_{k})}$.
+Then we can invoke the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
 Define
 $\rho_{k}=\frac{1}{\#\mathcal{I}_{(k,Q_{k})}}\sum_{x\in\mathcal{I}_{(k,Q_{k})}}\delta_{x}.$
-For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$, respectively.
+For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ respectively.
 Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$.
 It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$.
-Thus $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$, which implies that $\rho(\mathbb{F})=1$.
-By (3.9), we have
-#$\#\mathcal{I}_{k}\geq\exp\left[M_{k}\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
+Thus one has $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$.
+This implies that $\rho(\mathbb{F})=1$.
+Recall that $M_{(0,0)}=n_{(0,0)}=1$, by (3.12) and (3.13), we have
+#$\#\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
 Lemma 3.5.
-Let $k_{0}\geq 1$, and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball which intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
+Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball that intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
 $\rho_{k}\left(B_{M_{k_{0}}}\left(z,\frac{\epsilon^{\ast}}{2}\right)\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}.$
-Proof.
-If $k=k_{0}+1$, it is easy to check that
+Proof of Lemma 3.5.
+Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not=\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also nonempty by the construction of $\mathbb{F}$ and #$\#(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$.
+Notice that $\mathbb{F}_{k_{0}+1}$ descends from $\mathbb{F}_{k_{0}}$.
+Then one has
 #$\#(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
@@
 In general, for any $k>k_{0}$,
-$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
-For large enough $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
+$\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
+For sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
 $\rho\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)$ $\leq\liminf_{k\to\infty}\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}$
-$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
+$\leq\exp\left[-(M_{k_{0}}-1)(1-\zeta)^{2}\cdot(H-3\zeta)\right]$
+$\leq\exp\left[-M_{k_{0}}(1-\zeta)^{3}\cdot(H-3\zeta)\right]$
 Using Lemma 2.2, one has
-$(1-\zeta)^{2}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
+$(1-\zeta)^{3}\cdot(H-3\zeta)\leq h_{top}^{P}(f,\mathbb{F},\frac{\epsilon^{\ast}}{12})\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
 The arbitrariness of $\zeta$ implies that
 $\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$
-Now dividing $|\log\epsilon^{\ast}|$ in both sides and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
+Dividing in both sides of the above inequality by $|\log\epsilon^{\ast}|$, and then taking $\limsup_{\epsilon^{\ast}\to 0}$, we have
 $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d).$ (3.15)
 Finally, by (3.2) and (3.15) we get
-$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+$\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+which proves Theorem 1.1.
 3.2 Bowen metric mean dimension of saturated sets
-We prove Theorems 1.2 and 1.3 in this subsection.
+In this subsection, we prove Theorems 1.2 and 1.3.
 3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
-The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already given in [Y23, Lemma 3.3,(3)]
+The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already obtained in [Y23, Lemma 3.3,(3)].
 Lemma 3.6.
-Let $(X,f)$ be a t.d.s.. Let $K\subset M_{f}(X)$ be a nonempty closed connected subset. Then for any $\epsilon>0$, we have
+Let $(X,f)$ be a t.d.s., and let $K\subset M_{f}(X)$ be a nonempty closed subset.
+  • (1)
+    For any $\epsilon>0$, we have
+    $h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap K\not=\emptyset\},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+  • (2)
+    For any $\epsilon>0$, we have
     $h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Thus, by the Lemma 3.6 we obtain an upper bound of the Bowen metric mean dimension of saturated set:
+Thus, by the Lemma 3.6 the upper bound of the Bowen metric mean dimension of saturated set is given by:
 $\overline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.16)
 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
-The proof of the lower bound is similar to that of in subsection 3.1.2. So we only give a sketch for the estimation of lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
-We shall show that for any small enough $\epsilon^{\ast}>0$, it holds that
+The proof for the lower bound is similar to that in subsection 3.1.2. So we only give a sketch for the estimation of the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
+We show that for any small enough $\epsilon^{\ast}>0$, it holds that
 $\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$ (3.17)
@@
 $K\subset\cup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})<\frac{1}{k},\ D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})<\frac{1}{k}.$
-Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow
-$\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$
-with the lexicographic order again.
+Put $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, $\mathcal{A}=\{(k,0):k\geq 0\}\cup\mathcal{A}^{\prime}$, and endow these sets with the lexicographic order again.
 Using the fixed notation in subsection 3.1.2, we have the data $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ and $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$.
@@
 $\frac{n_{(k,i)}^{\prime}}{n_{(k,i)}+m_{(k,i)}}>\max\{1-\zeta,1-\frac{1}{k}\},$
-where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$.
-$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$ which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
+where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$;
+$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
 #$\#E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
@@
 […]
-Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
+with gap $m_{(k,i)}$. Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
 #$\#\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
@@
 Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$.
-For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{(k,i,2)},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
+For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
 $\{y_{k,i,l},f(y_{k,i,l}),\ldots,f^{n_{(k,i)}-1}(y_{k,i,l})\},\ l=1,\cdots,N_{(k,i)},$
 with the same gap $m_{(k,i)}$.
-Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points, which is $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
+Denote by $\mathcal{S}_{(k,i)}$ the set of such shadowing points.
+It is a $(c_{k,i},3\epsilon^{\ast})$-separated set, where $c_{k,i}=n_{(k,i)}N_{(k,i)}+(N_{(k,i)}-1)m_{(k,i)}$.
 Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$.
@@
 Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$.
-Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with $\mathbb{F}$, it holds that
+Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with the fractal set $\mathbb{F}$, it holds that
 $\rho(B_{n}(z,\frac{\epsilon^{\ast}}{2}))\leq\exp\left[-n(1-\zeta)^{2}(H-2\zeta)\right].$
-Then by the Lemma 2.1, we have
+Then, by the Lemma 2.1, we have
 $(1-\zeta)^{2}(H-2\zeta)\leq h_{top}^{B}(f,\mathbb{F},\frac{\epsilon^{\ast}}{2})\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$
@@
 Therefore, by (3.16) and (3.17) we get
-$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$ (3.15)
+$\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}^{B}(f,G_{K},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}\xi<\epsilon}h_{\mu}(f,\xi).$
 3.2.3 Proof of Theorem 1.3
+With the help of Theorems 1.1 and 1.2, we prove Theorem 1.3.
 Firstly, we show
@@
 Let $\gamma>0$.
-Then according to Theorem 1.2 we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for any $k$,
+Then, by Theorem 1.2, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
 $\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
-Assume that $\mu_{k}\rightarrow\mu$. Then $\mu\in M_{f}(X)$ since $K$ is closed. This shows
-$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
+Assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$. Then $\mu\in K$ since $K$ is closed. This shows that
+$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
 Letting $\gamma\to 0$, we have
@@
 This completes the proof of (3.18).
-Using formula (3.2), the inequality
-$\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d).$
-is clear.
-On the other hand, let $\gamma>0$ be sufficiently small and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}$.Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence of $\{\mu_{k}\}$ in $K$ such that for sufficiently large $k$,
+Next we show the second equality:
+$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
+Using formula (3.3), the inequality $\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is clear.
+On the other hand, let $\gamma>0$ be sufficiently small and put
+$M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$
+Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
 $M_{\gamma}<\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi).$
-Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows
-$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
+Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows that
+$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}.$
 Letting $\gamma\to 0$, we have
@@
 3.3 Upper capacity metric mean dimension of saturated sets
-We prove the Theorem 1.4 in this subsection.
-The inequality
-$\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)\leq\overline{\rm mdim}_{M}(f,X,d)$
+In this subsection, we prove the Theorem 1.4.
+By Proposition 2.1, the inequality
+$\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K},d)\leq\overline{\rm mdim}_{M}(f,X,d)$
 is clear. Thus, it suffices to show
-$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}\left(f,G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},d\right).$
-Once we prove that for $\epsilon^{\ast}>0$, $\zeta>0$ small enough and $\delta\in(0,1)$, it holds that
-$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}),$
-then by Theorem 2.4, the proof of Theorem 1.4 will be done.
-Since the proof is similar to the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ in subsection 3.1.2, we give a sketch for it.
-We write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
+$\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d).$
+In light of Theorem 2.4, we need to show that for $\epsilon^{\ast}>0$, $\zeta>0$ sufficiently small and $\delta\in(0,1)$, one has
+$\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}).$ (3.19)
+However, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
+Write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)$ and assume that $H>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
 $\underline{h}_{\mu^{\ast}}^{K}(f,3\epsilon^{\ast},\delta)>H-\zeta.$
-Then for large enough $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
+Then, for sufficiently large $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
 #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
 Now fix such $\mathcal{N}$, and write $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
-Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}\subset K$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
+Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
 $K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{k,i},\frac{1}{k}),\ D(\mu_{k,i},\mu_{k,i+1})<\frac{1}{k},\ \text{and }D(\mu_{k,Q_{k}},\mu_{k+1,1})<\frac{1}{k}.$
-Consider the sets $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$, and
-$\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$
-with the lexicographic order.
-The sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed similarly as in subsection 3.1.2.
+Consider the set $\mathcal{A}^{\prime}=\{(k,i):k\geq 1,1\leq i\leq Q_{k}\}$ and endow $\mathcal{A}=\{(k,0):k\in\mathbb{Z}\ \text{and }k\geq-1\}\cup\mathcal{A}^{\prime}$ with the lexicographic order.
+Similarly, the sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed as in subsection 3.1.2.
 For the nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$.
@@
 $d_{M_{(0,0)}}(x,x^{\prime})>2\epsilon^{\ast}.$
-Recall that $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$.
-Hence for $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$, and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$ one has
+Recall $M_{(0,0)}=1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}$.
+Let $y$, $y^{\prime}\in\mathbb{F}$ with $y\in\overline{B}_{M_{(0,0)}}(x,2\epsilon_{(0,0)})$ and $y^{\prime}\in\overline{B}_{M_{(0,0)}}(x^{\prime},2\epsilon_{(0,0)})$.
+Then one has
 $d_{M_{(0,0)}}(y,y^{\prime})>d_{M_{(0,0)}}(x,x^{\prime})-d_{M_{(0,0)}}(x,y)-d_{M_{(0,0)}}(x^{\prime},y^{\prime})>\epsilon^{\ast}.$
-Hence for any large enough $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$.
+Hence, for sufficiently large $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$.
 Thus
 $h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
-4 Applications
-In this section, we apply the main results to multifractal analysis and mean Li-Yorke chaos. The main results are Theorems 4.1 and 4.3.
+This implies the inequality (3.19).
+3.4 An extension to metric mean dimensions of sup sets and sub sets
+In this section, we consider two variants of saturated sets, namely sup sets and sub sets, and establish variational principles for metric mean dimensions of sup sets and sub sets.
+The main results are presented in Theorems 3.7, 3.8 and 3.9.
+In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the sense of weak-topology.
+The set of such points is denoted by $I(f)$, known as the historic set [^1] This terminology is due to Rulle’s original paper [Rue01].
+of $f$.
+By Birkhoff ergodic theorem, although the historic set is usually regarded as negligible in the measure-theoretic context, its topological dynamics can be quite rich.
+Especially for systems with specification-like properties, it can even have full topological entropy [CTS05] and full metric mean dimension [LV21, LL24].
+To understand the underlying geometry structure of $I(f)$, we define the following sets based on the framework introduced and further developed by Olsen [O02, O03, O06, O08] and Olsen and Winter [OW03].
+More precisely, given a nonempty subset $C\subset M_{f}(X)$, we define the sup set, equ set [^2] In this context, we do not take into account the deformation of $\mathcal{E}_{n}(x)$ via a continuous affine map from $M_{f}(X)$ to a topological vector space $Y$.
+Hence, we prefer to use the term “saturated set” rather than “equ set”.
+and sub set of $C$ as
+${}^{C}G:$ =$=\{x\in X:V_{f}(x)\supset C\},$
+$G_{C}:$ =$=\{x\in X:V_{f}(x)=C\},$
+$G^{C}:$ =$=\{x\in X:V_{f}(x)\subset C\},$
+respectively.
+Note that the equ set is precisely the saturated set associated with $C$ defined previously. More generally, by using two nonempty subsets $C_{1},C_{2}\subset M_{f}(X)$, we define the set
+$G(C_{1},C_{2}):=\{x\in X:C_{1}\subset V_{f}(x)\subset C_{2}\},$
+which covers the three different notions above in an unified manner.
+Now, we are ready to establish variational principles for the three given types of level sets $G^{C},^{C}G,G(C_{1},C_{2})$ by means of the Bowen metric mean dimension.
+Theorem 3.7.
+Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty subset $C\subseteq M_{f}(X)$,
+$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Proof.
+We claim that for every $\epsilon>0$,
+$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
+where $\overline{{\rm Conv}}(C)$ denotes the convex closure of set $C$.
+Indeed, since $C\subset\overline{{\rm Conv}}(C)$, it suffices to show
+$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Let $\mu\in\overline{{\rm Conv}}(C)$.
+We can choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $\lambda_{n}^{(j)}>0$ and $\sum_{j=o}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$.
+The upper semi-continuity of $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ on $M_{f}(X)$ (cf.
+[FW16, Lemma 2.3, (3)]) implies that
+$\limsup_{n\to\infty}~{}{\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{n}}(f,\xi)}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Combining this and $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is concave in $\mu$, we can deduce that
+$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma$
+holds for any for $\gamma>0$. This yields the desired inequality and hence (3.20) holds.
+Notice that $G_{\overline{{\rm Conv}}(C)}\subset^{C}G$, we have
+$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C)},d\right)$
+[…]
+$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Conversely, for any $\mu\in C$ and $\epsilon>0$, by Lemma 3.6 one has
+$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+This shows $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ for every $\epsilon>0$. ∎
+Theorem 3.8.
+Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty closed subset $C\subseteq M_{f}(X)$, we have
+$\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Proof.
+For every $\epsilon>0$, by Lemma 3.6 again we have
+$h_{top}^{B}(f,G^{C},\epsilon)\leq h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap{C}\not=\emptyset\},\epsilon)\leq\sup_{\mu\in{C}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+By (3.17), there exists a constant $c>0$ such that
+$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon)$
+holds for all $\mu\in C$ and any sufficiently small $\epsilon>0$. Combining with the fact that $G_{\mu}\subset G^{C}$ holds for each $\mu\in C$, one gets
+$\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G^{C},\epsilon).$
+Theorem 3.9.
+Let $(X,d,f)$ be a t.d.s.
+with specification property.
+Suppose that $C_{1}\subseteq M_{f}(X)$ is a nonempty set satisfying that $\overline{{\rm Conv}}(C_{1})$ is a connected component of $C_{2}\subset M_{f}(X)$.
+Then
+$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Proof.
+Since $C_{1}\subset\overline{{\rm Conv}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
+$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C_{1})},d\right)$
+$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm Conv}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
+$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Furthermore, we have $\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,^{C_{1}}G,d\right)$. Using Theorem 3.7, we get
+$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+This completes the proof. ∎
+The Theorems 3.7, 3.8 and 3.9 imply the following corollaries, which allows us to compare the metric mean dimensions of sup sets and sub sets with the saturated sets.
+Corollary 3.10.
+Let $(X,d,f)$ be a t.d.s. with specification property.
+  • (1)
+    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
+    $\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right).$
+  • (2)
+    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
+    $\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right).$
+Proof.
+(1) follows by Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
+Corollary 3.11.
+Let $(X,d,f)$ be a t.d.s. with specification property.
+  • (1)
+    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
+    $\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right).$
+  • (2)
+    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
+    $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,^{C}G,d\right).$
+  • (3)
+    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
+    $\overline{\rm mdim}_{M}^{UC}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,^{C}G,d\right).$
+Proof.
+$(1)$ is clear. For $(2)$, it follows from Theorem 1.1 that
+$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right).$
+On the other hand, one has $h_{top}^{P}(f,G^{C},\epsilon)\leq\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ by Lemma 3.2.
+Hence, we have $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)$.
+$(3)$ follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
+4 Some applications of main results
+In this section, we apply the main results to multifractal analysis of level sets and mean Li-Yorke chaos in infinite entropy systems.
+The main results in this section are presented in Theorems 4.1 and 4.3.
 4.1 Variational principle of level sets
-In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ in the sense of an abstract framework.
-Let $(X,f)$ be a dynamical system, and let $\varphi:M_{f}(X)\to\mathbb{R}$ be a continuous function. Let
-$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right],\$
-For $\alpha\in L$, let $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$, and let
-$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\}$
-denote the level set of $\alpha$ w.r.t.
-$\varphi$ consisting of the set of points of $X$ such that $\varphi$ is taking constant value $\alpha$ on $V_{f}(x)$.
-Then the continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
-Condition A: Given a continuous function $\varphi$ on $M_{f}(X)$, for every $\alpha\in L_{\varphi}$ we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
-Some examples of $\varphi$ satisfying Condition A are given as follows:
+In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ within the context of an abstract framework.
+Given a dynamical system $(X,f)$, assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function. We define
+$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right].$
+For any $\alpha\in L_{\varphi}$, we define $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$.
+The continuity of $\varphi$ then implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
+We define the level set of $\alpha$ w.r.t.
+$\varphi$ as
+$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\},$
+which consists of the set of points of $X$ such that $\varphi$ takes constant value $\alpha$ on $V_{f}(x)$.
+Condition A: For every $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
+Several examples of the function $\varphi$ that satisfy Condition A are stated below:
   • (a)
@@
   • (b)
-    Let $\psi$, $\phi\in C(X)$ such that $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ holds for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$). We define
-    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu}.$
+    We define
+    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu},$
+    where $\psi$, $\phi\in C(X)$ and $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$).
     Then for any $\alpha\in L_{\varphi}$,
@@
   • (c)
-    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i.e.
-for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$.
+    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, that is, for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$.
 We define
     $\varphi:M_{f}(X)\to\mathbb{R},~{}\mu\mapsto\lim_{n\to\infty}\int\frac{\phi_{n}(x)}{n}\mathop{}\!\mathrm{d}\mu(x).$
-    Then $\varphi$ is continuous (cf. [FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for any $\Phi$ and $\alpha\in L_{\Phi}$.
-    Specially, when the sequence $\{\phi_{n}\}$ is an additive function $\{S_{n}\psi\}$ with $\psi\in C(X)$, in this case we have
+    Then $\varphi$ is continuous (cf.
+[FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for all $\alpha\in L_{\Phi}$.
+Specially, let $\psi\in C(X)$ and define $\phi_{n}=S_{n}\psi$, then $\Phi=\{\phi_{n}\}$ is an additive potential, in this case we have
     $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\psi\mathop{}\!\mathrm{d}\mu=\alpha\},$
     $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
-Given a nonempty open subset $U\subset X$, we put
+Given a nonempty open subset $U\subset X$, we similarly define the mixed $\alpha$-level set
 $K_{\alpha}^{\omega,U}=K_{\alpha}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}.$
 Theorem 4.1.
-Let $(X,d,f)$ be a dynamical system with specification property and $U\subset X$ be a nonempty open subset.
-Suppose that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A.
-Then for any $\alpha\in L$ and $S\in\{B,P\}$, we have
+Let $(X,d,f)$ be a t.d.s.
+with the specification property, and let $U\subset X$ be a nonempty open subset.
+Assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A.
+Then for any $\alpha\in L_{\varphi}$ and $S\in\{B,P\}$,
 $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha}^{\omega,U},d)=$ $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha},d)$
@@
 Proof.
-Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$, which is not dependent on $\mu$, such that
-$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$ (4.1)
-holds for all $\mu\in M_{f}(X)$ and any small enough $\epsilon>0$. Together with the fact $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
+Define $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.17), there exists a constant $c>0$ such that
+$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$
+holds for all $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$. Together with the fact that $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
 $\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha}^{\omega,U},\epsilon)$
-for all sufficiently small $\epsilon>0$, which implies that
+for all sufficiently small $\epsilon>0$. This implies that
 $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha}^{\omega,U},d).$ (4.1)
 Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$.
-Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, then $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$.
+Fix $x\in X$.
+Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, it follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$.
 By Lemma 3.2 $(1)$, we get
@@
 By (4.1), (4.2), and the fact stated in Proposition 2.1 $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)$ for any $Z\subset X$, we complete the proof. ∎
-A direct consequence of Theorem 4.1 is the variational principles for Bowen metric mean dimension of level sets obtained by Backes and Rodrigues [BR23, Theorem A], but we can strength the result which states that the variational principle is still valid for packing metric mean dimension of level sets.
+A direct consequence of Theorem 4.1 is the variational principles for the Bowen metric mean dimension of level sets, as derived by Backes and Rodrigues [BR23, Theorem A].
+However, we are able to strengthen this result by demonstrating that the variational principle remains valid for the packing metric mean dimension of level sets [^3] Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure..
 Corollary 4.2.
-Let $(X,d,f)$ be a dynamical system with specification property and $\psi$ be a continuous function on $X$. Let
+Let $(X,d,f)$ be a t.d.s. with the specification property and $\psi$ be a continuous function on $X$. Let
 $E_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}$
 denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then
-$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
 […]
@@
 Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$.
-We get the desired result by Theorem 4.1.
+We obtain the desired result by Theorem 4.1.
 ∎
-Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure.
-Then, by Theorem 4.1, one can get analogous variational principles for the metric mean dimensions of level sets of some special functions, which are determined by the dynamical behaviors of the system.
 4.2 Linking metric mean dimension and mean Li-Yorke chaotic
-In this subsection, we link the concepts of metric mean dimension and mean Li-Yorke chaos in the sense of infinite entropy.
-We give a quantitative characterization for the chaotic phenomenon of infinite entropy systems.
-Namely, we show the packing metric mean dimension of the set of mean Li-Yorke pairs is bounded from blow by the metric mean dimension of the phase space (Theorem 4.3).
-Let $(X,d,f)$ be a dynamical system.
-This naturally gives a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is compact endowed with the product metric $d\times d$ defined by $d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\}$, and the map
-$f\times f:X\times X\rightarrow X\times X$
-given by $(f\times f)(x,y):=(fx,fy)$ is continuous.
-It is well-known that if two systems admit the specification property, then the product system also has the specification property(cf.
-[DGS76, Proposition 21.4]).
+In the present subsection, we establish a connection between the concepts of metric mean dimension and mean Li-Yorke chaos within the context of infinite entropy.
+We provide a quantitative description of the chaotic behavior exhibited by infinite entropy systems.
+Specifically, we demonstrate that the packing and upper capacity metric mean dimensions of the set of mean Li-Yorke pairs are equal to twice the metric mean dimension of the phase space, as stated in Theorem 4.3.
+Given a t.d.s. $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric defined as
+$d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\},$
+and the mapping $f\times f:X\times X\rightarrow X\times X$, given by $(f\times f)(x,y):=(fx,fy)$, is continuous.
 A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
-$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
-The set of mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$.
+$\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,~{}\text{and}\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
+The set of all mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$.
 A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair.
 We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set.
+The well-known fact is that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]).
 Theorem 4.3.
-Let $(X,d,f)$ be a dynamical system with specification property. Then
-$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d),$
-$\overline{\rm mdim}_{M}^{UC}(f\times f,{\rm MLY}(f),d\times d)=\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d).$
+Let $(X,d,f)$ be a t.d.s. with specification property. Then for $S\in\{UC,P\}$ we have
+$\overline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$
+$\underline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~{}\underline{\rm mdim}_{M}(f,X,d).$
 Proof.
-Claim: there exists $\mu_{0}\in M_{f\times f}(X\times X)$ with $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
-Proof of the claim.
-Specification property implies that the system has positive entropy.
-Then by the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$.
-We put $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$.
-Then any an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
-∎
-Now take $\nu\in M_{f}(X)$ and put $\mu_{1}=\phi_{\ast}\nu\in M_{f\times f}(X\times X)$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$.
-Then one has $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$.
-Assume that $\overline{\rm mdim}_{M}(f,X,d)$ is finite.
-For any $\gamma>0$, by Theorem 2.4 we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $\rho_{k}\in M_{f}(X)$ such that for each $k$,
-$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$ (4.4)
-Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then for every $k$,
-$\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)=\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta).$ (4.5)
-Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ so that ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
-$H_{\rho_{k}^{\ast}}(\eta^{n})$ $=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=-\sum_{C\in\xi^{n}}\rho_{k}^{\ast}(C\times C)\log\rho_{k}^{\ast}(C\times C)$
-$=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=H_{\rho_{k}}(\xi^{n}),$
-which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\leq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$.
-Similarly, if $\eta$ is a finite Borel partition of $X\times X$, then $\xi=\{\phi^{-1}(D\cap\Delta):D\in\eta\}$ is a finite partition of $\Delta=\{(x,x):x\in X\}$, such that ${\rm diam}(\xi)\leq{\rm diam}(\eta)$.
-Moreover, one has
-$H_{\rho_{k}}(\xi^{n})=-\sum_{C\in\xi^{n}}\rho_{k}(C)\log\rho_{k}(C)=-\sum_{\tilde{C}\in\eta^{n}}\rho_{k}^{\ast}(\tilde{C})\log\rho_{k}^{\ast}(\tilde{C})=H_{\rho_{k}^{\ast}}(\eta^{n}),$
-which implies that $\inf_{{\rm diam}(\eta,d\times d)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)\geq\inf_{{\rm diam}(\xi,d)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)$. This shows (4.5).
-Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e.
-the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$.
+It suffices to show the first equality since the second one is obtained similarly.
+We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
+Notice that the specification property implies that the system has positive topology entropy.
+Since positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that
+$\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$
+Set $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$.
+Then, in the weak-topology any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
+Now take $\nu\in M_{f}(X)$ arbitrarily and put $\mu_{1}:=\phi_{\ast}\nu=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$.
+It is easy to check that $\mu_{1}\in M_{f\times f}(X\times X)$, and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$.
+For sufficiently small $\gamma>0$, let
+$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)-\gamma\}.$
+By Theorem 2.4, we pick a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and a sequence of invariant measures $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
+$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$ (4.3)
+Let $K=\overline{{\rm Conv}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}$, i.e., the convex closure of the set $\{\mu_{0},\mu_{1}\}\cup\{m_{k}:k\geq 1\}$.
 Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$.
-For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$.
-Hence
+It is easy to see that if two systems admit the specification property, so does the product system.
+By [PS07], we have $G_{K}\not=\emptyset$.
+For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}_{n\in\mathbb{N}}$ and $\{t_{n}^{1}\}_{n\in\mathbb{N}}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively.
+Hence,
 $\lim_{n\to\infty}\frac{1}{t_{n}^{0}}\sum_{i=0}^{t_{n}^{0}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}>0,$
@@
 $\lim_{n\to\infty}\frac{1}{t_{n}^{1}}\sum_{i=0}^{t_{n}^{1}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0.$
-This shows $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
+This implies that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
 $\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq$ $\overline{\rm mdim}_{M}^{P}(f\times f,G_{K},d\times d)$
-$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
-$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
-$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}\geq\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
-Since $\gamma$ is arbitrary, by Proposition 2.1 one has
-$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
-The case of $\overline{\rm mdim}_{M}(f,X,d)=\infty$ is still valid by considering a slight modification.
-For sufficiently small $\gamma>0$, let $c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}$.
-By Theorem 2.4, we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
-$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f,\xi)>c_{\gamma}.$ (4.6)
-Put $\hat{K}=\overline{{\rm Conv}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}}$. Then the similar argument gives us
-$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d),$
-which implies the desired result by Proposition 2.1.
-Finally, we give an example to verify the Corollary 4.2 and Theorem 4.3.
+$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
+$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)}{|\log\epsilon_{k}|}\geq c_{\gamma}.$
+Since $\gamma$ is arbitrary, one has
+$\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
+$\leq$ $\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)$
+$\leq$ $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d),$
+where we used the Proposition 2.1 for the second inequality.
+Together with $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$, we obtain the desired equality.
+∎
+Finally, we give an example to verify these results presented in Corollary 4.2 and Theorem 4.3.
 Example 4.4.
-Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric
+Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric:
 $d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
-and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})\mapsto(x_{n+1})_{n\in\mathbb{Z}}$.
-It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf.
-[DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
+and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})=(x_{n+1})_{n\in\mathbb{Z}}$.
+It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
   • (1)
-    Thus, by the product formula of metric mean dimension one has ${\rm mdim}_{M}(\sigma\times\sigma,[0,1]^{\mathbb{Z}}\times[0,1]^{\mathbb{Z}},d\times d)=2$. Consequently, by Theorem 4.3, we have
-    $\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
+    Thus, by Theorem 4.3, we have
+    $\overline{\rm mdim}_{M}^{P}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
   • (2)
-    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, we have
+    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, by Corollary 4.2 we have
     $\overline{\rm mdim}_{M}^{B}(\sigma,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
 […]
-5 Examples
-Applying the main results, we give the specific formulae of the metric mean dimensions of generic points of full shift over compact metric space.
+4.3 Generic points of full shift over compact metric space
+In this subsection, based on the main results, we present some specific formulae for the metric mean dimensions of the set of generic points of full shifts over compact metric spaces.
+We begin by reviewing some definitions and results for our statements. Let $(X,d,f)$ be a t.d.s., $\mu\in M_{f}(X)$, $\epsilon>0$ and $x\in X$. We define the Brin-Katok local entropy at $x$ as
+$\overline{h}_{\mu}^{BK}(x,\epsilon)=\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n},$
+and the Brin-Katok local $\epsilon$-entropy of $\mu$ as
+$\overline{h}_{\mu}^{BK}(f,\epsilon):=\int\overline{h}_{\mu}^{BK}(x,\epsilon)\mathop{}\!\mathrm{d}\mu,$
+respectively. Specially, if $\mu\in M_{f}^{e}(X)$, then $h_{\mu}^{BK}(x,\epsilon)=h_{\mu}^{BK}(f,\epsilon)$ holds for $\mu$-a.e. $x\in X$.
+Example 4.5.
+Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
+$d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
+It is well-known that the system $(K^{\mathbb{Z}},\sigma)$ has the specification property. Moreover, it was proved in [VV17, Theorem 5] and [S22, Theorem 15] that,
+$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d),$ (4.4)
+where $\overline{\rm dim}_{B}(K,d)$ is the upper box dimension of $K$ under the metric $d$.
+Then, by Corollary 1.5, the metric mean dimensions of the generic points of the full shift over compact metric space are fully determined by the boxing dimension of the underlying space and mean Rényi information dimension of invariant measures.
+Theorem 4.6.
+Under the setting of Example 4.5, for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ one has
+$\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d)$
+$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
+Next, we present two examples of full shifts with the maximal metric mean dimension measure.
 Example 4.7.
 Consider the Example 4.4 again.
-Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ be the product measure, which is a $\sigma$-ergodic measure on $[0,1]^{\mathbb{Z}}$.
-Then the ergodicity of $\mu$ enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
-$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in[0,1]^{\mathbb{Z}}.$
-Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12]. Together with [YCZ24, Theorem 1.1], we have
+Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$, and define $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ as the product measure on $[0,1]^{\mathbb{Z}}$.
+Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\sigma,\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12].
+Together with [YCZ25, Theorem 1.1], we have
 $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d).$
@@
 ${\rm\overline{mdim}}_{M}^{B}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{P}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{UC}(\sigma,G_{\mu},d)=1.$
-We extend the Example 5.1 to more general full shift.
 Example 4.8.
-Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
-$d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
-For $n\geq 0$, let $\pi_{n}$ be the projection from $K^{\mathbb{Z}}$ to its first $n$ coordinates.
-It was proved in [S22, Theorem 15]
-$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K),$
-where $\overline{\rm dim}_{B}(K)$ is the upper box dimension of $K$ under the metric $d$.
+Endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$, and let $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ $1\leq i\leq l$ be contracting similitudes, that is,
+$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}$
+for some $r_{i}\in(0,1)$.
+One says that the iterated function system $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there is a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset,\ \text{for }i\neq j.$
+Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure such that
+  • (1)
+    $K=\cup_{i=1}^{l}T_{i}(K)$;
+  • (2)
+    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, where $\alpha$ is the unique solution of equation
+    $\sum_{i=1}^{l}r_{i}^{\alpha}=1,$
+    and ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. $\alpha$ is called the similarity dimension of the IFS;
+  • (3)
+    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$;
+  • (4)
+    ${\rm supp}(\mu)=K$.
+Many famous fractal sets, for instance, the middle thirds Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition.
+Now take the attractor $K$ obtained above as in the Example 4.5. Then the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$ is invariant with respect to the left shift $\sigma$.
 Theorem 4.9.
-Let $K$ be a compact metric space with a metric $d$. Then there exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
-$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K)$
-$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
+Under the setting of Example 4.8, $\nu$ is the maximal metric mean dimension such that the three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
+$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha,$
+where $S\in\{B,P,UC\}$
 Proof.
-It is known that the system $(K^{\mathbb{Z}},\sigma)$ has specification property. Let $S_{\epsilon}$ be a maximal $\epsilon$-separated subset of $K$ with respect to metric $d$. For $n\geq 1$, we put
-${\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
-and consider the set $\hat{\mathcal{S}}_{n}\subset K^{\mathbb{Z}}$ consisting of one preimage of each element in ${\mathcal{S}_{n}}$ under $\pi_{n}$.
-Then $\hat{\mathcal{S}}_{n}$ is an $(n,\epsilon)$-separated set of $K^{\mathbb{Z}}$ and #$\#{\hat{\mathcal{S}}_{n}}=(\#S_{\epsilon})^{n}$.
-Define
-$\mu_{n}:=\frac{1}{\#\hat{\mathcal{S}}_{n}}\sum_{x\in\hat{\mathcal{S}}_{n}}\mathcal{E}_{n}(x).$
-We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(K^{\mathbb{Z}})$. Then, by Lemma 3.1, for each $\epsilon>0$ we have
-$\log\#S_{\epsilon}=\limsup_{n\to\infty}\frac{1}{n}\log\#\hat{\mathcal{S}}_{n}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(\sigma,\xi).$
-Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$ as $\epsilon\to 0$. Combining with the Theorem 1.3, one has
-$\overline{\rm dim}_{B}(K)=\limsup_{\epsilon\to 0}\frac{\log\#S_{\epsilon}}{|\log\epsilon|}\leq\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})\leq\overline{\rm dim}_{B}(K).$
-Finally, by the Corollary 1.5 we complete the proof.
-Availability of data and materials: Not applicable.
-Competing interests: The authors declare that they have no competing interests.
+Hutchinson [Hut81] showed that for any $z\in K$ and $\epsilon\in(0,1)$, one has
+$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha},$ (4.5)
+where $C>0$ is a constant independent of the choice of the point  $z\in K$ and  $\epsilon$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$ and $\epsilon\in(0,1)$, we define
+[…]
+Then the Bowen ball $B_{n}(x,\epsilon)$ is contained in $L_{n}(x,\epsilon)$ and $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5). Thus, one has
+$\nu(B_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha},\ \forall x\in K^{\mathbb{Z}},$
+which yields that $\overline{h}_{\nu}^{BK}(\sigma,\epsilon)\geq\alpha|\log\epsilon|-\log C.$ By the Shannon-McMillan-Breiman theorem, the following inequality is valid for any t.d.s. $(X,d,f)$:
+$\overline{h}_{\mu}^{BK}(f,\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),\ \forall\mu\in M_{f}(X).$ (4.6)
+Hence
+[…]
+[…]
+which implies the desired result. ∎
 Acknowledgments: The first and second authors were supported by NNSF of China (12101340).
-The third author was supported by the China Postdoctoral Science Foundation (Grant Number: 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University.
+The third author was supported by the China Postdoctoral Science Foundation (No.
+2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University.
 The authors also thank Prof.

### v4

@@
 Abstract.
-Saturated sets and its reduced case, the set of generic points, constitute two significant types of fractal-like sets in multifractal analysis of dynamical systems.
-In the context of infinite entropy systems, we aim to provide some qualitative aspects of saturated sets and the set of generic points from both topological and measure-theoretic perspectives.
-For systems with the specification property, we establish certain variational principles for saturated sets in terms of Bowen and packing metric mean dimensions, and show the upper capacity metric mean dimension of saturated sets has full metric mean dimension.
-As applications, we further present some qualitative aspects of the metric mean dimensions of level sets and the set of mean Li-Yorke pairs in infinite entropy systems.
-Keywords and phrases: Metric mean dimension; Saturated set; Generic point; Variational principle; Multifractal analysis
+From a geometric perspective, we employ metric mean dimension to investigate the set of generic points of invariant measures and saturated sets in infinite entropy systems.
+For systems with the specification property, we establish certain variational principles for the Bowen and packing metric mean dimensions of saturated sets in terms of Kolmogorov-Sinai $\epsilon$-entropy, and prove that the upper capacity metric mean dimension of saturated sets has full metric mean dimension.
+Consequently, the Bowen and packing metric mean dimensions of the set of generic points of invariant measures coincide with the mean Rényi information dimension, and the upper capacity metric mean dimension of the set of generic points of invariant measures also has full metric mean dimension.
+As applications, for systems with the specification property, we present the qualitative characterization of the metric mean dimensions of level sets, the set of mean Li-Yorke pairs in infinite-entropy systems, and the set of generic points of invariant measures in full shifts over compact metric spaces.
+Keywords: multifractal analysis; metric mean dimension; saturated set; generic point; variational principle
 1 Introduction
-Let $(X,d,f)$ be a topological dynamical system (t.d.s.
-for short), where $(X,d)$ is a compact metric space with a metric $d$, and $f:X\to X$ is a homeomorphism.
-Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures endowed with weak-topology, $f$-invariant Borel probability measures, and ergodic $f$-invariant Borel probability measures on $X$, respectively.
-Quantitative analysis using some entropy-like quantities is of great significance in understanding the dynamics of abstract dynamical systems.
-Measure-theoretic entropy and topological entropy are two crucial invariants that helps us understand the topological complexity of dynamical systems, which are related by the classical variational principle:
+Let $(X,d,f)$ be a topological dynamical system (TDS for short), where $(X,d)$ is a compact metric space, and $f:X\to X$ is a homeomorphism.
+Let $M(X)$, $M_{f}(X)$, $M_{f}^{e}(X)$ denote the sets of Borel probability measures on $X$ endowed with the weak-topology, $f$-invariant, and $f$-ergodic Borel probability measures on $X$, respectively.
+Quantitative analysis of dynamical systems aims to use entropy-like quantities to understand the dynamics of the abstract dynamical systems.
+Among these entropy-like quantities, measure-theoretic entropy and topological entropy are two topological invariants to characterize the topological complexity of dynamical systems, which are related by the classical variational principle:
 $h_{top}(f)=\sup_{\mu\in M_{f}(X)}{h_{\mu}(f)},$
-where $h_{top}(f)$ is the topological entropy of $X$, $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$, and the supremum ranges over all $f$-invariant measures of $X$.
-In 1973, resembling the definition of Hausdorff dimension Bowen [Bow73] introduced Bowen topological entropy for any subset of the phase space $X$.
-The pioneering work has directly contributed to the development of dimension theory in dynamical systems, especially serving as a powerful tool for studying the “size” of fractal-like sets in multifractal analysis.
-The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors(or known as multifractal spectrum), and then, from a geometrical or topological perspective, describes the “size” of those subsets by using fractal-like dimensions and dimensional-like entropies.
-Such a procedure allows us to get more information about the dynamics and even recovers the partial dynamical behaviors of the system, a phenomenon known as multifractal rigidity.
-We aim to study the multifractal analysis of saturated sets in infinite entropy systems.
-Let us mention some relevant results about the dimensional entropies of saturated sets to clarify the motivation.
-Given $x\in X$, by
-$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}\in M(X),$
-where $\delta_{x}$ is the Dirac mass at $x$, we denote the $n$-th empirical measure of $x$.
-Then, in the sense of weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a nonempty compact connected subset of $M_{f}(X)$ (cf.
-[DGS76, Proposition 3.8]).
-Specially, when $V_{f}(x)=\{\mu\}$ for some invariant measure $\mu$, then $x$ is called a generic point of $\mu$.
+where $h_{top}(f)$ denotes the topological entropy of $X$, and $h_{\mu}(f)$ is the measure-theoretic entropy of $\mu$.
+To extend the concept of topological entropy to any non-empty subsets, resembling the definition of Hausdorff dimension, Bowen [Bow73] in his pioneering work introduced Bowen topological entropy on subsets.
+It turns out that the Bowen topological entropy provides a powerful tool for studying the “size” of fractal-like sets and has directly contributed to the development of multifractal analysis in dynamical systems.
+The multifractal analysis, with its origin dating back to Besicovitch, takes into account the decomposition of the whole phase space into smaller subsets consisting of points with the similar dynamical behaviors.
+And then, from a geometrical and topological perspectives, we use the fractal dimension and entropy-like quantity to describe the “size” of these subsets.
+Such a procedure allows us to obtain more information about the dynamics and even recovers the partial dynamics.
+This phenomenon is also known as multifractal rigidity.
+Typically, infinite entropy systems exhibit rather complicated topological complexity.
+As a result, the concept of topological entropy fails to offer more information about the dynamics of the system.
+In 2000, Lindenstrauss and Weiss [LW00] introduced the metric mean dimension, which is a dynamical analogue of box dimension.
+The measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21].
+The analogous variational principles for metric mean dimension can be found in the references [LT18, GS21, YCZ25, S22].
+Given $x\in X$, denote by
+$\mathcal{E}_{n}(x):=\frac{1}{n}\sum_{i=0}^{n-1}\delta_{f^{i}x}$
+the $n$-th empirical measure of $x$, where $\delta_{x}$ is the Dirac measure at $x$.
+It is well-known that, in the weak-topology, the limit-point set $V_{f}(x)$ of the sequence $\{\mathcal{E}_{n}(x)\}_{n\geq 1}$ is a non-empty compact connected subset of $M_{f}(X)$ .
+In particular, if $V_{f}(x)=\{\mu\}$ for some $f$-invariant measure $\mu$, then $x$ is called a generic point of $\mu$.
 Denote by $G_{\mu}$ the set of all generic points of $\mu$.
-The Birkhoff’s ergodic theorem implies that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
-In 1973, Bowen [Bow73] proved that for any ergodic measure $\mu$, the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
+The Birkhoff ergodic theorem shows that $\mu(G_{\mu})=1$ if and only if $\mu$ is ergodic.
+We aim to study the multifractal analysis of the set of generic points in infinite entropy systems.
+Let us mention some results about the topological entropy of the set of generic points.
+For every ergodic measure $\mu$, Bowen [Bow73] proved that the Bowen topological entropy of $G_{\mu}$ equals $h_{\mu}(f)$, i.e.,
 $h_{top}^{B}(f,G_{\mu})=h_{\mu}(f).$
-For the case of non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there even exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$.
+For the non-ergodic measure $\mu$, one has $\mu(G_{\mu})=0$, and there exist the examples showing that $G_{\mu}=\emptyset$ but $h_{\mu}(f)>0$.
 Therefore, Bowen’s result fails for non-ergodic measures.
-It is natural to ask whether there are some proper conditions such that the Bowen’s equality still holds for non-ergodic measures.
-A positive answer to the above question is due to Pfister and Sullivan [PS07], who provided it within a more general framework.
-A subset $K$ of $X$ is called a saturated set if $x\in K$ and $V_{f}(x)=V_{f}(y)$, then $y\in K$.
-More generally, given a nonempty closed connected subset $K\subset M_{f}(X)$, we have a saturated set by defining $G_{K}=\{x\in X:V_{f}(x)=K\}$ associated with $K$.
-For any topological dynamical system (t.d.s) possessing the specification property, Sigmund [Sig72, Sig74] firstly proved the existence of the saturated sets, and then was later extended to non-uniformly hyperbolic and non-uniformly expanding systems by several authors [LST13, TV17].
-In 2007, Pfister and Sullivan [PS07] introduced the $g$-almost product property that is strictly weaker than specification property, which solely requires a point partially $\epsilon$-shadowing the given finite pieces of orbit-segments and the uniform separation property (e.g.
-expansive and more generally asymptotically $h$-expansive systems), and established the variational principle of Bowen topological entropy of nonempty compact connected set $K\subset M_{f}(X)$:
-$h_{top}^{B}(T,G_{K})=\inf\{h_{\mu}(f):\mu\in K\}.$
-Furthermore, under only the assumption of $g$-almost product property, they [PS07] showed that for all $\mu\in M_{f}(X)$,
-$h_{top}^{B}(f,G_{\mu})=h_{\mu}(f),$
-which extends Bowen’s remarkable result to non-ergodic measures.
-Consequently, the work of [Bow73, PS07] leads to an intensive study toward the estimation of different types of dimensional-like entropies of saturated sets.
-Under the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained the variational principle of packing topological entropy of nonempty compact connected set $K\subset M_{f}(X)$,
-$h_{top}^{P}(f,G_{K})=\sup\{h_{\mu}(f):\mu\in K\}.$
-By weakening the specification property and considering systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] not only demonstrated that the variational principle $h_{top}^{P}(f,G_{K}^{C})=\sup\{h_{\mu}(f):\mu\in K\}$ holds for (mixed) saturated sets with more complicated dynamics, but also showed that the upper capacity topological entropy of the saturated sets has full topological entropy, i.e.,
+In 2007, Pfister and Sullivan [PS07] proved that Bowen’s equality still holds for non-ergodic measures within a more general framework.
+More generally, given a non-empty closed connected subset $K\subset M_{f}(X)$, the saturated set $G_{K}$ associated with $K$ is defined by
+$G_{K}=\{x\in X:V_{f}(x)=K\}.$
+By introducing the $g$-almost product property (e.g.
+$\beta$-shifts), which is strictly weaker than the specification property, and the uniform separation property (e.g.
+expansive and more generally asymptotically $h$-expansive systems), they established a variational principle for $G_{K}$:
+$h_{top}^{B}(T,G_{K})=\inf_{\mu\in K}h_{\mu}(f).$
+Furthermore, under only the assumption of the $g$-almost product property, they [PS07] showed that $h_{top}^{B}(f,G_{\mu})=h_{\mu}(f)$ for all $\mu\in M_{f}(X)$.
+For systems with the specification property and positive expansive property, Zhou, Chen and Cheng [ZCC12] also obtained a (dual) variational principle for packing topological entropy of $G_{K}$:
+$h_{top}^{P}(f,G_{K})=\sup_{\mu\in K}h_{\mu}(f).$
+Later, for systems with the $g$-almost product property, Hou, Tian and Zhang [HTZ23] proved that the upper capacity topological entropy of the saturated sets with more complicated dynamics has full topological entropy, i.e.,
 $h_{top}^{UC}(f,G_{K}^{C})=h_{top}(f),$
 where $G_{K}^{C}:=G_{K}\cap\{x\in X:C_{f}(X)\subset\omega_{f}(x)\},$ $\omega_{f}(x):=\bigcap_{n=0}^{\infty}\overline{\bigcup_{k=n}^{\infty}\{f^{k}x\}}$ is the $\omega$-limit set of $x$, and $C_{f}(X)$ is the measure center given by $C_{f}(X):=\overline{\cup_{\mu\in M_{f}(X)}{\rm supp}{(\mu)}}$, ${\rm supp}(\mu)$ is the supported set of $\mu$.
-Additionally, Huang, Tian and Wang [HTW19] showed that Pfister and Sullivan’s variational principle is still valid for transitive-saturated set $G_{K}^{T}:=G_{K}\cap{\rm Tran}$, where ${\rm Tran}$ means the set of transitive points.
-The generalizations of the aforementioned results can be found in [JCZ22, LLS17] for the scenarios of nonuniformly hyperbolic systems, and in [ZC18, DZZ23, Z18, RTZ23, ZRZ23] for the cases of the actions of amenable groups
-Based on the possible values of topological entropy, dynamical systems can be can be classified into three types: zero entropy systems, finite positive entropy systems and infinite entropy systems.
-Infinite entropy systems typically imply that these systems may exhibit rather complicated topological complexity.
-As a result, the concept of topological entropy fails to offer more information about the dynamics of the system.
-In 2000, Lindenstrauss and Weiss [LW00] introduced a refined notion for topological entropy, called metric mean dimension, which is a dynamical analogue of box dimension to measure how fast the $\epsilon$-topological entropy diverges to $\infty$ as $\epsilon$ goes to 00.
-It has been shown that every system with finite entropy has a metric mean dimension of zero.
-See some applications of metric mean dimension in the analog compression of information theory [GS20], and in estimating the upper bound of mean dimension of some complex systems [Tsu18a, Tsu18b].
-The notion of measure-theoretic counterpart of metric mean dimension, called upper and lower mean Rényi information dimensions, is due to Gutman and Śpiewak[GS21].
-More precisely, the lower and upper mean Rényi information dimensions of $\mu\in M_{f}(X)$ are respectively defined by
-$\underline{\rm MRID}_{\mu}(f,X,d):$ $=\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
-$\overline{\rm MRID}_{\mu}(f,X,d):$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
-where the infimum ranges over the finite Borel measurable partition with diameter at most $\epsilon$.
-This measures how fast the Kolmogorov-Sinai $\epsilon$-entropy $\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ diverges to $\infty$ as $\epsilon\to 0$.
-Several variational principles regarding the metric mean dimension of $\mathbb{Z}$-actions can be found in the works [LW00, GS21, S22, YCZ23a, YCZ25].
 Then, in the context of the infinite entropy systems, the aforementioned work suggests the following questions for saturated sets:
@@
 If not, can metric mean dimensions of saturated sets have the full metric mean dimension?
-For (1), the authors in [YCZ25] showed that the packing metric mean dimension of the set of generic points is equal to the mean Rényi information dimension.
-However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems (cf.[YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07] to the framework of infinite entropy systems encounters some nontrivial challenges.
-Under the assumption of $g$-almost product property, for any non-empty compact connected set $K$ contained in the convex closure of finite invariant measures, Yuan [Y23] investigated the Bowen metric mean dimension of $G_{K}$.
-Prior to the comparison of the different types of specification-like property, we concentrate solely on the systems possessing the specification property and provide a comprehensive answer to question (2).
+For (1), the authors in [YCZ25] proved that the packing metric mean dimension of the set of generic points equals the mean Rényi information dimension.
+However, this equality holds for Bowen metric mean dimension under the assumption of an analogous Brin-Katok formula within the context of infinite entropy systems ([YCZ25, §5, Question,(1)]); for (2), since the uniform separation property implies the system has finite topological entropy, extending the work of [PS07, ZCC12, HTZ23] to the framework of infinite entropy systems encounters some nontrivial challenges.
 Actually, since the asymptotic behavior of orbits ${\rm orb}_{f}(x)=\{f^{n}(x):n\geq 0\}$ is a fundamental topic in ergodic theory, the subsequent statements of main results involve more general (mixed) saturated sets.
@@
 $G_{K}^{\omega,U}:=G_{K}\cap U\cap\{x\in X:\omega_{f}(x)=X\},$
-where $U\subset X$ is a nonempty open subset, $K\subset M_{f}(X)$ is a nonempty closed connected (or convex) subset.
-If $G_{K}^{\omega,U}\not=\emptyset$, then there exists a transitive point in the open set $U$ such that $V_{f}(x)=K$.
-Saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions.
-There are indeed certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3.
-Although the $\mu$-measure of $G_{\mu}$ is zero for every $\mu\in M_{f}(X)$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
+where $U\subset X$ is a non-empty open subset, $K\subset M_{f}(X)$ is a non-empty closed connected (or convex) subset.
+If $G_{K}^{\omega,U}\not$$\emptyset$, then there exists a transitive point in the sufficiently “small” open set $U$ such that $V_{f}(x)=K$.
+For systems with the specification property, we provide a complete answer to question (2).
+Our main results clarify that the saturated sets exhibit different dynamical behaviors in terms of different types of metric mean dimensions.
+There are certain variational principles for the Bowen and packing metric mean dimensions of saturated sets, as presented in Theorems 1.1, 1.2 and 1.3.
+Let $\overline{\rm mdim}_{M}^{B}(f,Z,d)$, $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ and $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ denote the Bowen/packing/upper capacity metric mean dimension of $Z$, respectively.
 Theorem 1.1.
-Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, and $U\subset X$ is a nonempty open set, then
-$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
-where $\overline{\rm mdim}_{M}^{P}(f,Z,d)$ is the packing upper metric mean dimension of $Z$, $h_{\mu}(f,\xi)$ is the measure-theoretic entropy of $\mu$ w.r.t. $\xi$.
+Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact convex set, and $U\subset X$ is a non-empty open set, then
+$\overline{\rm mdim}_{M}^{P}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 Theorem 1.2.
-Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact connected set, and $U\subset X$ is a nonempty open set, then
-$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
-where $\overline{\rm mdim}_{M}^{B}(f,Z,d)$ is the Bowen upper metric mean dimension of $Z$.
-Perhaps, more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$).
-However, up to now, the authors remain unaware of whether there exist any counterexamples among infinite entropy systems with the specification property for which the following equalities hold strictly
-$\sup_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d),$
-and
-$\inf_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d).$
-Inspired by the work of [GS21, CPV24], we define two variants of mean Rényi information dimension of invariant measures.
-Let $K$ be a closed set of $M_{f}(X)$.
-For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology.
-We define the (modified) lower and upper mean Rényi information dimensions of $\mu$ as
-$\underline{\rm mrid}_{\mu}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
-$\overline{\rm mrid}_{\mu}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}$
-respectively. Then this allows us to obtain the the variational principles in terms of the (modified) lower and upper mean Rényi information dimensions.
+Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact connected set, and $U\subset X$ is a non-empty open set, then
+$\overline{\rm mdim}_{M}^{B}\left(f,G_{K},d\right)=$ $\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 Theorem 1.3.
-Let $(X,d,f)$ be a t.d.s. with specification property. If $K\subseteq M_{f}(X)$ is a nonempty compact convex set, then
-$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\},$
-$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
-We remark that Theorem 1.3 does not require the connectedness of $K$ since any convex set in a topological vector space is automatically connected.
+Let $(X,d,f)$ be a TDS with the specification property. If $K\subseteq M_{f}(X)$ is a non-empty compact convex set, then
+$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d),$
+$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d),$
+where $\underline{\rm mrid}_{\mu}^{B}(f,K,d)$ and $\overline{\rm mrid}_{\mu}^{P}(f,K,d)$ denotes the Bowen and packing Rényi information dimensions of $\mu$, respectively.
+We remark that any convex set in a topological vector space is automatically connected.
+Although the $\mu$-measure of $G_{\mu}$ is zero for every non-ergodic measure $\mu$, its topological complexity can be sufficiently large, that is, the upper capacity metric mean dimension of saturated sets carries the full metric mean dimension, as stated in Theorem 1.4.
 Theorem 1.4.
-Let $(X,d,f)$ be a t.d.s. with specification property. Let $K\subseteq M_{f}(X)$ be a nonempty compact connected set, and $U\subset X$ be a nonempty open set. Then
+Let $(X,d,f)$ be a TDS with the specification property. Let $K\subseteq M_{f}(X)$ be a non-empty compact connected set, and $U\subset X$ be a non-empty open set. Then
 $\overline{\rm mdim}_{M}^{UC}\left(f,G_{K},d\right)=\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)=\overline{\rm mdim}_{M}(f,X,d),$
 where $\overline{\rm mdim}_{M}^{UC}(f,Z,d)$ is the upper capacity metric mean dimension of $Z$.
-As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K$ consists of exactly a single point.
-The Corollary 1.5 extends the result given in [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with specification property.
+As an immediate consequence of Theorems 1.1, 1.2 and 1.4, the most interesting case occurs when $K\subset M_{f}(X)$ consists of exactly a single point.
+The following Corollary 1.5 extends the result [YCZ25, Theorem 1.1] to all non-ergodic measures for systems with the specification property.
 Corollary 1.5.
-Let $(X,d,f)$ be a dynamical system with specification property. Then for every $\mu\in M_{f}(X)$,
+Let $(X,d,f)$ be a TDS with the specification property. Then for every $\mu\in M_{f}(X)$,
 $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),$
 $\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
-Consequently, if $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure in the sense of $\overline{\rm mdim}_{M}(f,X,d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$, then we have
-$\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}^{P}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}^{UC}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
-The “Consequently” part can be realized for some dynamical systems.
-For example, the full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$ with the compatible product metric $d(x,y)=\sum_{n\in\mathbb{Z}}\frac{||x_{n}-y_{n}||_{\mathbb{R}^{m}}}{2^{|n|}}$, and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is product measure of the Lebesgue measure $\mathcal{L}$ on $[0,1]^{m}$.
-The concept of maximal metric mean dimension measure was proposed in [YCZ23a].
-However, determining what types of infinite entropy systems can possess a maximal metric mean dimension measure remains a challenging task.
-See [LR24] for the existence of maximal metric mean dimension measure for typical conservative homeomorphisms.
-By Gutman- Śpiewak variational principle (as presented in Theorem 2.4) and Corollary 1.5, we have an abstract criterion for maximal metric mean dimension measure for identifying maximal metric mean dimension measures in systems with the specification property.
-Corollary 1.6.
-Let $(X,d,f)$ be a t.d.s.
-with specification property.
-An invariant measure $\mu\in M_{f}(X)$ is a maximal metric mean dimension measure if and only if $\overline{\rm mdim}^{B}_{M}(f,G_{\mu},d)=\overline{\rm mdim}_{M}(f,X,d).$
-Some comments are in order.
-  • (1)
-    There are plenty of infinite entropy systems with specification property, for instance, the usual full shift over the infinite product of $m$-dimensional cubes $[0,1]^{m}$, and the examples presented in [BR23, Section IV], and our main results can be applied to these systems.
-  • (2)
-    In Theorem 1.3, we do not know whether the equalities hold for Bowen upper metric mean dimension and packing lower metric mean dimension.
-The other theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
-  • (3)
-    We believe that it is trivial to extend the aforementioned results to metric mean dimensions with potential by making some modifications to the current proofs.
-The novelty of our work not only gives a proper and full extension of the previous work in [PS07, ZCC12, ZC13, HTW19] to the setting of infinite entropy systems of $\mathbb{Z}$-actions with specification property, but also may be helpful for studying the multifractal analysis of other types of fractal-like sets in infinite entropy systems.
-It remains unclear whether there are some possible substitutions of the uniform separation property so that all the above results are true for $g$-almost product property, or even for cases beyond $\mathbb{Z}$.
+Except Theorem 1.3, all theorems mentioned above are valid for the corresponding lower metric mean dimensions by considering changing the $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
+Besides, it becomes unclear whether the aforementioned results for weaker specification-like property.
 The paper is organized as follows.
-In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property.
+In section 2, we recall the precise definitions of metric mean dimensions and the associated properties, and two types of measure-theoretic $\epsilon$-entropies, and the specification property of TDSs.
 In section 3, we prove the main results.
-In section 4, we consider some applications of the main results, and apply it to investigate the level sets of Birkhoff average of a continuous potential, Li-Yorke chaos and the metric mean dimensions of generic points of the full shift over compact metric spaces.
+In section 4, we consider some applications of the main results, and apply it to investigate the level sets of the Birkhoff average of a continuous potential, Li-Yorke chaos in finite entropy systems and the metric mean dimensions of generic points of the full shift over compact metric spaces.
 2 Preliminaries
-In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of specification property in dynamical systems.
+In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of the specification property in dynamical systems.
 2.1 Three types of metric mean dimensions
 In this subsection, we review the definitions of upper capacity metric mean dimension of subsets[LW00], Bowen metric mean dimension[LV21, W21] and packing metric mean dimension [YCZ22] of subsets.
-Let $(X,d,f)$ be a t.d.s.. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on the space $X$ as
-$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\},$
+Let $(X,d,f)$ be a TDS. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on $X$ as
+$d_{n}(x,y):=\max_{0\leq i\leq n-1}d(f^{i}x,f^{i}y),$
 where $x,y\in X$. Then the Bowen open ball and closed ball centered at $x$ with radius $\epsilon$ in the metric $d_{n}$ are given by
-$B_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)<\epsilon\},$
-$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
+$B_{n}(x,\epsilon):$ $=\{y\in X:d_{n}(x,y)<\epsilon\},$
+$\overline{B}_{n}(x,\epsilon):$ $=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
 respectively.
-Let $Z$ be a nonempty subset of $X$.
+Let $Z$ be a non-empty subset of $X$.
 Fix $\epsilon>0$.
@@
 $h_{top}^{UC}(f,Z,\epsilon)=\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(Z,\epsilon).$
-Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, i.e.,
+Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 0, i.e.,
 $h_{top}^{UC}(f,Z)=\lim_{\epsilon\to 0}h_{top}^{UC}(f,Z,\epsilon)=\sup_{\epsilon>0}h_{top}^{UC}(f,Z,\epsilon).$
-Specifically, topological entropy fails to distinguish different levels of topological complexity within the infinite entropy systems
-Analogous to the definition of box dimension in fractal geometry, we formulate the refined entropy-like quantity for infinite entropy systems using $\epsilon$-upper capacity topological entropy.
+To get more information about the dynamics of infinite entropy systems, analogous to the definition of box dimension in fractal geometry, we formulate a refined entropy-like quantity using the $\epsilon$-upper capacity topological entropy.
 Definition 2.1.
@@
 $\overline{\rm mdim}_{M}^{UC}(f,Z,d):=\limsup_{\epsilon\to 0}\frac{h_{top}^{UC}(f,Z,\epsilon)}{|\log\epsilon|}.$
-In the case where $Z$ is equal to $X$, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$ respectively.
-Then the upper capacity metric mean dimension of $X$ is reduced to Lindenstrauss-Weiss’s definition introduced in [LW00].
-Such a formulation extends the concept of the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
-Notice that the (upper capacity) metric mean dimension is defined by detecting how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon$ tends to 00.
+If $Z=X$, the upper capacity metric mean dimension of $X$ is reduced to the metric mean dimension introduced by Lindenstrauss-Weiss [LW00]; in this case, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively.
+Such a formulation extends the concept of metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
+Notice that the (upper capacity) metric mean dimension is defined by measuring how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon\to 0$.
 That is, for sufficiently small $\epsilon>0$, we may think of
 $h_{top}^{UC}(f,Z,\epsilon)\approx\overline{\rm mdim}_{M}^{UC}(f,Z,d)\cdot|\log\epsilon|.$
-It is straightforward to observe that any system with finite topological entropy has zero metric mean dimension.
-This, in turn, shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
-The classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogue of the Hausdorff dimension and packing dimension in fractal geometry.
-The corresponding dimensional characterizations of metric mean dimensions, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the Bowen and packing metric mean dimensions.
-One may refer to [BR23, CLS21, LL24, FKO24, YCZ25] for the results involving Bowen and packing metric mean dimensions.
-For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
+It is easy to see that any system with finite topological entropy has zero metric mean dimension.
+This shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
+The Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogues of the Hausdorff dimension and packing dimension in fractal geometry.
+The corresponding dimensional characterizations for metric mean dimension, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the following Bowen and packing metric mean dimensions.
+For $Z\subset X$, $s>0$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
 $M(Z,s,N,\epsilon)=\inf\left\{\sum_{i\in I}\exp(-n_{i}s)\right\},$
-where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
-The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases, this ensures that the limit
+where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$ for all $i\in I$.
+The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So the limit
 $M(Z,s,\epsilon)=\lim_{N\to\infty}M(Z,s,N,\epsilon)$
-exists.
-There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00.
-That is, the critical value is given by
-$h_{top}^{B}(f,Z,\epsilon)=\inf\{s>0:M(Z,s,\epsilon)=0\}=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
+exists. There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 0, i.e.,
+$h_{top}^{B}(f,Z,\epsilon)$ $=\inf\{s>0:M(Z,s,\epsilon)=0\}$
+$=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
 Definition 2.2.
@@
 $\overline{\rm mdim}^{B}_{M}(f,Z,d)=\limsup_{\epsilon\to 0}\frac{h_{top}^{B}(f,Z,\epsilon)}{|\log\epsilon|}.$
-For $Z\subseteq X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
+For $Z\subseteq X$, $s>0$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
 $P(Z,s,N,\epsilon)=\sup\left\{\sum_{i\in I}\exp(-sn_{i})\right\},$
@@
 $P(Z,s,\epsilon)=\lim_{N\to\infty}P(Z,s,N,\epsilon).$
-Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So some modifications are needed to obtain the property:
+Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So the modifications are needed to obtain this property:
 $\mathcal{P}(Z,s,\epsilon):=\inf\left\{\sum_{i=1}^{\infty}P(Z_{i},s,\epsilon):\bigcup_{i=1}^{\infty}Z_{i}\supseteq Z\right\}.$
-There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00.
+There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 0.
 That is, the critical value is given by
@@
 If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$.
-We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6].
+We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23, Proposition 2.8].
 Proposition 2.1.
@@
 $(2)$ If $Z=\bigcup_{i=1}^{\infty}Z_{i}$, then for every $\epsilon>0$,
-$h_{top}^{B}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{B}(f,Z_{i},\epsilon)$
+$h_{top}^{B}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{B}(f,Z_{i},\epsilon),$
 $h_{top}^{P}(f,Z,\epsilon)$ $=\sup_{i\geq 1}h_{top}^{P}(f,Z_{i},\epsilon),$
@@
 $\overline{\rm mdim}_{M}^{S}(f,Z,d)$ $=\max_{1\leq i\leq N}\overline{\rm mdim}_{M}^{S}(f,Z_{i},d).$
-$(3)$ If $Z$ is a nonempty subset of $X$, then for every $10<\epsilon<$ 1, we have
+$(3)$ If $Z$ is a non-empty subset of $X$, then for every $0<\epsilon<1$, we have
 $h_{top}^{B}(f,Z,3\epsilon)\leq h_{top}^{P}(f,Z,\epsilon)\leq h_{top}^{UC}(f,Z,\epsilon).$
 Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$.
-Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions are equal.
-Given a subset of phase space, it is not easy to calculate its precise Bowen and packing metric mean dimensions.
-However, in the multifractal analysis of dynamical systems the Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets.
-Consequently, the powerful tool yields the lower bounds of metric mean dimension of the fractal-like sets.
-$\epsilon$-Bowen and packing topological entropies of Entropy Distribution Principles are stated as follows:
+Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions coincide.
+Given a certain subset of phase space, in general it is not easy to calculate its precise Bowen and packing metric mean dimensions.
+The Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets, and hence yields the lower bounds of metric mean dimension of the fractal-like sets.
+The Entropy Distribution Principles of $\epsilon$-Bowen and packing topological entropies are stated as follows:
 Lemma 2.1.
-[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s.
-and $\epsilon>0$.
-Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$.
-If there exist a $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and for every $n$ and for each Bowen ball $B_{n}(x,\epsilon)$ with $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, one has $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$, then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
+[BR23, Lemma 13] Let $(X,d,f)$ be a TDS and $\epsilon>0$.
+Suppose that $Z$ is a Borel set of $X$ and $0<s<\infty$.
+If there exist $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for every each Bowen ball $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, then
+$h_{top}^{B}(f,Z,\epsilon)\geq s.$
 Lemma 2.2.
-Let $(X,d,f)$ be a t.d.s.
-and $\epsilon>0$.
-Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$.
-If there exist a $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ that converges to $\infty$ satisfying $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
+Let $(X,d,f)$ be a TDS and $\epsilon>0$.
+Suppose that $Z$ is a Borel set of $X$ and $0<s<\infty$.
+If there exist $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ such that $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then
+$h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s.$
 Proof.
-By definition, it is enough to prove that for sufficiently small $\delta>0$,
+By definition of the $\epsilon$-packing topological entropy, it is enough to prove that for every sufficiently small $\delta>0$,
 $P\left(E,s-\delta,\frac{\epsilon}{6}\right)=\infty$
-holds for any Borel subset $E\subset Z$ with $\mu(E)>0$.
-For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of the set $E$.
-Then, by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ Hence, one has
+for any Borel subset $E\subset Z$ with $\mu(E)>0$.
+For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$.
+Then, by $5r$-covering lemma [Mat95, Theorem 2.1], there exists a countable pairwise disjoint sub-collection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ In particular, the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint.
+Hence, one has
 $P\left(E,s-\delta,n_{i},\frac{\epsilon}{6}\right)\geq\sum_{j\in J}\exp[-n_{i}(s-\delta)]\geq\frac{e^{n_{i}\delta}}{C}\sum_{j\in J}\mu(B_{n_{i}}(x_{j},\epsilon))\geq\frac{e^{n_{i}\delta}}{C}\mu(E).$
@@
 In this subsection, we recall the definitions of Kolmogorov-Sinai $\epsilon$-entropy and Katok $\epsilon$-entropy of invariant measures.
-Let $C(X)$ denote the space consisting of real-valued continuous functions defined on $X$, equipped with the supremum norm.
+Let $C(X)$ denote the space of real-valued continuous functions on $X$, equipped with the supremum norm $||\cdot||$.
 By virtue of the Riesz representation theorem for compact metric spaces, we are able to endow $M(X)$ with the weak-topology, which is metrizable by the following metric:
@@
 Thus, the diameter of $M(X)$ in the sense of the metric $D$ is always less than $1$, and for fixed $\mu,\nu\in M(X)$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$.
-It follows from the weak-topology on $M(X)$ that the natural embedding from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous.
-Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the sense of the weak-topology.
-Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\cap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\cup_{i=1}^{k}C_{i})=0$.
-Then the partition entropy of $\xi$ w.r.t.
-$\mu$ is defined by
-$H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$
-We define the measurable partition of $\mu\in M_{f}(X)$ w.r.t. $\xi$ as
+Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the weak-topology.
+Given $\mu\in M_{f}(X)$ and a finite Borel partition $\xi=\left\{C_{1},\cdots,C_{k}\right\}$, the partition entropy of $\xi$ w.r.t.
+$\mu$ is defined by $H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$ We define the measure-theoretic entropy of $\mu$ w.r.t.
+$\xi$ as
 $h_{\mu}(f,\xi)=\lim_{n\to+\infty}\frac{1}{n}H_{\mu}(\xi^{n}),$
-where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
+Denote by $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the preimage partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
 $h_{\mu}(f)=\sup_{\xi}h_{\mu}(f,\xi),$
-where the supremum is taken over all finite Borel measurable partitions of $X$.
-For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions is the measure-theoretic metric mean dimensions of invariant measures by measuring the divergent rates of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others.
-A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf.
-[YCZ25, Theorem 1.1]).
+where the supremum ranges over all finite Borel partitions of $X$.
+For topological dynamical systems, the measure-theoretic counterpart of metric mean dimension is the measure-theoretic metric mean dimension of invariant measures.
+It measures the divergent rate of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others.
+A systemic study about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25].
 Here, we only recall two types of the measure-theoretic $\epsilon$-entropies.
+Let $\mu\in M_{f}(X)$ and $\epsilon>0$.
   • (a)
-    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu\in M_{f}(X)$ is given by
-    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+    The Kolmogorov-Sinai $\epsilon$-entropy [GS21] of $\mu$ is given by
+    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+    where the infimum is taken over all finite Borel partitions of $X$ with diameter less than $\epsilon$.
   • (b)
-    For $\mu\in M_{f}(X)$, $\epsilon>0$, $\delta\in(0,1)$ and $n\in\mathbb{N}$, we put
+    For $\delta\in(0,1)$ and $n\in\mathbb{N}$, we put
     $N_{\mu}(n,\epsilon,\delta):=\min\{\#E:E\subset X~\emph{and}~\mu(\cup_{x\in E}B_{n}(x,\epsilon))>1-\delta\}.$
-    The lower Katok $\epsilon$-entropy of $\mu$ is given by
+    The Katok $\epsilon$-entropy of $\mu$ [Kat80] is defined by
     $\underline{h}_{\mu}^{K}(f,\epsilon,\delta)=\liminf_{n\to\infty}\frac{1}{n}\log N_{\mu}(n,\epsilon,\delta).$
@@
 Lemma 2.3.
-[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$.
-Then for any $\delta\in(0,1)$, one has
-$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
-Consequently, $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
-The following variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
+[S22, Lemma 4.1] Let $(X,d,f)$ be a TDS and $\mu\in M_{f}^{e}(X)$, If $\mathcal{U}$ is a finite open cover of $X$ with the diameter ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and the Lebesgue number ${\rm Leb}(\mathcal{U})\geq\epsilon$, then for any $\delta\in(0,1)$,
+$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta),$
+where the infimum $\xi$ ranges over all finite Borel partitions of $X$ refining than $\mathcal{U}$.
+Consequently, for every $\epsilon>0$, one has $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
+We state the variational principles between the metric mean dimension and the above two types of measure-theoretic $\epsilon$-entropies.
 Theorem 2.4.
-(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with a metric $d$. Then for any $\delta\in(0,1)$,
+(cf. [YCZ25, Theorems 1.2 and 1.3]) Let $(X,d,f)$ be a TDS. Then for any $\delta\in(0,1)$,
 $\overline{\rm mdim}_{M}(f,X,d)$ $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(X)}F(\mu,\epsilon),$
@@
 where $F(\mu,\epsilon)\in\{\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(f,\xi),\underline{h}^{K}_{\mu}(f,\epsilon,\delta)\}$.
-The results are also valid for lower metric mean dimension.
-A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is summarized in [YCZ25, Theorems 1.2 and 1.3].
-2.3 Specification property
-Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as specification property.
+The results are also valid for $\underline{\rm mdim}_{M}(f,X,d)$ by changing $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
+2.3 Specification property of TDSs
+In general, it is difficult to find a real orbit with the desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as the specification property in TDSs.
 Definition 2.4.
-A t.d.s.
-$(X,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
+One says that a TDS $(X,d,f)$ has the specification property if for any $\epsilon>0$, there is a positive integer $m=m(\epsilon)$ such that for any finite collection of intervals $I_{j}=[a_{j},b_{j}]\subset\mathbb{N}$ with $a_{j+1}-b_{j}\geq m(\epsilon)$, $j=1,\dots,k-1$, and any collection of $k$ points $\{x_{1},\dots,x_{k}\}$ of $X$, there exists a point $x\in X$ such that
 $d(f^{p}x_{j},f^{p+a_{j}}x)<\epsilon$
@@
 $\{x_{j},f(x_{j}),\ldots,f^{b_{j}-a_{j}}(x_{j})\},\ j=1,\ldots,k.$
-For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $[0,1]^{\mathbb{Z}}$ have specification property.
-More examples of infinite entropy systems with specification property can be found in [BR23, Section IV].
+For instance, every topological mixing locally maximal hyperbolic set, and the full shift on $0,1]^{\mathbb{Z}}$ have the specification property.
+More examples of infinite entropy systems with the specification property can be found in [BR23, Section IV].
 A point $x\in X$ is called almost periodic if for any open neighborhood $U$ of $x$, there exists $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$, there exists an integer $k$ with $n\leq k\leq n+N$ for which $f^{k}(x)\in U$.
-Let $AP(X)$ denote the set of all almost periodic points of $X$.
-The following statement, proved in [HTZ23, Proposition 2.11] for systems with $g$-almost product property, gives a characterization of the structure of $AP(X)$.
+Denote by $AP(X)$ the set of all almost periodic points of $X$.
 Proposition 2.2.
-If the dynamical system $(X,f)$ has specification property, then the almost periodic set AP(X) is dense in $X$.
+[HTZ23, Proposition 2.11] If a TDS $(X,d,f)$ has the specification property, then the almost periodic set AP(X) is dense in $X$.
 3 Proofs of main results
-This section proves the main results of this paper.
-The challenging aspect lies in the estimation of the lower bound of $\overline{\rm mdim}_{M}^{S}(f,G_{K}^{\omega,U},d)$, where $S\in\{B,P,UC\}$.
-We will construct a Moran fractal set, which is geometrically constructed and is contained in $G_{K}^{\omega,U}$.
-Consequently, this set is well-suited for computing the lower bound of the metric mean dimensions of saturated sets.
+In this section, we prove the main results of this paper.
 3.1 Packing metric mean dimension of saturated sets
-We first prove Theorem 1.1. To arrive at this goal, we divide its proof into two parts.
+To prove Theorem 1.1, we divide its proof into two parts.
 3.1.1 Upper bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K},d)$
-We employ a general method presented in [PS07, Theorem 4.1], [ZCC12, Theorem 1.1, $(2)$] and [JCZ22, Proposition 4] to derive the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$.
 Lemma 3.1.
-[Y23, Lemma 3.1] Let $(X,f)$ be a t.d.s.
-and $\epsilon>0$.
+[Y23, Lemma 3.1] Let $(X,d,f)$ be a TDS and $\epsilon>0$.
 Let $\{F_{n}\}$ be a sequence of non-empty subsets of $X$, and let $E_{n}$ be an $(n,\epsilon)$-separated subset of $F_{n}$ with the largest cardinality.
@@
 Proof.
-This essentially adheres to the proof of the upper bound of the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
-Recall $G_{K}=\{x\in X:V_{f}(x)=K\}$. For $K\subset M_{f}(X)$, we set
-$G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
-Obviously, we have $G_{K}\subset G^{K}$.
+It essentially follows from the proof of the upper bound for the classical variational principle (cf.[Wal82, Theorem 8.6]). ∎
 Lemma 3.2.
-Let $(X,f)$ be a t.d.s..
-$(1)$ Let $K\subset M_{f}(X)$ be a nonempty closed convex subset. Then for any $\epsilon>0$,
-$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-$(2)$ If $\mu\in M_{f}(X)$, then for any $\epsilon>0$,
+Let $(X,d,f)$ be a TDS, and let $K\subset M_{f}(X)$ be a non-empty closed convex subset. Then for any $\epsilon>0$,
+$h_{top}^{P}(f,G_{K},\epsilon)\leq h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+where $G^{K}:=\{x\in X:V_{f}(x)\subset K\}.$
+Consequently, for every $\mu\in M_{f}(X)$ and $\epsilon>0$,
 $h_{top}^{P}(f,G_{\mu},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 Proof.
-It suffices to show $(1)$ since one can take $K=\{\mu\}$ as in $(1)$ and use $G_{\mu}=G^{\{\mu\}}$ to get the desired $(2)$.
+It suffices to show the inequality
+$h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 For $n\in\mathbb{N}$ and $\delta>0$, we put
-$\mathcal{R}_{K}(\delta,k):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~{}\forall n\geq k\right\},$
-$\mathcal{Q}_{K}(\delta,n):$ =$=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
-and
+$\mathcal{R}_{K}(\delta,k):$ $=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta),~\forall n\geq k\right\},$
+$\mathcal{Q}_{K}(\delta,n):$ $=\left\{x\in X:\ \mathcal{E}_{n}(x)\in B(K,\delta)\right\},$
+and define
 $\Theta(K,\epsilon):=\lim_{\delta\to 0}\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon),$
 where $s_{n}(\mathcal{Q}_{K}(\delta,n),\epsilon)$ is the largest cardinality of $(n,\epsilon)$-separated sets of $\mathcal{Q}_{K}(\delta,n)$.
-Fix $x\in G^{K}$.
-Since we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$, then the index $n$ with $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite.
+For every $x\in G^{K}$, we have $V_{f}(x)\subset B(K,\delta)$ for every $\delta>0$.
+Then the cardinality of $n$ such that $\mathcal{E}_{n}(x)\in M_{f}(X)\backslash B(K,\delta)$ must be finite.
 This shows that $x\in\mathcal{R}_{K}(\delta,k)$ for some $k$.
-Then, by $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
+By $G^{K}\subset\cup_{k=1}^{\infty}\mathcal{R}_{K}(\delta,k)$ for every $\delta>0$, it follows from Proposition 2.1 that
 $h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{k}h_{top}^{P}(f,\mathcal{R}_{K}(\delta,k),\epsilon)\leq\sup_{k}h_{top}^{UC}(f,\mathcal{R}_{K}(\delta,k),\epsilon).$ (3.1)
@@
 $\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
-For any $k\geq 1$, by definitions, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
+For any $k\geq 1$, it holds that $\mathcal{R}_{K}(\delta_{0},k)\subset\mathcal{Q}_{K}(\delta_{0},n)$ for any $n\geq k$. Thus, we have
 $h_{top}^{UC}(f,\mathcal{R}_{K}(\delta_{0},k),\epsilon)\leq\limsup_{n\to\infty}\frac{1}{n}\log s_{n}(\mathcal{Q}_{K}(\delta_{0},n),\epsilon)<\Theta(K,\epsilon)+\tau.$
-Combining (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
-Next we shall prove for any $\gamma>0$, there is $\mu\in K$ such that
-$\Theta(K,\epsilon)-\gamma\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Take a strictly decreasing sequence $\{\delta_{k}\}$ of positive real numbers with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
+Using (3.1) and letting $\tau\to 0$, we conclude that $h_{top}^{P}(f,G^{K},\epsilon)\leq\Theta(K,\epsilon)$.
+Take a strictly decreasing sequence $\{\delta_{k}\}$ with $\delta_{k}\to 0$ and a strictly increasing sequence $\{n_{k}\}$ with $n_{k}\to\infty$ such that
 $s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)>\exp[n_{k}(\Theta(K,\epsilon)-\gamma)].$
-Let $F_{k}$ be a maximal $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$. We set
+Let $F_{k}$ be a $(n_{k},\epsilon)$-separated subset of $\mathcal{R}_{K}(\delta_{k},n_{k})$ with the largest cardinality. We set
 $\mu_{k}:=\frac{1}{\#F_{k}}\sum_{x\in F_{k}}\mathcal{E}_{n_{k}}(x).$
 Then $\mu_{k}\in B(K,\delta_{k})$ since $K$ is convex.
-In the weak-topology, choose a subsequence $k_{j}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$.
+Now choose a subsequence $\{n_{k_{j}}\}$ such that $\mu_{k_{j}}\to\mu$ as $j\to\infty$.
 Then $\mu\in K$ since $K$ is a closed.
 By Lemma 3.1, we have
-$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k}}(\mathcal{Q}_{K}(\delta_{k},n_{k}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+$\Theta(K,\epsilon)-\gamma\leq\limsup_{j\to\infty}\frac{1}{n_{k_{j}}}\log s_{n_{k_{j}}}(\mathcal{Q}_{k_{j}}(\delta_{k_{j}},n_{k_{j}}),\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 This implies that $h_{top}^{P}(f,G^{K},\epsilon)\leq\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
-Thus, by the Lemma 3.2 the upper bound of the packing metric mean dimension of $G_{K}$ is given by
+Thus, by Lemma 3.2 the upper bound of $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is bounded above by
 $\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.2)
 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
-For any small enough $\epsilon^{\ast}>0$, we prove that
+The challenging direction is how to estimate the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$.
+This shall be done by constructing a geometric Moran fractal set that is contained in $G_{K}^{\omega,U}$.
+For any sufficiently small $\epsilon^{\ast}>0$, we prove that
 $\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{12}).$ (3.3)
@@
 Since $K$ is a connected and compact set, we choose a sequence $\{\mu_{(k,i)}:k\in\mathbb{N},1\leq i\leq Q_{k}\}\subseteq K$ such that
-[…]
+$K\subset\bigcup_{i=1}^{Q_{k}}B(\mu_{(k,i)},\frac{1}{k}),\ D(\mu_{(k,i)},\mu_{(k,i+1)})$ $\leq\frac{1}{k},\ 1\leq i<Q_{k}$
 $D(\mu_{(k,Q_{k})},\mu_{(k+1,1)})$ $\leq\frac{1}{k},\ \mu_{(k,Q_{k})}=\mu^{\ast}.$
@@
 We shall simply write $\delta_{k}=\delta_{(k,Q_{k})}$.
-For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$, where $m$ is chosen as in the definition of specification property.
-For any nonempty open set $U$, there exist $\tilde{\epsilon}>0$ and a point $a_{0}\in X$ such that $\overline{B(a_{0},\tilde{\epsilon})}\subseteq U$.
-One may assume that $\epsilon^{\ast}<\tilde{\epsilon}$.
-Set
+For $(k,j)\in\mathcal{A}$, let $\epsilon_{(k,j)}:=\frac{\epsilon^{\ast}}{2^{\iota(k,j)+2}}$, and $m_{(k,j)}:=m(\epsilon_{(k,j)})$ be chosen as in the definition of the specification property.
+For any non-empty open set $U$, there exist $0<\tilde{\epsilon}<\epsilon^{*}$ and a point $a_{0}\in X$ such that ${\overline{B}(a_{0},\tilde{\epsilon})}\subseteq U$. Set
 $\mathcal{C}_{(0,0)}=\{a_{0}\},\ n_{(0,0)}=1,\ N_{(0,0)}=1.$
@@
 For each $k\in\mathbb{N}$, there exists $\nu_{k}\in M_{f}(X)$ satisfying
-$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\{\lambda_{i}^{k}:1\leq i\leq j_{k}\}$ are rational positive numbers, $\forall 1\leq i\leq j_{k}$ with $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$ and $\nu_{i}^{k}\in M_{f}^{e}(X)$;
+$(1)$ $v_{k}=\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\nu_{i}^{k}$, where $\nu_{i}^{k}\in M_{f}^{e}(X)$, $\lambda_{i}^{k}\in\mathbb{Q}^{+}$ for every $1\leq i\leq j_{k}$, and $\sum_{i=1}^{j_{k}}\lambda_{i}^{k}=1$;
 $(2)$ one has
@@
 $(2)$ there exists a subset $E_{k,i}$ that is a $(\lambda_{i}^{k}n_{k}^{\prime},5\epsilon^{\ast})$-separated set of $Y_{i}(N)$ satisfying
-#$\#E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
+$E_{k,i}\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\underline{h}_{\nu_{i}^{k}}^{K}(f,5\epsilon^{\ast},\zeta)-\zeta\right)\right]\geq\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\cdot\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right],$
 where the second inequality is derived from Lemma 2.3.
@@
 $D(\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x),\nu_{i}^{k})<\delta_{k}.$
-By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ so that
-[…]
+By the specification property, for $k\geq 1$ and each $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}$, there exists $y=y(\underline{x}_{k})\in X$ such that
+$d\left(f^{T_{i}+t_{i}}(y),f^{t_{i}}(x_{i})\right)<\epsilon_{(k,Q_{k})}<\frac{\epsilon^{\ast}}{2^{k+2}},\ 1\leq i\leq j_{k},\ 0\leq t_{i}\leq\lambda_{i}^{k}n_{k}^{\prime}-1,$
 where $T_{1}=0$ and $T_{i}=\lambda_{1}^{k}n_{k}^{\prime}+\cdots+\lambda_{i-1}^{k}n_{k}^{\prime}+(i-1)m_{k}$ for $2\leq i\leq j_{k}$.
-One may assume that $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$ since the map $x\mapsto\delta_{x}$ is uniformly continuous.
 For any two distinct $\underline{x}_{k}=(x_{1},x_{2},\ldots,x_{j_{k}})$, $\underline{x}_{k}^{\prime}=(x_{1}^{\prime},x_{2}^{\prime},\ldots,x_{j_{k}}^{\prime})$ in $E_{k,1}\times\cdots\times E_{k,j_{k}}$, one has $y(\underline{x}_{k})\neq y^{\prime}(\underline{x}_{k}^{\prime})$.
@@
 $<\frac{\epsilon^{\ast}}{2^{k+1}}+d_{\lambda_{i}^{k}n_{k}^{\prime}}\left(f^{T_{i}}y,f^{T_{i}}y^{\prime}\right),$
-which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$. Let
+which implies $d_{n_{k}}(y,y^{\prime})>4\epsilon^{\ast}$.
+Let
 $\mathcal{C}_{(k,Q_{k})}=\left\{y(x_{1},x_{2},\ldots,x_{j_{k}}):\ (x_{1},x_{2},\ldots,x_{j_{k}})\in E_{k,1}\times E_{k,2}\times\cdots\times E_{k,j_{k}}\right\}.$
 Then $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set and
-#$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
+$\#\mathcal{C}_{(k,Q_{k})}$ $=\Pi_{i=1}^{j_{k}}\#E_{k,i}\geq\Pi_{i=1}^{j_{k}}\exp\left[\lambda_{i}^{k}n_{k}^{\prime}\left(\inf_{\xi\succ\mathcal{U}}h_{\nu_{i}^{k}}(f,\xi)-\zeta\right)\right]$
 […]
@@
 $\leq$ $\left[\left|\frac{1}{n_{k}^{\prime}}-\frac{1}{n_{k}}\right|\cdot n_{k}^{\prime}+\frac{n_{k}-n_{k}^{\prime}}{n_{k}}\right]\cdot\|\varphi\|<\frac{2}{k}\|\varphi\|,$
-which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)))<\frac{2}{k}$. Since
+which shows that $D(\mathcal{E}_{n_{k}}(y),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}f^{T_{i}}(y))<\frac{2}{k}$.
+Since the map $x\mapsto\delta_{x}$ is uniformly continuous, we may assume that for any $x,y\in X$ with $d(x,y)<\frac{\epsilon^{\ast}}{2^{k+2}}$ implies $D(\delta_{x},\delta_{y})<\frac{1}{k}$.
+Since
 $D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i}),\nu_{k}\right)<\delta_{k},\ \ D\left(\nu_{k},\mu^{\ast}\right)<\delta_{k},$
@@
 $D\left(\mathcal{E}_{n_{k}}(y),\mu^{\ast}\right)\leq\frac{2}{k}+D\left(\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(f^{T_{i}}(y)),\sum_{i=1}^{j_{k}}\lambda_{i}^{k}\mathcal{E}_{\lambda_{i}^{k}n_{k}^{\prime}}(x_{i})\right)+2\delta_{k}<\frac{3}{k}+2\delta_{k}.$ (3.8)
-Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k$. We proceed to consider the cases for $1\leq j<Q_{k}$.
-Since the system has specification property, the generic point $G_{\mu}\neq\emptyset$ holds for any $\mu\in M_{f}(X)$ [PS07].
-Applying this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, then there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
+Until now, by virtue of the specification property, we have obtained two sets $\mathcal{C}_{(k,0)}$ and $\mathcal{C}_{(k,Q_{k})}$ for each $k\geq 1$.
+We proceed to consider the cases for $1\leq j<Q_{k}$.
+Since the system has the specification property, the generic point $G_{\mu}\neq\emptyset$ holds for all $\mu\in M_{f}(X)$ [PS07].
+Using this fact for any $k\geq 1$ and $1\leq j<Q_{k}$, there exist $x_{k,j}\in G_{\mu_{(k,j)}}$ and $n_{(k,j)}\in\mathbb{N}$ such that
 $D(\mathcal{E}_{n_{(k,j)}}(x_{k,j}),\mu_{(k,j)})<\delta_{(k,j)},$
@@
 Now set $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
-With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready for the construction:
+With these data $\{\mathcal{C}_{(k,j)}\}$ and $\{n_{(k,j)}\}$, we are ready to proceed the next construction:
 Step $1$. Construct the intermediate sets $\mathcal{S}_{(k,j)}$.
@@
 $\begin{split}\sum_{0\leq k^{\prime}\leq k+1}(n_{(k^{\prime},0)}+m_{(k^{\prime},0)})+n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}&\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})],\\ \sum_{(k^{\prime},j^{\prime})\leq(k,j)}[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})]&\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k},\\ n_{k}N_{(k,Q_{k})}&\geq(1-\zeta)\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}].\end{split}$ (3.9)
-For $(k,j)\in\mathcal{A}$ with $j=0$, we set $\mathcal{S}_{(k,j)}=\mathcal{C}_{(k,j)}$ and $c_{k,j}=n_{(k,j)}$.
-We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
-Case $1$. $j\neq Q_{k}$.
-Recall that $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$ is defined above.
-By specification property again, there exists $y_{k,j}\in X$ so that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$.
+For $(k,0)\in\mathcal{A}$, we set $\mathcal{S}_{(k,0)}:=\mathcal{C}_{(k,0)}$ and $c_{k,0}:=n_{(k,0)}$. We divide the following two cases to define the rest $\mathcal{S}_{(k,j)}$ indexed by $j\neq 0$.
+Case $1$. $1\leq j<Q_{k}$.
+Recall $\mathcal{C}_{(k,j)}=\{x_{k,j}\}$.
+By the specification property again, there exists $y_{k,j}\in X$ such that the pieces of orbits $\{x_{k,j},\cdots,f^{n_{(k,j)-1}}(x_{(k,j)})\}$ repeating $N_{(k,j)}$ times, can be $\epsilon_{(k,j)}$-traced by $y_{(k,j)}$ with the gap $m_{(k,j)}$.
 Now we set $\mathcal{S}_{(k,j)}=\{y_{k,j}\}$, and $c_{k,j}=n_{(k,j)}N_{(k,j)}+(N_{(k,j)}-1)m_{(k,j)}$.
@@
 $\mathcal{C}_{(k,Q_{k})}=\{y_{i}^{k}:i=1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}.$
-For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$.
+For any $(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in\{1,2,\ldots,\#\mathcal{C}_{(k,Q_{k})}\}^{N_{(k,Q_{k})}}$, let $y(i_{1},\ldots,i_{N_{(k,Q_{k})}})\in X$ be a point given by the specification property, such that the orbit $\epsilon_{(k,Q_{k})}$-shadows, with the gap $m_{(k,Q_{k})}$, the pieces of orbits $\{y_{i_{l}}^{k},f(y_{i_{l}}^{k}),\ldots,f^{n_{k}-1}(y_{i_{l}}^{k})\}$, $l=1,2,\ldots,N_{(k,Q_{k})}$.
 We write
@@
 Since $\mathcal{C}_{(k,Q_{k})}$ is an $(n_{k},4\epsilon^{\ast})$-separated set, this implies that $\mathcal{S}_{(k,Q_{k})}$ is a $(c_{k,Q_{k}},3\epsilon^{\ast})$-separated set.
-Then, by (3.1.2) and (3.9), one has the estimation for the cardinality of $\mathcal{S}_{(k,Q_{k})}$:
-#$\#\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.10)
-Similar to the arguments in (3.8) one has
-$D(\mathcal{E}_{c_{k,j}}(y_{k,j}),\mu_{(k,j)})<\frac{2}{k}+\delta_{(k,j)},\ \forall\ y_{k,j}\in\mathcal{S}_{(k,j)}~{}\text{with}~{}k\geq 1\ \text{and }1\leq j<Q_{k},$
+Then, by (3.1.2) and (3.9), the cardinality of $\mathcal{S}_{(k,Q_{k})}$ is bounded blew by
+$\mathcal{S}_{(k,Q_{k})}=\#\mathcal{C}_{(k,Q_{k})}^{N_{(k,Q_{k})}}\geq\exp\left[\sum_{0\leq j\leq Q_{k}}[(n_{(k,j)}+m_{(k,j)})N_{(k,j)}]\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$ (3.10)
+Similar to the arguments in (3.8), one has
+[…]
 and
@@
 Hence, noting that $\mu_{(k,Q_{k})}=\mu^{*}$, in an unified manner we always have
-$D(\mathcal{E}_{c_{k,j}}(y),\mu_{(k,j)})<\frac{5}{k}+2\delta_{(k,j)},\ \forall\ y\in\mathcal{S}_{(k,j)}\ \text{with }k\geq 1,\ 1\leq j\leq Q_{k}.$ (3.11)
+[…]
 Step $2$. Construct the $\mathbb{F}_{(k,j)}$ and the Moran set $\mathbb{F}$.
-One may regard this process as an analogue of the construction of the middle three Cantor set of the unit interval $[0,1]$.
+This step can be regarded as an analogue of the construction of the middle-third Cantor set of $[0,1]$.
 Let $\mathcal{I}_{(0,0)}=\mathcal{S}_{(0,0)}$ and $M_{(0,0)}=n_{(0,0)}$.
@@
 $M_{(k,j)}$ $=M_{(k,j)^{\ast}}+m_{(k,j)}+c_{k,j};$
-[…]
-Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in formula (3.10) has the estimation:
-#$\#\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)
+$M_{(k,j)}^{i}$ $=M_{(k,j)^{\ast}}+i(m_{(k,j)}+n_{(k,j)}),\ 0\leq i\leq N_{(k,j)}.$
+Note that $M_{(k,j)}^{N_{(k,j)}}=M_{(k,j)}$ and the number of $\mathcal{S}_{(k,Q_{k})}$ given in (3.10) has the estimation:
+$\mathcal{S}_{(k,Q_{k})}\geq\exp\left[(M_{(k,Q_{k})}-M_{(k-1,Q_{k-1})})(1-\zeta)^{2}(H-3\zeta)\right].$ (3.12)
 For $x\in\mathcal{I}_{(k,j)^{\ast}}$ and $y\in\mathcal{S}_{(k,j)}$, by the specification property let $z=z(x,y)\in X$ be some point such that
@@
 which leads to
-#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
+$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
 Moreover, according to (3.11), for $z\in\mathcal{I}_{(k,j)}$ with $j\neq 0$, we have
@@
 $\overline{B}_{M_{(k,j)}}(z,2\epsilon_{(k,j)})\subset\overline{B}_{M_{(k,j)^{\ast}}}(x,2\epsilon_{(k,j)^{\ast}}).$
-This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$. Let
+This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing non-empty closed subsets of $X$. Let
 $\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}=\bigcap_{k\in\mathbb{N}}\mathbb{F}_{(k,Q_{k})}.$
@@
 $\mathbb{F}\subset\mathbb{F}_{(0,0)}=\overline{B}_{n(0,0)}(a_{0},\frac{\epsilon^{\ast}}{8})\subset{\overline{B}(a_{0},\tilde{\epsilon})}\subset U.$
-For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$.
-Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$.
-This implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$.
+For $p\in\mathbb{F}$, we have $p\in\mathbb{F}_{(k,0)}$ for any $k\geq 1$.
+Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<4\epsilon_{(k,0)}$.
+This implies that the orbit of $p$ is $5\epsilon_{(k,0)}$-dense in $X$.
 By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have
 $\mathbb{F}\subset\{x\in X:\omega_{f}(x)=X\}.$
-For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
-Case $(1)$. There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
+Let $p\in\mathbb{F}$. For sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
+Case $(1)$: There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
 Then there is some $z\in\mathcal{I}_{(k,Q_{k})}$ such that $p\in\overline{B}_{M_{(k,Q_{k})}}(z,\frac{\epsilon_{(k,Q_{k})}}{2})$, and
-$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{c_{k,Q_{k}}}{n}D(\mathcal{E}_{c_{k,Q_{k}}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast})$
-$+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast}).$
-Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.14), we have $D(\mathcal{E}_{c_{k,Q_{k}}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$.
+$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast})$
+$+\frac{c_{k,Q_{k}}+m_{(k,Q_{k})}}{n}D(\mathcal{E}_{c_{k,Q_{k}}+m_{(k,Q_{k})}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast}).$
+Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.14), we have $D(\mathcal{E}_{c_{k,Q_{k}}+m_{(k,Q_{k})}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$.
 Thus,
 $D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,Q_{k})}.$
-Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
+Case $(2)$: There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
 Then there is some $z\in\mathcal{I}_{(k,j)}$ such that $p\in\overline{B}_{M_{(k,j)}}(z,\frac{\epsilon_{(k,j)}}{2})$, and
-$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{c_{k,j}}{n}D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
-$+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)}).$
+$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)})$
+$+\frac{c_{k,j}+m_{(k,j)}}{n}D(\mathcal{E}_{c_{k,j}+m_{(k,j)}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
 The similar argument as in Case $(1)$ yields that
 $D(\mathcal{E}_{n}(p),K)\leq\frac{9}{k}+2\delta_{(k,j)}.$
-Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
+Case $(3)$: There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
 One has
 $D($ $\mathcal{E}_{n}(p),\mu_{(k,1)})\leq\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}D(\mathcal{E}_{M_{(k-1,Q_{k-1})^{\ast}}}(p),\mu_{(k,1)})$
-$+\frac{c_{k-1,Q_{k-1}}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})+\frac{c_{k,0}}{n}D(\mathcal{E}_{c_{k,0}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
+$+\frac{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$
+$+\frac{c_{k,0}+m_{(k,0)}}{n}D(\mathcal{E}_{c_{k,0}+m_{(k,0)}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
 $+\frac{n_{(k,1)}+m_{(k,1)}}{n}\sum_{l=0}^{i-1}D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})$
@@
 Note that
-$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
+$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}+m_{(k,0)}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
 $D(\mathcal{E}_{n_{(k,1)}+m_{(k,1)}}(f^{M_{(k,1)}^{l}}(p)),\mu_{(k,1)})\leq\frac{5}{k}+2\delta_{(k,1)},$
-and
-$D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ $\leq D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})$
-$\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
+and $D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ is smaller than
+$D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
 This gives us
 $D(\mathcal{E}_{n}(p),K)\leq\frac{16}{k-1}+4\delta_{(k-1,0)}.$
-Case $(4)$. There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
+Case $(4)$: There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
 In this case, using the expression
-$\mathcal{E}_{n}(z)=$ $\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}}{n}\mathcal{E}_{c_{k,j-1}}(f^{M_{(k,j-1)^{\ast}}}(p))$
+$\mathcal{E}_{n}(z)$ $=\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}+m_{(k,j-1)}}{n}\mathcal{E}_{c_{k,j-1}+m_{(k,j-1)}}(f^{M_{(k,j-1)^{\ast}}}(p))$
 $+\frac{n_{(k,j)}+m_{(k,j)}}{n}\sum_{l=0}^{i-1}\mathcal{E}_{n_{(k,j)}+m_{(k,j)}}(f^{M_{(k,j)}^{l}}(p))+\frac{n-M_{(k,1)}^{i}}{n}\mathcal{E}_{n-M_{(k,1)}^{i}}(f^{M_{(k,1)}^{i}}(p)),$
@@
 Based on the above arguments, one can also show that the sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points.
-Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$.
+Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ for any fixed $(k,j)\in\mathcal{A}$.
 Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$.
@@
 Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying the $\rho$-measure of some appropriate dynamical balls decays exponentially fast along the sequence $\{M_{k}\}$, where $M_{k}:=M_{(k,Q_{k})}$.
-Then we can invoke the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
+Then we can invoke the Entropy Distribution Principle stated in Lemma 2.2 to obtain a lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
 Define
@@
 It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$.
-Thus one has $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$.
+Thus, $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$.
 This implies that $\rho(\mathbb{F})=1$.
-Recall that $M_{(0,0)}=n_{(0,0)}=1$, by (3.12) and (3.13), we have
-#$\#\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
+Recall that $M_{(0,0)}=n_{(0,0)}=1$.
+By (3.12) and (3.13), we have
+$\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
 Lemma 3.5.
-Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball that intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
+Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be a Bowen ball that intersects $\mathbb{F}$. Then, for any $k>k_{0}$, one has
 $\rho_{k}\left(B_{M_{k_{0}}}\left(z,\frac{\epsilon^{\ast}}{2}\right)\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}.$
 Proof of Lemma 3.5.
-Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not=\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also nonempty by the construction of $\mathbb{F}$ and #$\#(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$.
+Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not$$\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also non-empty by the construction of $\mathbb{F}$ and $(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$.
 Notice that $\mathbb{F}_{k_{0}+1}$ descends from $\mathbb{F}_{k_{0}}$.
 Then one has
-#$\#(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
+$(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
 and hence
 $\rho_{k_{0}+1}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})}}{\#\mathcal{I}_{k_{0}+1}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
-In general, for any $k>k_{0}$,
+In general, for any $k>k_{0}$, one has
 $\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{\Pi_{k_{0}+1\leq k^{\prime}\leq k}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}}{\#\mathcal{I}_{k}}=\frac{1}{\#\mathcal{I}_{k_{0}}}.$
-For sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
+By Lemma 3.5, for sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
 $\rho\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)$ $\leq\liminf_{k\to\infty}\rho_{k}\left(B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\right)\leq\frac{1}{\#\mathcal{I}_{k_{0}}}$
@@
 3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
-The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already obtained in [Y23, Lemma 3.3,(3)].
+The following lemma allows us to get an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$.
 Lemma 3.6.
-Let $(X,f)$ be a t.d.s., and let $K\subset M_{f}(X)$ be a nonempty closed subset.
+[Y23, Lemma 3.3,(3)] Let $(X,d,f)$ be a TDS, and let $K\subset M_{f}(X)$ be a non-empty closed subset.
   • (1)
@@
     $h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Thus, by the Lemma 3.6 the upper bound of the Bowen metric mean dimension of saturated set is given by:
+Thus, by Lemma 3.6 we have
 $\overline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in K}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (3.16)
 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
-The proof for the lower bound is similar to that in subsection 3.1.2. So we only give a sketch for the estimation of the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
-We show that for any small enough $\epsilon^{\ast}>0$, it holds that
+The proof for the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$ is similar to that in subsection 3.1.2. So we only give a sketch for the proof.
+For any sufficiently small $\epsilon^{\ast}>0$, we show
 $\inf_{\mu\in K}\inf_{{\rm diam}(\xi)<20\epsilon^{\ast}}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$ (3.17)
@@
 Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$.
-For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination of ergodic $f$-invariant probability measures with rational coefficients such that
+For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination with rational coefficients of $f$-ergodic probability measures such that
 $D(\mu_{(k,i)},\nu_{(k,i)})<\delta_{(k,i)},\ \ \inf_{\xi\succ\mathcal{U}}h_{\mu_{(k,i)}}(f,\xi)\leq\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)+\zeta.$
-Applying Katok’s $\epsilon$-entropy formula and repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
+Repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
 $(1)$ for any $1\leq l\leq p_{k,i}$, $\lambda_{k,i,l}n_{(k,i)}^{\prime}$ is an integer and
@@
 where $n_{(k,i)}:=n_{(k,i)}^{\prime}+(p_{k,i}-1)m_{(k,i)}$;
-$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
-#$\#E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
+$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated satisfying
+$E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
 $(3)$ for any $x\in E_{k,i,l}$,
@@
 where $y(x_{1},\ldots,x_{p_{k,i}})$ is the point that $\epsilon_{(k,i)}$-shadows the pieces of orbits
-[…]
+$\{x_{l},\ldots,f^{\lambda_{k,i,l}n_{(k,i)}^{\prime}}(x_{l})\},\ 1\leq l\leq p_{k,i},$
 with gap $m_{(k,i)}$. Then $\mathcal{C}_{(k,i)}$ is a $(n_{(k,i)},4\epsilon^{\ast})$-separated set with
-#$\#\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
+$\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
 For $k\geq 0$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $N_{(k,0)}$ be defined as in subsection 3.1.2.
@@
 Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$.
-For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
+For every $(k,i)\in\mathcal{A}^{\prime}$ and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
 $\{y_{k,i,l},f(y_{k,i,l}),\ldots,f^{n_{(k,i)}-1}(y_{k,i,l})\},\ l=1,\cdots,N_{(k,i)},$
@@
 Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$.
-Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with the fractal set $\mathbb{F}$, it holds that
+Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every sufficiently large $n$ and $B_{n}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}$, it holds that
 $\rho(B_{n}(z,\frac{\epsilon^{\ast}}{2}))\leq\exp\left[-n(1-\zeta)^{2}(H-2\zeta)\right].$
-Then, by the Lemma 2.1, we have
+Then, by the Lemma 2.1 we have
 $(1-\zeta)^{2}(H-2\zeta)\leq h_{top}^{B}(f,\mathbb{F},\frac{\epsilon^{\ast}}{2})\leq h_{top}^{B}(f,G_{K}^{\omega,U},\frac{\epsilon^{\ast}}{2}).$
@@
 3.2.3 Proof of Theorem 1.3
-With the help of Theorems 1.1 and 1.2, we prove Theorem 1.3.
-Firstly, we show
-$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}.$ (3.18)
-Fix $\mu\in K$ and $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$. Then, by Lemma 3.6, for every $\epsilon>0$,
+The more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$) of Theorems 1.1 and 1.2.
+However, up to now, the authors remain unaware of whether there exist any counter-examples among infinite entropy systems with the specification property for which the following equalities hold strictly, i.e.,
+$\sup_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{P}(f,G_{K}^{\omega,U},d),$
+$\inf_{\mu\in K}\left\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\right\}<\overline{\rm mdim}_{M}^{B}(f,G_{K}^{\omega,U},d).$
+Inspired by the definition of mean Rényi information dimension [GS21], we define two variants of mean Rényi information dimension of invariant measures.
+Let $K$ be a closed set of $M_{f}(X)$. For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology.
+Definition 3.1.
+For any $\mu\in K$, we define the Bowen and packing Rényi information dimensions of $\mu$ as
+$\underline{\rm mrid}_{\mu}^{B}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
+$\overline{\rm mrid}_{\mu}^{P}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
+respectively.
+With the help of Bowen and packing Rényi information dimensions, we establish the variational principles for Bowen and packing metric mean dimensions of saturated sets, whose form are close to the ones of Bowen and packing topological entropies of saturated sets [PS07, ZCC12].
+Now we are ready to prove Theorem 1.3. We first prove the variational principle:
+$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d).$
+Fix $\mu\in K$. For every $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$, by Lemma 3.6 it holds that
 $h_{top}^{B}(f,G_{K},\epsilon)\leq\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi).$
-This shows
+This shows that
 $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)$
-and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\underline{\rm mrid}_{\mu}(f,K,d)$ for every $\mu\in K$.
-On the other hand, we may assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove.
+and hence $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\leq\inf_{\mu\in K}\limits\underline{\rm mrid}_{\mu}^{B}(f,K,d)$.
+To show the reverse inequality
+$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)\geq\inf_{\mu\in K}\limits\underline{\rm mrid}_{\mu}^{B}(f,K,d),$
+we assume that $\underline{\rm mdim}_{M}^{B}(f,G_{K},d)$ is finite; otherwise there is nothing left to prove.
 Let $\gamma>0$.
-Then, by Theorem 1.2, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
-$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
-Assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$. Then $\mu\in K$ since $K$ is closed. This shows that
-$\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
+By Theorem 1.2, we can choose a strictly decreasing sequence $\epsilon_{k}\to 0$ as $k\to 0$ and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for any $k$,
+$\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi)<\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$ (3.18)
+Without loss of generality, we assume that $\mu_{k}\rightarrow\mu$ as $k\to\infty$.
+Then $\mu\in K$ since $K$ is closed.
+For every $\epsilon>0$, if $\epsilon\in(\epsilon_{k+1},\epsilon_{k}]$ for some $k\geq 1$, we define $\nu_{\epsilon}:=\epsilon_{k}$.
+Then $(\nu_{\epsilon})\in M_{K}(\mu)$.
+Using (3.18), we have
+$\underline{\rm mrid}_{\mu}^{B}(f,K,d)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d)+\gamma.$
 Letting $\gamma\to 0$, we have
-$\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
-This completes the proof of (3.18).
-Next we show the second equality:
-$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
-Using formula (3.3), the inequality $\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d)$ is clear.
-On the other hand, let $\gamma>0$ be sufficiently small and put
-$M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$
-Then we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real number that converges to 00 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
+$\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d)\leq\underline{\rm mdim}_{M}^{B}(f,G_{K},d).$
+Next we show the variational principle:
+$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)=\sup_{\mu\in K}\limits\overline{\rm mrid}_{\mu}^{P}(f,K,d).$
+Using (3.3), there exists a constant $c>0$ for every sufficiently small $\epsilon>0$,
+$\sup_{\mu\in K}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{P}(f,G_{K},c\epsilon).$
+This yields the inequality
+$\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d)\leq\overline{\rm mdim}_{M}^{P}(f,G_{K},d).$
+On the other hand, let $\gamma>0$ and put $M_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{P}(f,G_{K},d)-\gamma\}.$ By Theorem 1.1, we can choose a sequence $\{\epsilon_{k}\}_{k}$ of positive real numbers that converges to 0 and a sequence $\{\mu_{k}\}$ of measures in $K$ such that for sufficiently large $k$,
 $M_{\gamma}<\frac{1}{|\log\epsilon_{k}|}\inf_{\mathrm{diam}(\xi)<\epsilon_{k}}h_{\mu_{k}}(f,\xi).$
-Again, we assume that $\mu_{k}\rightarrow\mu\in K$. This shows that
-$M_{\gamma}\leq\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\}.$
-Letting $\gamma\to 0$, we have
-$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\{\overline{\rm mrid}_{\mu}(f,K,d)\}.$
+By the similar arguments, we have
+$\overline{\rm mdim}_{M}^{P}(f,G_{K},d)\leq\sup_{\mu\in K}\overline{\rm mrid}_{\mu}^{P}(f,K,d).$
 3.3 Upper capacity metric mean dimension of saturated sets
-In this subsection, we prove the Theorem 1.4.
-By Proposition 2.1, the inequality
+In this subsection, we give the proof of Theorem 1.4.
+By Proposition 2.1, it is clear that
 $\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K},d)\leq\overline{\rm mdim}_{M}(f,X,d)$
@@
 $\overline{\rm mdim}_{M}(f,X,d)\leq\overline{\rm mdim}_{M}^{UC}(f,G_{K}^{\omega,U},d).$
-In light of Theorem 2.4, we need to show that for $\epsilon^{\ast}>0$, $\zeta>0$ sufficiently small and $\delta\in(0,1)$, one has
+In light of Theorem 2.4, we need to show that for every sufficiently small $\epsilon^{\ast}>0$, $\zeta>0$ and $\delta\in(0,1)$, one has
 $\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta\leq h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast}).$ (3.19)
-However, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
+Here, we will give a sketch for the proof of (3.19) since it is essentially analogous to that of the lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$ presented in subsection 3.1.2.
 Write $H=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)$ and assume that $H>0$. Take $\mu^{\ast}\in M_{f}^{e}(X)$ such that
@@
 Then, for sufficiently large $\mathcal{N}\in\mathbb{N}$, there exists a $(\mathcal{N},3\epsilon^{\ast})$-separated subset $S_{\mathcal{N}}$ with
-#$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
-Now fix such $\mathcal{N}$, and write $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
+$S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)].$
 Since $K$ is a connected closed subset, a standard argument shows that there exists a family $\{\mu_{k,i}:k\geq 1,1\leq i\leq Q_{k}\}$ of measures in $K$ such that for all $k\geq 1$, $1\leq i<Q_{k}$,
@@
 Similarly, the sequences $\{\epsilon_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{\delta_{(k,j)}\}_{(k,j)\in\mathcal{A}}$, $\{m_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ are constructed as in subsection 3.1.2.
-For the nonempty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with $\overline{B(a_{0},\epsilon_{0})}\subset U$.
-Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$.
+For the non-empty open set $U$, there exist $a_{0}\in X$ and $\epsilon_{0}\in(0,\epsilon^{\ast})$ with ${\overline{B}(a_{0},\epsilon_{0})}\subset U$.
+Set $\mathcal{C}_{(-1,0)}=\{a_{0}\}$, $n_{(-1,0)}=c_{(-1,0)}=N_{(-1,0)}=1$, and $\mathcal{C}_{(0,0)}=S_{\mathcal{N}}$, $c_{(0,0)}=\mathcal{N}$, $N_{(0,0)}=1$.
 For $k\geq 1$, let $\mathcal{C}_{(k,0)}$, $n_{(k,0)}$, $c_{(k,0)}$ and $N_{(k,0)}$ be the same as in subsection 3.1.2.
@@
 $\begin{split}&n_{(k,j)_{\ast}}+m_{(k,j)_{\ast}}\leq\frac{1}{k}\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right],\\ &\sum_{(k^{\prime},j^{\prime})\leq(k,j)}\left[N_{(k^{\prime},j^{\prime})}\cdot(n_{(k^{\prime},j^{\prime})}+m_{(k^{\prime},j^{\prime})})\right]\leq\frac{N_{(k,j)_{\ast}}\cdot n_{(k,j)_{\ast}}}{k}.\end{split}$
-Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain the fractal set
+Based on the data $\mathcal{C}_{(k,j)}$, $N_{(k,j)}$ for $(k,j)\in\mathcal{A}$, similar to the previous subsections, one can construct the intermediate sets $\mathcal{S}_{(k,j)}$, $\mathcal{I}_{(k,j)}$, $\mathbb{F}_{(k,j)}$ to obtain a fractal set
 $\mathbb{F}=\bigcap_{(k,j)\in\mathcal{A}}\mathbb{F}_{(k,j)}\subset G_{K}^{\omega,U}.$
-It is worth mentioning that for any $(k,j)>(0,0)$, #$\#\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
+It is worth mentioning that for any $(k,j)>(0,0)$, $\mathcal{I}_{(k,j)}=\#S_{\mathcal{N}}$. Moreover, for any distinct $x$, $x^{\prime}\in\mathcal{I}_{(0,0)}$, one has
 $d_{M_{(0,0)}}(x,x^{\prime})>2\epsilon^{\ast}.$
@@
 $d_{M_{(0,0)}}(y,y^{\prime})>d_{M_{(0,0)}}(x,x^{\prime})-d_{M_{(0,0)}}(x,y)-d_{M_{(0,0)}}(x^{\prime},y^{\prime})>\epsilon^{\ast}.$
-Hence, for sufficiently large $\mathcal{N}$, there is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset in $\mathbb{F}$ with cardinality #$\#S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$.
-Thus
+Hence, $\mathbb{F}$ is a $(1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N},\epsilon^{\ast})$-separated subset with $S_{\mathcal{N}}\geq\exp[\mathcal{N}(H-2\zeta)]$. Thus,
 $h_{top}^{UC}(f,G_{K}^{\omega,U},\epsilon^{\ast})\geq\limsup_{\mathcal{N}\to\infty}\frac{\mathcal{N}(H-2\zeta)}{1+m(\frac{\epsilon^{\ast}}{8})+\mathcal{N}}=\sup_{\mu\in M_{f}^{e}(X)}\underline{h}_{\mu}^{K}(f,3\epsilon^{\ast},\delta)-2\zeta.$
-This implies the inequality (3.19).
-3.4 An extension to metric mean dimensions of sup sets and sub sets
+This implies the inequality (3.19) by letting $\zeta\to 0$.
+3.4 Metric mean dimensions of sup sets and sub sets
 In this section, we consider two variants of saturated sets, namely sup sets and sub sets, and establish variational principles for metric mean dimensions of sup sets and sub sets.
-The main results are presented in Theorems 3.7, 3.8 and 3.9.
-In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the sense of weak-topology.
+In general, given a point $x\in X$, the limit $\lim_{n\to\infty}\mathcal{E}_{n}(x)$ may not exist in the weak-topology.
 The set of such points is denoted by $I(f)$, known as the historic set [^1] This terminology is due to Rulle’s original paper [Rue01].
 of $f$.
-By Birkhoff ergodic theorem, although the historic set is usually regarded as negligible in the measure-theoretic context, its topological dynamics can be quite rich.
-Especially for systems with specification-like properties, it can even have full topological entropy [CTS05] and full metric mean dimension [LV21, LL24].
 To understand the underlying geometry structure of $I(f)$, we define the following sets based on the framework introduced and further developed by Olsen [O02, O03, O06, O08] and Olsen and Winter [OW03].
-More precisely, given a nonempty subset $C\subset M_{f}(X)$, we define the sup set, equ set [^2] In this context, we do not take into account the deformation of $\mathcal{E}_{n}(x)$ via a continuous affine map from $M_{f}(X)$ to a topological vector space $Y$.
-Hence, we prefer to use the term “saturated set” rather than “equ set”.
-and sub set of $C$ as
-${}^{C}G:$ =$=\{x\in X:V_{f}(x)\supset C\},$
-$G_{C}:$ =$=\{x\in X:V_{f}(x)=C\},$
-$G^{C}:$ =$=\{x\in X:V_{f}(x)\subset C\},$
-respectively.
-Note that the equ set is precisely the saturated set associated with $C$ defined previously. More generally, by using two nonempty subsets $C_{1},C_{2}\subset M_{f}(X)$, we define the set
+More precisely, given a non-empty subset $C\subset M_{f}(X)$, we define the sup set and sub set of $C$ as
+$^{C}G:$ $=\{x\in X:V_{f}(x)\supset C\},$
+$G^{C}:$ $=\{x\in X:V_{f}(x)\subset C\},$
+respectively. More generally, for two non-empty subsets $C_{1},C_{2}\subset M_{f}(X)$, we can define
 $G(C_{1},C_{2}):=\{x\in X:C_{1}\subset V_{f}(x)\subset C_{2}\},$
-which covers the three different notions above in an unified manner.
-Now, we are ready to establish variational principles for the three given types of level sets $G^{C},^{C}G,G(C_{1},C_{2})$ by means of the Bowen metric mean dimension.
+to recover the above notions in an unified manner.
+Now, we are ready to establish variational principles for $G^{C},^{C}G$.
 Theorem 3.7.
-Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty subset $C\subseteq M_{f}(X)$,
+Let $(X,d,f)$ be a TDS with the specification property. Then for any non-empty subset $C\subseteq M_{f}(X)$,
 $\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
 We claim that for every $\epsilon>0$,
-$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
-where $\overline{{\rm Conv}}(C)$ denotes the convex closure of set $C$.
-Indeed, since $C\subset\overline{{\rm Conv}}(C)$, it suffices to show
-$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm Conv}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Let $\mu\in\overline{{\rm Conv}}(C)$.
-We can choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $\lambda_{n}^{(j)}>0$ and $\sum_{j=o}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$.
+$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\inf_{\mu\in\overline{{\rm co}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$ (3.20)
+where $\overline{{\rm co}}(C)$ denotes the closure of the convex hull of $C$.
+Indeed, since $C\subset\overline{{\rm co}}(C)$, it suffices to show
+$\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\inf_{\mu\in\overline{{\rm co}}(C)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
+Let $\mu\in\overline{{\rm co}}(C)$.
+Choose $\mu_{n}=\sum_{j=1}^{k_{n}}\lambda_{n}^{(j)}v_{n}^{(j)}$ with $v_{n}^{(j)}\in C$, $0<\lambda_{n}^{(j)}<1$ and $\sum_{j=0}^{k_{n}}\lambda_{n}^{(j)}=1$ such that $\mu_{n}\rightarrow\mu$ as $n\to\infty$.
 The upper semi-continuity of $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ on $M_{f}(X)$ (cf.
@@
 $\limsup_{n\to\infty}~{\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu_{n}}(f,\xi)}\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Combining this and $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is concave in $\mu$, we can deduce that
-$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma$
-holds for any for $\gamma>0$. This yields the desired inequality and hence (3.20) holds.
-Notice that $G_{\overline{{\rm Conv}}(C)}\subset^{C}G$, we have
-$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C)},d\right)$
+Since $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ is a concave function on $M_{f}(X)$, for every $\gamma>0$ we can deduce that
+$\inf_{\nu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(f,\xi)<\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)+\gamma.$
+This yields the desired inequality (3.20).
+Notice that $G_{\overline{{\rm co}}(C)}\subset^{C}G$, we have
+$\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm co}}(C)},d\right)$
 […]
 $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-Conversely, for any $\mu\in C$ and $\epsilon>0$, by Lemma 3.6 one has
-$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-This shows $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$ for every $\epsilon>0$. ∎
+Fix $\mu\in C$ and $\epsilon>0$. By Lemma 3.6, one has
+$h_{top}^{B}(f,^{C}G,\epsilon)\leq h_{top}^{B}(f,\{x\in X:\mu\in V_{f}(x)\},\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+and hence $h_{top}^{B}(f,^{C}G,\epsilon)\leq\inf_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$. ∎
 Theorem 3.8.
-Let $(X,d,f)$ be a t.d.s. with specification property. Then for any nonempty closed subset $C\subseteq M_{f}(X)$, we have
+Let $(X,d,f)$ be a TDS with the specification property. Then for any non-empty closed subset $C\subseteq M_{f}(X)$, we have
 $\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 Proof.
-For every $\epsilon>0$, by Lemma 3.6 again we have
+For every $\epsilon>0$, by Lemma 3.6 we have
 $h_{top}^{B}(f,G^{C},\epsilon)\leq h_{top}^{B}(f,\{x\in X:V_{f}(x)\cap{C}\not=\emptyset\},\epsilon)\leq\sup_{\mu\in{C}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-By (3.17), there exists a constant $c>0$ such that
-$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon)$
-holds for all $\mu\in C$ and any sufficiently small $\epsilon>0$. Combining with the fact that $G_{\mu}\subset G^{C}$ holds for each $\mu\in C$, one gets
+By (3.17), there exists a constant $c>0$ such that for all $\mu\in C$ and any sufficiently small $\epsilon>0$,
+$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon).$
+Combining this inequality with the fact that $G_{\mu}\subset G^{C}$ for each $\mu\in C$, we have
 $\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G^{C},\epsilon).$
 Theorem 3.9.
-Let $(X,d,f)$ be a t.d.s.
-with specification property.
-Suppose that $C_{1}\subseteq M_{f}(X)$ is a nonempty set satisfying that $\overline{{\rm Conv}}(C_{1})$ is a connected component of $C_{2}\subset M_{f}(X)$.
-Then
+Let $(X,d,f)$ be a TDS with the specification property.
+Let $C_{1}\subseteq M_{f}(X)$ be a non-empty set and $C_{1}\subseteq M_{f}(X)$.
+If $\overline{{\rm co}}(C_{1})$ is a connected component of $C_{2}$, then
 $\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
 Proof.
-Since $C_{1}\subset\overline{{\rm Conv}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
-$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm Conv}}(C_{1})},d\right)$
-$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm Conv}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
+Since $C_{1}\subset\overline{{\rm co}}(C_{1})\subset C_{2}$, by Theorem 1.2 we have
+$\overline{\rm mdim}_{M}^{B}\left(f,G(C_{1},C_{2}),d\right)$ $\geq\overline{\rm mdim}_{M}^{B}\left(f,G_{\overline{{\rm co}}(C_{1})},d\right)$
+$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in\overline{{\rm co}}(C_{1})}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)$
 $=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mu\in C_{1}}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
 Corollary 3.10.
-Let $(X,d,f)$ be a t.d.s. with specification property.
+Let $(X,d,f)$ be a TDS with the specification property.
   • (1)
-    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
+    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
     $\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)=\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right).$
   • (2)
-    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
+    If $C$ is a non-empty compact convex set of $M_{f}(X)$, then
     $\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right).$
 Proof.
-(1) follows by Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
+(1) follows from Theorems 1.2 and 3.7; (2) is due to Theorems 1.1 and 3.8. ∎
 Corollary 3.11.
-Let $(X,d,f)$ be a t.d.s. with specification property.
+Let $(X,d,f)$ be a TDS with the specification property.
   • (1)
-    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
+    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
     $\overline{\rm mdim}_{M}^{B}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{B}\left(f,^{C}G,d\right)\leq\overline{\rm mdim}_{M}^{B}\left(f,G^{C},d\right).$
   • (2)
-    If $C$ is a nonempty compact convex set of $M_{f}(X)$, then
+    If $C$ is a non-empty compact convex set of $M_{f}(X)$, then
     $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,^{C}G,d\right).$
   • (3)
-    If $C$ is a nonempty compact connected set of $M_{f}(X)$, then
+    If $C$ is a non-empty compact connected set of $M_{f}(X)$, then
     $\overline{\rm mdim}_{M}^{UC}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,G^{C},d\right)=\overline{\rm mdim}_{M}^{UC}\left(f,^{C}G,d\right).$
 Proof.
-$(1)$ is clear. For $(2)$, it follows from Theorem 1.1 that
+$(1)$ is clear.
+$(2)$. It follows from Theorem 1.1 that
 $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in C}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)=\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)\leq\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right).$
@@
 Hence, we have $\overline{\rm mdim}_{M}^{P}\left(f,G_{C},d\right)=\overline{\rm mdim}_{M}^{P}\left(f,G^{C},d\right)$.
-$(3)$ follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
-4 Some applications of main results
-In this section, we apply the main results to multifractal analysis of level sets and mean Li-Yorke chaos in infinite entropy systems.
-The main results in this section are presented in Theorems 4.1 and 4.3.
-4.1 Variational principle of level sets
-In this subsection, we establish variational principles for level sets of continuous functions on $M_{f}(X)$ within the context of an abstract framework.
-Given a dynamical system $(X,f)$, assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function. We define
-$L_{\varphi}=\left[\inf_{\mu\in M_{f}(X)}\varphi(\mu),\ \sup_{\mu\in M_{f}(X)}\varphi(\mu)\right].$
-For any $\alpha\in L_{\varphi}$, we define $M_{f}(\varphi,\alpha):=\left\{\mu\in M_{f}(X):\varphi(\mu)=\alpha\right\}$.
-The continuity of $\varphi$ then implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
-We define the level set of $\alpha$ w.r.t.
-$\varphi$ as
-$K_{\alpha}:=\left\{x\in X:\inf_{\mu\in V_{f}(x)}\varphi(\mu)=\sup_{\mu\in V_{f}(x)}\varphi(\mu)=\alpha\right\},$
-which consists of the set of points of $X$ such that $\varphi$ takes constant value $\alpha$ on $V_{f}(x)$.
-Condition A: For every $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
-Several examples of the function $\varphi$ that satisfy Condition A are stated below:
-  • (a)
-    If $\varphi=\alpha\in\mathbb{R}$ is a constant, then $L_{\varphi}=\{\alpha\}$, $K_{\alpha}=X$, $M_{f}(\varphi,\alpha)=M_{f}(X)$;
-  • (b)
-    We define
-    $\varphi(\mu):=\frac{\int\psi\mathop{}\!\mathrm{d}\mu}{\int\phi\mathop{}\!\mathrm{d}\mu},$
-    where $\psi$, $\phi\in C(X)$ and $\int\phi\mathop{}\!\mathrm{d}\mu\neq 0$ for all $\mu\in M_{f}(X)$ (e.g. $\phi>0$).
-    Then for any $\alpha\in L_{\varphi}$,
-    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\},$
-    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\int\phi\mathop{}\!\mathrm{d}\mu\};$
-  • (c)
-    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, that is, for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$.
-We define
-    $\varphi:M_{f}(X)\to\mathbb{R},~{}\mu\mapsto\lim_{n\to\infty}\int\frac{\phi_{n}(x)}{n}\mathop{}\!\mathrm{d}\mu(x).$
-    Then $\varphi$ is continuous (cf.
-[FH10, Lemma A.4]), and $M_{f}(\varphi,\alpha)$ is a convex subset for all $\alpha\in L_{\Phi}$.
-Specially, let $\psi\in C(X)$ and define $\phi_{n}=S_{n}\psi$, then $\Phi=\{\phi_{n}\}$ is an additive potential, in this case we have
-    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\psi\mathop{}\!\mathrm{d}\mu=\alpha\},$
-    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
-Given a nonempty open subset $U\subset X$, we similarly define the mixed $\alpha$-level set
-$K_{\alpha}^{\omega,U}=K_{\alpha}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}.$
+$(3)$. It follows from Theorem 1.4 and the relations $G_{C}\subset G^{C}$ and $G_{C}\subset^{C}G$. ∎
+4 The applications of main results
+In this section, we apply the main results to study the metric mean dimension of level sets, mean Li-Yorke chaos in infinite entropy systems and the metric mean dimension of the set of generic points of full shifts over compact metric spaces.
+4.1 Metric mean dimensions of level sets
+Let $(X,d,f)$ be a TDS and $\psi$ be a continuous function on $X$. If $\mu$ is a $f$-invariant probability measure on $X$, the Birkhoff ergodic theorem verifies that for $\mu$-a.e. $x\in X$, the limit
+$\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)$
+exists.
+However, given $\alpha\in\mathbb{R}$ it does not give more information about the “size” of the set of the points of $X$ that the time average of $\psi$ equals $\alpha$.
+For this reason, the $\alpha$-level set of $\psi$ is defined by
+$K_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}.$
+In [BR23], the authors proved that if $K_{\alpha}$ is non-empty and the TDS $(X,d,f)$ has the specification property, then
+$\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+where $M_{f}(\psi,\alpha):=\{\mu\in M_{f}(X):\int\psi d\mu=\alpha\}$.
+Actually, we are able to prove this variational principle by an alternative approach and give new characterization for the (packing/upper capacity) metric mean dimension of level sets.
 Theorem 4.1.
-Let $(X,d,f)$ be a t.d.s.
-with the specification property, and let $U\subset X$ be a nonempty open subset.
-Assume that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A.
-Then for any $\alpha\in L_{\varphi}$ and $S\in\{B,P\}$,
-$\overline{\rm mdim}_{M}^{S}(f,K_{\alpha}^{\omega,U},d)=$ $\overline{\rm mdim}_{M}^{S}(f,K_{\alpha},d)$
-$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
-[…]
+Let $(X,d,f)$ be a TDS with the specification property and $\psi$ be a continuous function on $X$. If $\alpha\in\mathbb{R}$ is a constant such that $K_{\alpha}\not$$\emptyset$, then
+$\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)$ $=\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
+$\overline{\rm mdim}_{M}^{UC}(f,K_{\alpha},d)$ $=\overline{\rm mdim}_{M}(f,X,d).$
+The corresponding result are also valid for the three types of lower metric mean dimensions by changing $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
 Proof.
-Define $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.17), there exists a constant $c>0$ such that
-$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu}^{\omega,U},\epsilon)$
-holds for all $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$. Together with the fact that $G_{\mu}^{\omega,U}\subset K_{\alpha}^{\omega,U}$ holds for any $\mu\in M_{f}(\varphi,\alpha)$, we have
-$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha}^{\omega,U},\epsilon)$
-for all sufficiently small $\epsilon>0$. This implies that
-$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha}^{\omega,U},d).$ (4.1)
+If $K_{\alpha}\not$$\emptyset$, then $M_{f}(\psi,\alpha)$ is a non-empty convex set.
+For $x\in X$, the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$.
 Recall $G^{M_{f}(\varphi,\alpha)}=\{x\in X:V_{f}(x)\subset M_{f}(\varphi,\alpha)\}$.
-Fix $x\in X$.
-Since the statement $\varphi(\mu)=\alpha$ for all $\mu\in V_{f}(x)$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\varphi,\alpha)$, it follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$.
+It follows that $K_{\alpha}=G^{M_{f}(\varphi,\alpha)}$.
 By Lemma 3.2 $(1)$, we get
 $\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)\leq\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$ (4.1)
-By (4.1), (4.2), and the fact stated in Proposition 2.1 $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)$ for any $Z\subset X$, we complete the proof. ∎
-A direct consequence of Theorem 4.1 is the variational principles for the Bowen metric mean dimension of level sets, as derived by Backes and Rodrigues [BR23, Theorem A].
-However, we are able to strengthen this result by demonstrating that the variational principle remains valid for the packing metric mean dimension of level sets [^3] Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure..
-Corollary 4.2.
-Let $(X,d,f)$ be a t.d.s. with the specification property and $\psi$ be a continuous function on $X$. Let
-$E_{\alpha}:=\{x\in X:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha\}$
-denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then
-$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\psi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
-[…]
-Proof.
-[…]
-$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
-Fix $x\in X$.
-Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$.
-We obtain the desired result by Theorem 4.1.
+By inequality (3.17), there exists a constant $c>0$ such that for every $\mu\in M_{f}(X)$ and sufficiently small $\epsilon>0$,
+$\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,G_{\mu},\epsilon).$
+Since $G_{\mu}\subset K_{\alpha}$ for any $\mu\in M_{f}(\varphi,\alpha)$, we have
+$\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<c\epsilon}h_{\mu}(f,\xi)\leq h_{top}^{B}(f,K_{\alpha},\epsilon),$
+which implies that
+$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi)\leq\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d).$ (4.2)
+By (4.1), (4.2) and $\overline{\rm mdim}_{M}^{B}(f,K_{\alpha},d)\leq\overline{\rm mdim}_{M}^{P}(f,K_{\alpha},d)$, we get the first equalities.
+Fix $\mu\in M_{f}(\psi,\alpha)$.
+Since the specification property implies $G_{\mu}\not$$\emptyset$ [PS07], we have $G_{\mu}\subset K_{\alpha}$.
+By Corollary 1.5, we obtain $\overline{\rm mdim}_{M}^{UC}(f,K_{\alpha},d)=\overline{\rm mdim}_{M}(f,X,d)$.
 ∎
-4.2 Linking metric mean dimension and mean Li-Yorke chaotic
-In the present subsection, we establish a connection between the concepts of metric mean dimension and mean Li-Yorke chaos within the context of infinite entropy.
-We provide a quantitative description of the chaotic behavior exhibited by infinite entropy systems.
-Specifically, we demonstrate that the packing and upper capacity metric mean dimensions of the set of mean Li-Yorke pairs are equal to twice the metric mean dimension of the phase space, as stated in Theorem 4.3.
-Given a t.d.s. $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric defined as
+Example 4.2.
+Endow the product space $0,1]^{\mathbb{Z}}$ with the product metric:
+$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
+and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})=(x_{n+1})_{n\in\mathbb{Z}}$.
+It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf.
+[DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18, E.
+Example].
+Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map given by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$.
+Then for any $\alpha\in[0,1]$, by Theorem 4.1 we have
+$\overline{\rm mdim}_{M}^{B}(\sigma,K_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,K_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
+where $K_{\alpha}:=\{x\in[0,1]^{\mathbb{Z}}:\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}x_{j}=\alpha\}$, and $M_{\sigma}(\varphi,\alpha):=\{\mu\in M_{\sigma}([0,1]^{\mathbb{Z}}):\int\varphi d\mu=\alpha\}$.
+Furthermore, we have
+$\overline{\rm mdim}_{M}^{UC}(\sigma,K_{\alpha},d)=1.$
+4.2 Mean Li-Yorke chaos in infinite entropy systems
+In this subsection, we provide a quantitative result involving the chaotic behavior of infinite entropy systems.
+Specifically, we prove that the (packing/upper capacity) metric mean dimension of the set of mean Li-Yorke pairs is equal to twice the metric mean dimension of the phase space.
+Given a TDS $(X,d,f)$, this naturally induces a product system $(X\times X,d\times d,f\times f)$, where $X\times X$ is a compact space equipped with the product metric:
 $d\times d((x_{1},y_{1}),(x_{2},y_{2}))=\max\{d(x_{1},x_{2}),d(y_{1},y_{2})\},$
-and the mapping $f\times f:X\times X\rightarrow X\times X$, given by $(f\times f)(x,y):=(fx,fy)$, is continuous.
-A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
+and $(f\times f)(x,y):=(fx,fy)$. A pair $(x,y)\in X\times X$ is called a mean Li-Yorke pair if
 $\liminf_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)=0,~\text{and}\ \limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}d(f^{i}x,f^{i}y)>0.$
-The set of all mean Li-Yorke pairs is denoted by ${\rm MLY}(f)$.
-A subset $S$ of $X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair.
-We say that the system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set.
-The well-known fact is that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]).
+Namely, the orbits of $x$ and $y$ is sufficiently close to each other for sufficiently $n$, but they may be apart from each other for infinitely many time.
+Denote by ${\rm MLY}(f)$ the set of all mean Li-Yorke pairs.
+A subset $S\subset X$ is called a mean Li-Yorke set if any two distinct points in $S$ form a mean Li-Yorke pair.
+The system $(X,f)$ is mean Li-Yorke chaotic if there is an uncountable mean Li-Yorke set.
+A well-known fact is that positive entropy systems are mean Li-Yorke chaotic [Dow14].
 Theorem 4.3.
-Let $(X,d,f)$ be a t.d.s. with specification property. Then for $S\in\{UC,P\}$ we have
+Let $(X,d,f)$ be a TDS with the specification property. Then for $S\in\{UC,P\}$, we have
 $\overline{\rm mdim}_{M}^{S}(f\times f,{\rm MLY}(f),d\times d)=2~\overline{\rm mdim}_{M}(f,X,d)$
@@
 Proof.
-It suffices to show the first equality since the second one is obtained similarly.
-We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0$.
+It suffices to prove the first equality since the second one can be obtained in a similar manner.
+We claim that there exists $\mu_{0}\in M_{f\times f}(X\times X)$ such that
+$\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{0}(x,y)>0.$
 Notice that the specification property implies that the system has positive topology entropy.
-Since positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that
-$\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$
-Set $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$.
-Then, in the weak-topology any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
-Now take $\nu\in M_{f}(X)$ arbitrarily and put $\mu_{1}:=\phi_{\ast}\nu=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$.
-It is easy to check that $\mu_{1}\in M_{f\times f}(X\times X)$, and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$.
-For sufficiently small $\gamma>0$, let
-$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)-\gamma\}.$
-By Theorem 2.4, we pick a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and a sequence of invariant measures $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
-$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$ (4.3)
-Let $K=\overline{{\rm Conv}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}$, i.e., the convex closure of the set $\{\mu_{0},\mu_{1}\}\cup\{m_{k}:k\geq 1\}$.
+Since positive entropy systems are mean Li-Yorke chaotic [Dow14], we choose two distinct points $x,y\in X$, a sequence $\{n_{i}\}$ of positive integers that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta.$ Put $\mu_{n_{i}}:=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$.
+Then, in the weak-topology, any accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim.
+Fix $\nu\in M_{f}(X)$ and put $\mu_{1}:=\nu\circ\phi^{-1}$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$.
+It is easy to see that $\mu_{1}\in M_{f\times f}(X\times X)$ and $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$.
+Fix $\gamma>0$ and let
+$c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}.$
+By Theorem 2.4, we choose a sequence of positive real numbers $\epsilon_{k}\to 0$ as $k\to\infty$ and a sequence $\{m_{k}\}$ of $f\times f$-invariant measures such that for every $k$,
+$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)>c_{\gamma}.$
+Let
+$K=\overline{{\rm co}}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}.$
 Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$.
-It is easy to see that if two systems admit the specification property, so does the product system.
-By [PS07], we have $G_{K}\not=\emptyset$.
-For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}_{n\in\mathbb{N}}$ and $\{t_{n}^{1}\}_{n\in\mathbb{N}}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively.
+Since $(X\times X,d\times dmf\times f)$ has the specification property, by [PS07] we have $G_{K}\not$$\emptyset$.
+For any $(x_{0},y_{0})\in G_{K}$, there exists two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ such that $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$, respectively.
 Hence,
@@
 $\lim_{n\to\infty}\frac{1}{t_{n}^{1}}\sum_{i=0}^{t_{n}^{1}-1}d(f^{i}x_{0},f^{i}y_{0})=\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0.$
-This implies that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1, one has
+This yields that $G_{K}\subset{\rm MLY}(f)$. Therefore, by Theorem 1.1 one has
 $\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq$ $\overline{\rm mdim}_{M}^{P}(f\times f,G_{K},d\times d)$
-$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
-$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)}{|\log\epsilon_{k}|}\geq c_{\gamma}.$
-Since $\gamma$ is arbitrary, one has
-$\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
+$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\nu\in K}\inf_{{\rm diam}(\eta,d\times d)<\epsilon}h_{\nu}(f\times f,\eta)$
+$\geq$ $\limsup_{k\to\infty}\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f\times f,\xi)\geq c_{\gamma}.$
+Letting $\gamma\to 0$, one has
+$\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)\leq$ $\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)$
 $\leq$ $\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)$
-$\leq$ $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d),$
-where we used the Proposition 2.1 for the second inequality.
-Together with $\overline{\rm mdim}_{M}^{UC}(f\times f,X\times X,d\times d)=2~{}\overline{\rm mdim}_{M}(f,X,d)$, we obtain the desired equality.
-∎
-Finally, we give an example to verify these results presented in Corollary 4.2 and Theorem 4.3.
-Example 4.4.
-Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric:
-$d((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}})=\sum_{n\in\mathbb{Z}}\frac{|x_{n}-y_{n}|}{2^{|n|}},$
-and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})=(x_{n+1})_{n\in\mathbb{Z}}$.
-It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
-  • (1)
-    Thus, by Theorem 4.3, we have
+$\leq$ $\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)$
+$=$ $2~\overline{\rm mdim}_{M}(f,X,d).$
+This completes the proof. ∎
+For instance, under the setting of Example 4.2 we have
 $\overline{\rm mdim}_{M}^{P}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
-  • (2)
-    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be a continuous map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, by Corollary 4.2 we have
-    $\overline{\rm mdim}_{M}^{B}(\sigma,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
-[…]
 4.3 Generic points of full shift over compact metric space
-In this subsection, based on the main results, we present some specific formulae for the metric mean dimensions of the set of generic points of full shifts over compact metric spaces.
-We begin by reviewing some definitions and results for our statements. Let $(X,d,f)$ be a t.d.s., $\mu\in M_{f}(X)$, $\epsilon>0$ and $x\in X$. We define the Brin-Katok local entropy at $x$ as
-$\overline{h}_{\mu}^{BK}(x,\epsilon)=\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n},$
-and the Brin-Katok local $\epsilon$-entropy of $\mu$ as
-$\overline{h}_{\mu}^{BK}(f,\epsilon):=\int\overline{h}_{\mu}^{BK}(x,\epsilon)\mathop{}\!\mathrm{d}\mu,$
-respectively. Specially, if $\mu\in M_{f}^{e}(X)$, then $h_{\mu}^{BK}(x,\epsilon)=h_{\mu}^{BK}(f,\epsilon)$ holds for $\mu$-a.e. $x\in X$.
-Example 4.5.
-Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
-$d^{\mathbb{Z}}\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d(x_{n},y_{n})}{2^{|n|}}.$
-It is well-known that the system $(K^{\mathbb{Z}},\sigma)$ has the specification property. Moreover, it was proved in [VV17, Theorem 5] and [S22, Theorem 15] that,
-$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d),$ (4.4)
-where $\overline{\rm dim}_{B}(K,d)$ is the upper box dimension of $K$ under the metric $d$.
-Then, by Corollary 1.5, the metric mean dimensions of the generic points of the full shift over compact metric space are fully determined by the boxing dimension of the underlying space and mean Rényi information dimension of invariant measures.
+For the full shifts over compact metric spaces, we present the precise formulae for the metric mean dimension of the set of generic points of invariant measures.
+Let $(K,d_{K})$ be a compact metric space. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with the following product metric
+$d\left((x_{n})_{n\in\mathbb{Z}},(y_{n})_{n\in\mathbb{Z}}\right)=\sum_{n\in\mathbb{Z}}\frac{d_{K}(x_{n},y_{n})}{2^{|n|}}.$
+It is well-known that the system $(K^{\mathbb{Z}},d,\sigma)$ has the specification property. It was proved in [VV17, Theorem 5] that,
+$\overline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d)$ $=\overline{\rm dim}_{B}(K,d_{K}),$ (4.3)
+$\underline{\rm mdim}_{M}(\sigma,K^{\mathbb{Z}},d)$ $=\underline{\rm dim}_{B}(K,d_{K}),$ (4.4)
+where $\overline{\rm dim}_{B}(K,d)$ and $\underline{\rm dim}_{B}(K,d)$ denote the upper and lower box dimensions of $K$ under the metric $d$, respectively.
+By Corollary 1.5, for the full shift over compact metric spaces the metric mean dimensions of the set of generic points of invariant measures are fully determined by the box dimension of the underlying space and mean Rényi information dimension of invariant measures.
 Theorem 4.4.
-Under the setting of Example 4.5, for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ one has
+Let $(K^{\mathbb{Z}},d,\sigma)$ be the full shift over $(K,d_{K})$. Then for any $\mu\in M_{\sigma}(K^{\mathbb{Z}})$,
 $\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K,d)$
 $\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
-Next, we present two examples of full shifts with the maximal metric mean dimension measure.
-Example 4.7.
-Consider the Example 4.4 again.
-Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$, and define $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ as the product measure on $[0,1]^{\mathbb{Z}}$.
-Shi proved that $\lim_{\epsilon\to 0}\frac{h_{\mu}^{BK}(\sigma,\epsilon)}{|\log\epsilon|}=1$ [S22, Example 12].
-Together with [YCZ25, Theorem 1.1], we have
-$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d).$
-Then, by Corollary 1.5, we have
-${\rm\overline{mdim}}_{M}^{B}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{P}(\sigma,G_{\mu},d)={\rm\overline{mdim}}_{M}^{UC}(\sigma,G_{\mu},d)=1.$
-Example 4.8.
-Endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$, and let $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ $1\leq i\leq l$ be contracting similitudes, that is,
-$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}$
-for some $r_{i}\in(0,1)$.
-One says that the iterated function system $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there is a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset,\ \text{for }i\neq j.$
-Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure such that
+Next, for certain full shift over self-similar set we show the metric mean dimension of the product of the self-similar measure is closely related to the similarity dimension of IFS.
+Let $m\geq 1$ be a positive integer and endow $\mathbb{R}^{m}$ with the supremum norm $\|\cdot\|_{\mathbb{R}^{m}}$.
+Assume that $T_{i}:\mathbb{R}^{m}\to\mathbb{R}^{m}$ is contracting similitude for $1\leq i\leq l$, i.e., for some $r_{i}\in(0,1)$,
+$\|T_{i}x-T_{i}y\|_{\mathbb{R}^{m}}=r_{i}\|x-y\|_{\mathbb{R}^{m}}.$
+The iterated function system (IFS for short) $\{T_{i}\}_{1\leq i\leq l}$ satisfies the Open Set Condition if there exists a bounded open subset $V\subset\mathbb{R}^{m}$ such that $T_{i}(V)\subset V,\ \forall 1\leq i\leq l$, and $T_{i}(V)\cap T_{j}(V)=\emptyset~\text{for}~i\neq j.$ For instance, the middle-third Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition.
+Under the OSC condition, it is well known that there are (unique) compact subset $K$, which we call the self-similar set, and a measure $\mu$, which we call the self-similar measure, such that
   • (1)
@@
   • (2)
-    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, where $\alpha$ is the unique solution of equation
+    ${\rm dim}_{H}(K,d)={\rm dim}_{B}(K,d)=\alpha$, and $\alpha$ is the unique solution of equation
     $\sum_{i=1}^{l}r_{i}^{\alpha}=1,$
-    and ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. $\alpha$ is called the similarity dimension of the IFS;
+    where ${\rm dim}_{H}$ denotes the Hausdorff dimension under the metric $d:=\|\cdot\|_{\mathbb{R}^{m}}$. The root $\alpha$ is called the similarity dimension of the IFS;
   • (3)
-    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$;
-  • (4)
-    ${\rm supp}(\mu)=K$.
-Many famous fractal sets, for instance, the middle thirds Cantor set, the Sierpinński triangle or gasket, are the attractors of iterated function systems satisfying the Open Set Condition.
-Now take the attractor $K$ obtained above as in the Example 4.5. Then the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$ is invariant with respect to the left shift $\sigma$.
+    $\mu=\sum_{i=1}^{l}r_{i}^{\alpha}\cdot\mu\circ T_{i}^{-1}$ and ${\rm supp}(\mu)=K$.
+Now take the attractor $K$ obtained as above and the product measure $\nu:=\mu^{\otimes\mathbb{Z}}$ on $K^{\mathbb{Z}}$.
 Theorem 4.5.
-Under the setting of Example 4.8, $\nu$ is the maximal metric mean dimension such that the three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
-$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha,$
-where $S\in\{B,P,UC\}$
+The three types of metric mean dimensions of $G_{\nu}$ coincide with the similarity dimension. i.e.,
+$\overline{\rm mdim}_{M}^{S}(\sigma,G_{\nu},d^{\mathbb{Z}})=\alpha=\limsup_{\epsilon\to 0}\frac{\inf_{{\rm diam}(\xi)<\epsilon}h_{\nu}(\sigma,\xi)}{|\log\epsilon|},$
+where $S\in\{B,P,UC\}$.
 Proof.
-Hutchinson [Hut81] showed that for any $z\in K$ and $\epsilon\in(0,1)$, one has
-$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha},$ (4.5)
-where $C>0$ is a constant independent of the choice of the point  $z\in K$ and  $\epsilon$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$ and $\epsilon\in(0,1)$, we define
-[…]
-Then the Bowen ball $B_{n}(x,\epsilon)$ is contained in $L_{n}(x,\epsilon)$ and $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5). Thus, one has
+In [Hut81], Hutchinson proved that there exists a constant $C>0$ such that for any $z\in K$ and $\epsilon\in(0,1)$, one has
+$\mu(B(z,\epsilon))\leq C\epsilon^{\alpha}.$ (4.5)
+Fix $\epsilon\in(0,1)$. For any $x=(x_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}$, we define
+$L_{n}(x,\epsilon)=\left\{(y_{k})_{k\in\mathbb{Z}}\in K^{\mathbb{Z}}:\|x_{k}-y_{k}\|_{\mathbb{R}^{m}}<\epsilon,\ 0\leq k\leq n-1\right\}.$
+Then $B_{n}(x,\epsilon)\subset L_{n}(x,\epsilon)$. Since $\nu(L_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha}$ by (4.5), one has
 $\nu(B_{n}(x,\epsilon))\leq C^{n}\epsilon^{n\alpha},\ \forall x\in K^{\mathbb{Z}},$
-which yields that $\overline{h}_{\nu}^{BK}(\sigma,\epsilon)\geq\alpha|\log\epsilon|-\log C.$ By the Shannon-McMillan-Breiman theorem, the following inequality is valid for any t.d.s. $(X,d,f)$:
+which yields that the Brin-Katok local $\epsilon$-entropy of $\nu$ is bounded blew by
+$\overline{h}_{\nu}^{BK}(\sigma,\epsilon):=\int\limsup_{n\to\infty}\frac{-\log\mu(B_{n}(x,\epsilon))}{n}d\nu\geq\alpha|\log\epsilon|-\log C.$
+Applying the Shannon-McMillan-Breiman theorem, the following inequality is valid for any TDS $(X,d,f)$:
 $\overline{h}_{\mu}^{BK}(f,\epsilon)\leq\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),\ \forall\mu\in M_{f}(X).$ (4.6)
-Hence
+Hence, we have
 […]
 […]
-which implies the desired result. ∎
-Acknowledgments: The first and second authors were supported by NNSF of China (12101340).
+This implies the desired result. ∎
+Acknowledgments: The first and second authors were supported by the National Natural Science Foundation of China (No.
+12101340).
 The third author was supported by the China Postdoctoral Science Foundation (No.
-2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University.
-The authors also thank Prof.
-Ercai Chen for many useful comments.
+2024M763856) and the Postdoctoral Fellowship Program of CPSF (No.
+GZC20252040); Rui Yang would like to thank Dr.
+Yunping Wang for many helpful discussions and the hospitality extended to him during his visit to Ningbo University.
+Special thanks also go to Prof.
+Ercai Chen for his many useful comments.